### v1

# Parameter-Efficient Fine-Tuning of State Space Models

Abstract
Deep State Space Models (SSMs), such as Mamba (Gu & Dao, 2024), have emerged as powerful tools for language modeling, offering high performance with efficient inference and linear scaling in sequence length. However, the application of parameter-efficient fine-tuning (PEFT) methods to SSM-based models remains largely unexplored. This paper aims to systematically study two key questions: (i) How do existing PEFT methods perform on SSM-based models? (ii) Which modules are most effective for fine-tuning? We conduct an empirical benchmark of four basic PEFT methods on SSM-based models. Our findings reveal that prompt-based methods (e.g., prefix-tuning) are no longer effective, an empirical result further supported by theoretical analysis. In contrast, LoRA remains effective for SSM-based models. We further investigate the optimal application of LoRA within these models, demonstrating both theoretically and experimentally that applying LoRA to linear projection matrices without modifying SSM modules yields the best results, as LoRA is not effective at tuning SSM modules. To further improve performance, we introduce LoRA with Selective Dimension tuning (SDLoRA), which selectively updates certain channels and states on SSM modules while applying LoRA to linear projection matrices. Extensive experimental results show that this approach outperforms standard LoRA.
1 Introduction
Over the past two years, Large Language Models (LLMs) such as ChatGPT (Achiam et al., 2023; Brown et al., 2020) have achieved groundbreaking performance and are now widely used in daily life. Many models use the Transformer architecture (Vaswani et al., 2017), with its attention mechanism essential in predicting subsequent tokens based on context. Each token computes attention scores with every preceding one, selectively focusing only on the most relevant during processing. This, however, creates quadratic time complexity, posing challenges when dealing with long sequences. In response, various alternative architectures like Linear Attention (Katharopoulos et al., 2020), RWKV (Peng et al., 2023), RetNet (Sun et al., 2023), and Mamba (Gu & Dao, 2024) have been developed to operate with subquadratic time complexity.
As the most popular subquadratic-time architecture currently serving as an alternative to Transformers, SSMs  (Gu et al., 2021, 2022b, 2022a; Gu & Dao, 2024) achieve efficient training and inference. SSMs are closely related to linear RNNs, which maintain a hidden state to encapsulate the information of previous tokens. When a new input token is introduced, the prediction of the next token involves only operations on this hidden state and the new token, which enhances inference efficiency. To overcome the limitation of RNNs, which cannot be trained in parallel, S4 (Gu et al., 2022b, a) leverages its linearity, enabling it to adopt a convolutional form during training, facilitating parallel computation. Consequently, SSMs are highly efficient and have demonstrated success in numerous long-sequence tasks (Gu et al., 2022b, a). Recently, a new series of SSM models, Mamba (Mamba-I (Gu & Dao, 2024) and Mamba-II (Dao & Gu, 2024)), have achieved Transformer-level performance in language modeling. Although numerous SSM-based models exist, we primarily focus on the deep S4 model and Mamba. The deep S4 model, serving as the foundational architecture, readily extends its properties to other variants, while Mamba has emerged as one of the most popular SSM-based models.
Consequently, we expect fine-tuning these pretrained SSMs for downstream tasks will become a crucial problem in the near future. While fine-tuning the entire model is expensive and inefficient, numerous Parameter-Efficient Fine-Tuning (PEFT) methods (Houlsby et al., 2019; Hu et al., 2021; He et al., 2021; Li & Liang, 2021; Lester et al., 2021; Zaken et al., 2022; Liu et al., 2021, 2022) have been developed for efficient adaptation under resource constraints. Notably, most popular PEFT methods fall into two categories: (i) prompt-based tuning, which involves modifying the input sequence (Lester et al., 2021) or tuning the sequence at each layer (Li & Liang, 2021); and (ii) parameter-based tuning, which directly updates the model parameters, such as LoRA (Hu et al., 2021), which modifies the weight matrices, and BitFit (Zaken et al., 2022), which updates only the bias terms.
Despite the success that existing PEFT methods have achieved in adapting Transformer-based models, their efficacy in adapting SSM-based models remains largely unexplored, leaving many interesting questions open. For instance, are existing popular PEFT methods still effective for SSM-based models? If they are applicable, what is the optimal way to apply these methods to SSM-based models, and which parameters should be updated? If not, can we develop variants specifically tailored for SSMs that perform better? To answer these questions, to the best of our knowledge, we conduct the first comprehensive study of PEFT on SSM-based models, both theoretically and empirically.
To the best of our knowledge, we are the first to benchmark existing PEFT methods on SSM-based models. Through extensive experiments, we demonstrate that (Finding 1) prompt-based PEFT methods are no longer effective for SSM-based models, and (Finding 2) LoRA remains effective on SSM-based models. Meanwhile, the two major components of SSM-based models are the SSM module, which functions analogously to attention in Transformers, and linear projection matrices, which are similar to feed-forward layers. We next investigate which part of the model is more effective for applying PEFT. We empirically find that (Finding 3) applying LoRA to linear projection matrices without modifying the SSM module is already effective, while the most effective linear projection matrices differ depending on the dataset. Notably, Findings 1 and 3 are supported by our theoretical analysis. While LoRA is not effective for tuning SSM modules, theoretically, tuning additional SSM modules increases expressivity. Finally, we analyze the architecture of SSM-based models using the theoretical framework of Giannou et al. (2023) and Zeng & Lee (2024). We show that, in addition to applying LoRA to linear projection matrices, Selectively updating the channel and state Dimensions of SSM modules further enhances performance. We dub this method as SDLoRA, the first PEFT method tailored for SSM-based models. Through extensive experiments, we observe that (Finding 4) SDLoRA outperforms LoRA alone in fine-tuning SSM-based models.
2 Related Works
State Space Models (SSMs).
Linear State-Space Layers (LSSL) represent one of the earliest SSM layers utilized in deep learning, functioning as continuous-time, recurrent, and convolutional models (Gu et al., 2021). LSSL employs HiPPO theory (Gu et al., 2020) to initialize the state matrix ${\bm{A}}$, enabling the capture of long dependencies. However, LSSL is computationally expensive, limiting its practical application. Gu et al. (2022b) introduced Structured State Space Models (S4), which optimize computation efficiency by employing a structured state matrix ${\bm{A}}$. Gupta et al. (2022) proposed DSS, which simplifies the model by using a diagonal matrix for ${\bm{A}}$ and empirically demonstrated that it suffices to achieve performance comparable to S4. Further, Gu et al. (2022a) provided a theoretical explanation for the effectiveness of the diagonal state matrix ${\bm{A}}$ in DSS and introduced S4D, which offers various initialization methods for ${\bm{A}}$. Subsequently, the diagonal structure of the state matrix ${\bm{A}}$ has been adopted in follow-up methods (Gu & Dao, 2024). Despite differences in optimization algorithms, we refer to S4 and its close variants, including DSS and S4D, collectively as S4. This terminology encompasses models that maintain the standard discrete-time SSM form with a diagonal state matrix.
Despite of the remarkable performance of SSMs on certain tasks of sequence modeling, SSMs still showed worse performance than Transformers on language modeling. Fu et al. (2022) transitioned from synthetic language modeling tasks to real language modeling tasks with SSMs. They proposed H3, which is inspired by Linear Attention (Katharopoulos et al., 2020), introducing both diagonal SSM and shift SSM. Recently, Mamba (Gu & Dao, 2024; Dao & Gu, 2024) escaped from linear time invariance (LTI) modeling by introducing input-dependent terms and achieved better performance than Transformer on language modeling. Furthermore, several hybrid models (Lieber et al., 2024; Park et al., 2024) tried to exploit the advantages of both SSMs and Transformers.
Parameter-Efficient Fine-Tuning (PEFT).
Due to the increase in model size, PEFT methods have gained increasing popularity as they achieve good performance while being much more efficient compared to full model updating (Houlsby et al., 2019; Hu et al., 2021; He et al., 2021; Li & Liang, 2021; Lester et al., 2021; Zaken et al., 2022; Liu et al., 2021, 2022). Most of the existing popular PEFT methods fall into two categories: (i) prompt-based methods (Li & Liang, 2021; Lester et al., 2021; Liu et al., 2021, 2022), and (ii) parameter tuning methods (Donahue et al., 2014; Yosinski et al., 2014; Hu et al., 2021; Zaken et al., 2022). Common prompt-based methods include prompt tuning (Lester et al., 2021), and prefix-tuning (Li & Liang, 2021). Prompt tuning prepends a sequence of learnable virtual tokens, which are continuous vectors. Prefix-tuning further expands on Prompt tuning by prepending tokens across the model’s depth, making it more powerful. Therefore, our analysis of prompt-based methods’ limitations will focus on prefix-tuning, with the findings also applicable to the other prompt-based methods. Conversely, parameter tuning methods, which originated from traditional transfer learning practices, typically involve freezing the initial layers and only tuning the last few layers (Donahue et al., 2014; Yosinski et al., 2014). In recent years, more effective and innovative parameter tuning approaches have emerged (Hu et al., 2021; Zaken et al., 2022). The widely used Low-Rank Adaptation (LoRA) updates a subset of parameters (e.g., attention layers of a Transformer) in a low-rank manner. BitFit (Zaken et al., 2022), focuses on tuning only the bias terms of a pretrained model. In Sec. A, we provide a more detailed description of these baseline methods.
Numerous efforts have been made to theoretically understand existing PEFT methods. For prompt-based methods, Wang et al. (2023b), Petrov et al. (2024), and Oymak et al. (2023) have theoretically analyzed the effectiveness and limitations of prompt tuning and prefix-tuning for Transformer-based models. For LoRA, Zeng & Lee (2024) explored its expressive power by demonstrating that even a randomly initialized model can be adapted to match any smaller target model using LoRA. Some of our theoretical analysis draws upon the framework established by Zeng & Lee (2024). Jang et al. (2024) conducted a theoretical exploration of LoRA within the neural tangent kernel (NTK) regime.
3 Preliminaries of State Space Models
Scalar-input Scalar-output SSM.
[…]
To adapt SSMs for deep learning, the continuous parameters $({\bm{A}},{\bm{B}})$ are transformed into discrete counterparts $(\overline{{\bm{A}}},\overline{{\bm{B}}})$ using a learnable step size $\Delta\in{\mathbb{R}}$. An example of a discretization rule is the zero-order hold, which defines $\overline{{\bm{A}}}=\exp(\Delta{\bm{A}}),\overline{{\bm{B}}}=(\Delta{\bm{A}})^{-1}(\exp(\Delta{\bm{A}})-{\bm{I}})\cdot\Delta{\bm{B}}$.
The discrete-time SSM is formulated as (2). For efficient and parallelizable training, the output ${\bm{y}}$ of a length-$N$ input ${\bm{x}}$ in the discrete-time SSM can be computed with a long convolution, as detailed in (3). This convolution operation can be efficiently computed in the frequency domain with FFT.
Vector-input Vector-output SSM.
Many deep learning tasks, such as language modeling, often use multi-channel inputs. When the input and output are vectors, denoted as $\bm{x},\bm{y}\in{\mathbb{R}}^{D}$, separate SSMs are used for each of the $D$ input channels. As such, a superscript $(d)$ is introduced to indicate parameters specific to each channel when necessary. This notation may be omitted for simplicity.
Structured State Space Sequence Model (S4).
S4 introduced by Gu et al. (2022b) represents one of the earliest applications of SSMs in deep learning. It features a diagonal structure for the state matrix ${\bm{A}}$, a design theoretically validated by Gu et al. (2022b) and practically implemented through its subsequent variants, DSS (Gupta et al., 2022) and S4D (Gu et al., 2022a).
Deep S4 Layer.
[…]
$\bm{y}_{t}=\operatorname{ReLU}({\bm{W}}\cdot\operatorname{S4}_{t}(\bm{x}_{1},\ldots,\bm{x}_{t})+{\bm{\beta}}+{\bm{u}}\otimes\bm{x}_{t}),$ (4)
where ${\bm{W}}\in{\mathbb{R}}^{D\times D}$ and ${\bm{\beta}}\in{\mathbb{R}}^{D}$ represent the linear projection matrix and bias, respectively, and ${\bm{u}}\in\mathbb{R}^{D}$ is the coefficient of the residual connection. Note that in a deep S4 layer, the trainable parameters are SSM parameters $({\bm{A}}^{(d)},{\bm{B}}^{(d)},{\bm{C}}^{(d)},\Delta^{(d)})$ across $D$ channels with ${\bm{A}}^{(d)}$ being diagonal and the parameters $({\bm{W}},{\bm{\beta}})$ for the linear layer and ${\bm{u}}$ for the residual connection.
Selective State Space Models (S6).
A key property of all SSMs mentioned above is linear time invariance (LTI), where model dynamics remain constant over time. However, LTI models face significant limitations: their constant dynamics fail to selectively extract relevant information from the context or influence the hidden state in an input-dependent manner. The S6 model, proposed by Gu & Dao (2024), addresses these limitations by making its parameters input-dependent.
In particular, at each time step $t$, given the input $\bm{x}_{t}\in{\mathbb{R}}^{D}$, they introduce input-dependency to step size $\bm{\Delta}_{t}=(\Delta^{(1)}_{t},\ldots,\Delta^{(D)}_{t})^{\top}\in{\mathbb{R}}^{D}$, input transition vectors ${\bm{B}}_{t}\in{\mathbb{R}}^{H\times 1}$ and the output mapping vectors ${\bm{C}}_{t}\in{\mathbb{R}}^{1\times H}$ via linear projection:
$\bm{\Delta}_{t}=\operatorname{softplus}(\bm{W}_{\bm{\Delta}}\bm{x}_{t}+{\bm{\beta}}_{\bm{\Delta}}),\quad{\bm{B}}_{t}={\bm{W}}_{{\bm{B}}}\bm{x}_{t},\quad{\bm{C}}_{t}={\bm{W}}_{{\bm{C}}}\bm{x}_{t},$ (5)
whereas the diagonal state matrices ${\bm{A}}^{(1)},\ldots,{\bm{A}}^{(D)}$ remain input-independent. Note that $\bm{W}_{\bm{\Delta}}\in{\mathbb{R}}^{D\times D}$ is implemented via a rank-$r$ low-rank parameterization, denoted by $\bm{W}_{\bm{\Delta}}={\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\bm{\Delta},\downarrow}$, where ${\bm{W}}_{\bm{\Delta},\uparrow}\in{\mathbb{R}}^{D\times r}$ and ${\bm{W}}_{\bm{\Delta},\downarrow}\in{\mathbb{R}}^{r\times D}$, which is a common method for reducing compute overheads (Wang et al., 2021, 2023a). To summarize, the trainable parameters in S6 include state matrices ${\bm{A}}^{(d)}$ across $D$ channels, parameters ${\bm{W}}_{\bm{\Delta},\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{\beta}}_{\bm{\Delta}}$ for computing $\bm{\Delta}_{t}$, and weight matrices ${\bm{W}}_{{\bm{B}}},{\bm{W}}_{{\bm{C}}}\in{\mathbb{R}}^{H\times D}$ for computing ${\bm{B}}_{t},{\bm{C}}_{t}$. The state matrices and the input transition vectors of S6 are then discretized according to $\overline{{\bm{A}}}_{t}^{(d)}=\exp(\Delta_{t}^{(d)}{\bm{A}}^{(d)}),\overline{{\bm{B}}}_{t}^{(d)}=\Delta_{t}^{(d)}{\bm{B}}_{t}.$ In contrast to S4, where $\overline{{\bm{B}}}^{(d)}$ varies independently across channels, the differences in $\overline{{\bm{B}}}^{(d)}$ in S6 are solely due to the scalar $\Delta_{t}^{(d)}$. Additionally, S6 uses the same ${\bm{C}}_{t}$ for all channels at each time step $t$, unlike S4, which has unique ${\bm{C}}^{(d)}$ for each channel.
Mamba.
Similar to the Transformer block, which consists of attention and linear layers, the Mamba block proposed by Gu & Dao (2024) features an S6 module, a point-wise 1D causal convolution layer (Conv1d) for token mixing, linear layers — including input ($\bm{W}_{\text{in}}$) and output ($\bm{W}_{\text{out}}$) projection layers and a gated MLP. Mamba, primarily allocating its parameters in $\bm{W}_{\text{in}}$ and $\bm{W}_{\text{out}}$, is inspired by H3 (Fu et al., 2022).
4 Benchmarking PEFT Methods on SSM-based Models
In this section, we examine the effectiveness of popular PEFT methods when applied naively to SSM-based models, specifically Mamba (130M and 1.4B).
Experiment Setup.
We consider two main categories of PEFT methods: parameter-based and prompt-based. From each category, we evaluate two representative methods. For parameter-based methods, we select BitFit (Zaken et al., 2022) and LoRA (Hu et al., 2021). For prompt-based methods, we choose prefix-tuning (Li & Liang, 2021) and prompt tuning (Lester et al., 2021). For BitFit, fine-tuning is performed on all bias terms present in the Mamba architecture, specifically the biases of the Conv1d and the linear projection layer of step size $\bm{\Delta}$. For prefix-tuning, we adopted the huggingface implementation (Mangrulkar et al., 2022) to construct a MLP, employing the overparameterization technique to ensure stable optimization.
We consider five datasets spanning diverse domains: the GLUE natural language understanding benchmark (Wang et al., 2019), the DART RDF-to-text generation benchmark (Nan et al., 2021), the Spider text-to-SQL generation benchmark (Yu et al., 2018), and CIFAR-10 for computer vision tasks (Krizhevsky et al., 2009). A more detailed introduction of the datasets considered in this paper is provided in Sec. B. Notably, prefix-tuning requires substantially more parameters than other PEFT methods, as it employs a multilayer perceptron at each layer to project a fixed sequence into soft tokens for training stability. For all other PEFT methods, we constrain the trainable parameters to fewer than 0.5% for language tasks and 1% for vision tasks, ensuring a fair comparison. The higher allowance for vision tasks accommodates the need for extensive fine-tuning for new modalities. Consequently, LoRA adapters are applied exclusively to linear projection matrices, leaving the SSM modules unchanged to comply with these parameter constraints.
Results.
Table 1 presents our results. Parameter-based PEFT methods generally outperform prompt-based methods significantly, despite using the same number of trainable parameters—except for prefix-tuning, which underperforms despite using more parameters. LoRA consistently achieves the best performance across all tasks and metrics, occasionally surpassing full fine-tuning while tuning less than 1% of parameters. We provide the full experimental results, including detailed results of subtasks on GLUE and Spider, in Sec. C.2.
These findings above raise two critical questions: (i) Why do existing prompt-based PEFT methods lose effectiveness when applied to SSM-based models? (ii) Can LoRA achieve better performance when applying on both linear projection matrices and SSM modules? To address these questions, we conduct both theoretical analysis and further empirical studies on prompt-based PEFT methods and LoRA in the context of SSMs.
Dataset GLUE DART SAMSum Spider CIFAR-10
Metric ($\uparrow$) Avg. Score METEOR BLEU R1 R2 RL Acc. Acc.
Prompt Tuning 63.8 66.2 39.8 50.1 25.6 41.6 43.6 30.4
Prefix-Tuning 68.6 66.6 42.5 50.6 26.5 42.1 39.7 41.0
BitFit 76.8 67.0 43.7 50.3 25.7 41.9 48.4 44.4
LoRA (Linear Projection Matrices) 80.5 70.4 49.1 50.9 27.0 42.3 57.5 61.1
Full Fine-Tuning 80.5 71.0 51.8 51.2 27.3 42.9 66.2 60.0
Table 1: Benchmarking popular Parameter-Efficient Fine-Tuning (PEFT) methods on five real-world datasets. R1/R2/RL stand for ROUGE-1/2/L. For all PEFT methods except prefix-tuning, we report the best results for cases where fewer than 0.5% of parameters are tunable for language tasks and fewer than 1% for vision tasks (i.e., CIFAR-10) after comprehensive hyperparameter search. Prefix-tuning is an exception, as it requires training a multilayer perceptron at each layer to project a fixed sequence into soft tokens for training stability, consuming more trainable parameters than our threshold. Bold numbers indicate outperformance over all PEFT methods, while underlined numbers indicate outperformance over full fine-tuning.
4.1 Limitations of Applying Existing Prompt-based Methods on SSMs
This part addresses our first question arised in this section: Why do existing prompt-based PEFT methods lose effectiveness when applied to SSM-based models? We approach this by establishing an upper bound on the performance of existing prompt-based PEFT methods.
A key feature of SSMs is their next-token prediction mechanism, which relies solely on the current token and hidden states, without considering previous tokens directly. The hidden states encapsulate all information from preceding tokens. Consequently, prepending tokens to an SSM is functionally equivalent to tuning the initial state, as demonstrated by the following proposition. The formal version and proof of Proposition 1 are presented in Section C.3.
Proposition 1 (Informal: Expressivity of Prefix-Tuning on SSMs).
The maximum expressiveness achievable via prefix-tuning on SSMs is equivalent to the expressiveness of solely tuning the initial hidden state $\bm{h}_{0}$.
To evaluate the performance of initial state tuning, we conducted experiments on the GLUE benchmark, comparing prompt-tuning, prefix-tuning, initial state tuning, and LoRA across seven GLUE tasks. Table 2 presents our findings. The results demonstrate that initial state tuning generally outperforms prefix-tuning, corroborating our analysis. However, LoRA significantly surpasses initial state tuning in performance. These observations lead us to conclude that the limitations of initial state tuning inherently constrain the potential of existing prompt-based methods, preventing them from outperforming LoRA in the context of SSM-based models. While the reason for the underperformance of initial state tuning is unclear, we identify explaining it as an interesting direction for future research. Nevertheless, we propose a plausible explanation. We hypothesize that SSM’s exclusive reliance on hidden states, without direct access to previous tokens or states, severely restricts the impact of initial state tuning, particularly for long sequences. This aligns with the findings of Fu et al. (2022), which demonstrate SSM’s limitations in recalling older tokens.
Task RTE MRPC CoLA SST-2 QNLI QQP MNLI Avg. Score
Prompt Tuning .560 .716 .120 .894 .768 .796 .615 .638
Prefix-Tuning .695 .757 .434 .915 .834 .831 .356 .686
Initial State Tuning .668 .751 .524 .924 .864 .861 .785 .768
LoRA (Linear Projection Matrices) .704 .828 .606 .924 .884 .877 .815 .805
Table 2: Comparison of prompt-tuning, prefix-tuning, initial state tuning, and LoRA on seven tasks from the GLUE benchmark. We report the Matthews correlation ($\uparrow$) for CoLA, overall (matched and mismatched) accuracy ()($\uparrow$)( ↑ ) for MNLI, and accuracy for other tasks. Initial State Tuning and LoRA are constrained to use less than 0.5% trainable parameters. Bold numbers indicate the best performance across all three methods, while underlined numbers show the highest score among prompt-based methods (prefix-tuning and initial state tuning). Initial state tuning outperforms prefix-tuning and prompt-tuning on five out of seven tasks, while LoRA consistently outperforms all prompt-based methods.
4.2 Optimal Application of LoRA in SSM-based Models
In our previous experiments, we applied LoRA exclusively to linear projection matrices. However, SSM-based models typically comprise various modules, including S4 (convolution layer), S6, and multiple distinct linear projection matrices. To investigate the impact of applying LoRA to different components, we conduct a comprehensive study across five datasets.
Model Mamba-130M Mamba-1.4B
Dataset Params. (%) GLUE DART CIFAR-10 Params. (%) SAMSum Spider
Metric ($\uparrow$) Avg. Score METEOR BLEU Acc. R1 R2 RL Acc.
SSM Modules .92 79.3 69.9 50.8 44.0 .46 50.5 26.4 42.2 56.3
Linear Projection Matrices 1.02 80.5 71.2 49.2 62.8 .51 50.8 26.9 42.8 54.7
Both 1.92 80.2 71.0 49.5 60.4 .97 50.8 26.6 42.7 56.4
Table 3: For LoRA, targeting only the linear projection matrices yields better performance than applying it to all modules in Mamba. Consistent rank is maintained across all three methods.
We examine LoRA’s performance when applied to SSM modules and linear projection matrices separately, as well as in combination. For linear projections, we test LoRA on all possible matrices. For SSM modules, we apply LoRA to all weight matrices (e.g., weight matrices of input-dependent step size $\bm{\Delta}$) in SSM modules. For the state transition matrices ${\bm{A}}$, given their diagonal structure for each channel, we treat them as vectors, concatenate the channels into a matrix, and apply LoRA. The results are presented in Table 3. We observe that applying LoRA to linear projection matrices achieves superior performance on six out of eight metrics. Interestingly, additional tuning of SSM modules lead to decreased performance in some cases. This suggests that LoRA might not be well-suited for tuning SSM modules, while being highly effective for linear projection matrices.
To further elucidate this concept, we present the following lemma, which examines a simplified model architecture consisting of S6 with a linear input projection matrix at each layer. We demonstrate that fine-tuning the projection matrix $\bm{W}_{\text{in}}$ encompasses the expressivity of fine-tuning the parameters $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\uparrow}$.
Lemma 2 (Expressivity of Fine-Tuning Projection Matrices).
Consider an S6 with an additional linear input projection matrix $\bm{W}_{\text{in}}$. Denote the input-dependent SSM parameters $\{\{\overline{{\bm{A}}}^{(d)}_{n}\}_{d=1}^{D},\overline{{\bm{B}}}_{n},{\bm{C}}_{n}\}_{n=1}^{N}$ as ${\bm{\theta}}(\cdot;\{{\bm{A}}^{(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow},\bm{W}_{\text{in}}).$ For any given $\overline{\bm{W}}_{{\bm{B}}}$, $\overline{\bm{W}}_{{\bm{C}}}$, and $\overline{\bm{W}}_{\Delta,\uparrow}$, there exists a $\widehat{\bm{W}_{\text{in}}}$ such that for any input sequences ${\bm{X}}\in{\mathbb{R}}^{D\times N}$,
${\bm{\theta}}({\bm{X}};\{{\bm{A}}^{(d)}\}_{d=1}^{D},\overline{\bm{W}}_{{\bm{B}}},\overline{\bm{W}}_{{\bm{C}}},\overline{\bm{W}}_{\Delta,\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow})={\bm{\theta}}({\bm{X}};\{{\bm{A}}^{(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow},\widehat{{\bm{W}}}_{\text{in}}).$ (6)
We expand upon this section in Sec. C.4, where we provide more detailed statements of the above assertion and its corresponding proofs. Additionally, we empirically examine applying LoRA to different weight matrices of Mamba, which incorporates multiple linear projection matrices in each layer, including output projection matrices $\bm{W}_{\text{out}}$ after the S6 module and input projection matrices $\bm{W}_{\text{in}}$ before the gating and convolutional layer. Our experiment results, however, reveal that applying LoRA to different matrices achieves similar performance, as detailed in Sec. C.4.
5 Dimension Selection for Tuning State-Space Models
In Sec. 4.2, we demonstrate the efficacy of LoRA in fine-tuning linear projection matrices. Theoretically, fine-tuning all components should offer greater expressive power. However, Table 3 indicates that applying LoRA to SSM modules might paradoxically decrease performance. Therefore, we aim to develop an algorithm specifically tailored for tuning SSM modules. To achieve this, we first seek to understand the relative importance of different parameters within SSM modules.
5.1 Understanding the Roles of State Matrix A, Input Transition Vector B, and Output Mapping Vector C for a Single Channel in S4 Modules
Problem Setting.
Inspired by Zeng & Lee (2024)’s theoretical analysis of LoRA’s expressive power, we adopt a similar framework to explore the expressive potential of various parameters in the S4 model. Specifically, we assume a target model that performs well on the intended task and a frozen model, which may be either pretrained or randomly initialized. Our goal is to identify a parameter-efficient method to update the frozen model so that it becomes functionally equivalent to the target model. In alignment with Zeng & Lee (2024), we assume that the frozen model’s capacity is equal to or exceeds that of the target model. This assumption is based on two main considerations: (i) analytical tractability, which necessitates that the frozen model must have the potential to match the functionality of the target model, and (ii) a practical rationale, given that the models typically used in practice are often overparameterized. Assume that both the target model and the frozen model are S4, with the target model having a hidden state dimension $H_{\star}$ and the frozen model having a hidden state dimension $H\geq H_{\star}$. Meanwhile, suppose that all the hidden dimensions of both models are valid, meaning that none of the parameter elements are zero. The target model, frozen model, and the updated model after tuning the parameters on the frozen model can be formulated using discretized parameters $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$ as follows:
(Target model) $f^{\star}({\bm{x}})_{n}=\sum\nolimits_{m=1}^{n}{\bm{C}}_{\star}\overline{{\bm{A}}}_{\star}^{m-n}\overline{{\bm{B}}}_{\star}x_{m},\text{ where}\operatorname{diag}(\overline{{\bm{A}}}_{\star}),\overline{{\bm{B}}}_{\star},{\bm{C}}_{\star}\in{\mathbb{R}}^{H_{\star}},$ (7)
(Frozen model) $f_{0}({\bm{x}})_{n}=\sum\nolimits_{m=1}^{n}{\bm{C}}\overline{{\bm{A}}}^{m-n}\overline{{\bm{B}}}x_{m},\text{ where }\operatorname{diag}(\overline{{\bm{A}}}),\overline{{\bm{B}}},{\bm{C}}\in{\mathbb{R}}^{H},$ (8)
(Updated model) $\hat{f}({\bm{x}})_{n}=\sum\nolimits_{m=1}^{n}\widehat{{\bm{C}}}\widehat{\overline{{\bm{A}}}}^{m-n}\widehat{\overline{{\bm{B}}}}x_{m},\text{ where }\operatorname{diag}(\widehat{\overline{{\bm{A}}}}),\widehat{\overline{{\bm{B}}}},\widehat{{\bm{C}}}\in{\mathbb{R}}^{H}.\vspace{-.1in}$ (9)
Parameter Efficiency Analysis on S4.
Let ${\mathcal{P}}^{H}$ denote the set of all $H\times H$ permutation matrices. Given this formulation, we present our first analysis of parameter efficiency for the S4 model in the following lemma. This analysis is based on the parameters after necessary discretization $(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})$.
Lemma 3 (Essential Discretized Parameter Set for S4).
Consider the parameters after discretization, i.e., $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\otimes\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
This lemma highlights the significance of identifying essential hidden state dimensions. The term $\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\otimes\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}$ underscores the importance of excluding redundant dimensions. This can be achieved by either directly removing these dimensions from the state matrix $\overline{{\bm{A}}}$, or by updating $\overline{{\bm{B}}}$ or ${\bm{C}}$ to ensure that only the selected hidden state dimensions are utilized during the input transition or output mapping phases. Once redundant dimensions are filtered out, tuning only the essential dimensions is sufficient to align the updated model with the target model.
5.2 SSM Dimension Selection Algorithm
Inspired by Lemma 3, we introduce the Dimension Selection algorithm to construct adapters on SSMs for fine-tuning. This algorithm first selects unimportant dimensions and sets them to zero, filtering out irrelevant information based on Lemma 3. For enhanced parameter efficiency, we then update only the most important channels and state dimensions within these selected subsets. Regardless of other selections, we consistently tune the coefficients of residual connections and biases in linear projections, as these components contain a negligible number of parameters. However, we will later demonstrate that in practice, tuning residual connections and biases is unnecessary. The detailed pseudo-code is presented in Alg. 1. Given that tuning ${\bm{C}}$ alone is as effective as tuning both ${\bm{B}}$ and ${\bm{C}}$ for S4 (Gupta et al., 2022), subsequent discussions on S4 will focus solely on ${\bm{C}}$, excluding $\{{\bm{B}}^{(d)}\}_{d=1}^{D}$ for simplicity, without loss of generality.
Input: Dataset ${\mathcal{D}}$, warmup epochs $E_{0}$, train epochs $E$, number of layers $L$, total channels $D$, total states $H$, initial state sparsity $\beta_{0}$, initial channel sparsity $\alpha_{0}$, state update fraction $\beta$, channel update fraction $\alpha$
Output: Model adapter
/* Warmup Epochs */
1 Update SSM modules using ${\mathcal{D}}$ for $E_{0}$ epochs;
/* Setup Adapters */
2 for $l=1$ to $L$ do
       /* Set dimensions as zero */
3       Sort channels based on magnitude of $\overline{{\bm{A}}}^{(d)}$ at each channel;
4       Set final $(1-\beta_{0})D$ as zero by letting ${\bm{C}}^{(d)}={\bm{0}}$, denote non-zero channels as set ${\mathbb{D}}$;
5       for $d\in{\mathbb{D}}$ do
6             Sort states based on magnitude of $\bar{A}^{(d)}_{h}$ at each state dimension;
7             Set final $(1-\alpha_{0})H$ as zero by letting corresponding $C^{(d)}_{h}=0$, denote non-zero states as set ${\mathbb{H}}$;
8
      /* Unfreeze dimensions */
9       Sort non-zero channels ${\mathbb{D}}$ based on magnitude of parameter changes at each channel;
10       Denote first $\beta|{\mathbb{D}}|$ as ${\mathbb{D}}^{\prime}$;
11       for $d\in{\mathbb{D}}^{\prime}$ do
12             Sort non-zero state dimensions based on magnitude of parameter changes;
13             Construct adapter to update first $\alpha|{\mathbb{H}}|$ states at $d$-th channel;
14
      /* Include Residual Connections and Bias */
15       Construct adapter for all residual connections and bias;
16
Algorithm 1 Dimension Selection Algorithm for S4
We refer to our method as SDLoRA. This approach extends beyond applying LoRA to linear projection matrices by Selectively updating certain subset of channels and states Dimensions, which are chosen by Alg. 1. Our analysis considers cases where each input token $x_{t}\in{\mathcal{X}}$, with ${\mathcal{X}}\in{\mathbb{R}}^{D}$ bounded, and the input sequence length is finite. The following theorem elucidates the expressive capacity of SDLoRA on deep S4 models. For proof and additional details, refer to Sec. D.2.
Theorem 4 (Expressive Power of SDLoRA on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDLoRA, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states on SSM modules, applying rank-$\lceil\frac{L}{L^{\star}}\rceil$ updates on linear projection matrices and updating residual connections and biases at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
This theorem demonstrates that a larger pretrained model requires selecting fewer channels and hidden states at each layer. Furthermore, if the target task is less complex — evidenced by a smaller target model with fewer layers $L^{\star}$ and hidden states $H^{\star}$ — the number of channels and hidden states needed is also reduced. This finding aligns with the theoretical analysis of LoRA presented in Zeng & Lee (2024), which shows that larger pretrained models require fewer learnable parameters (referred to as “lower rank” in their context) during fine-tuning, especially for simpler tasks. Although this theorem is constrained by the assumptions of linear activations and the absence of residual connections in the target model, while also requiring fully fine-tuning the linear project matrix of last layer, our findings have broader implications. Our following experimental results suggest that these findings generalize beyond these restrictions.
5.3 Empirical Evaluation on Deep S4 Models
In this experiment, we seek to validate the theoretical guarantees for SDLoRA under more general conditions, including residual connections in the target model and ReLU activations in both frozen model and target model, without full fine-tuning the linear projection matrix of the last layer. Additionally, we assess SDLoRA’s empirical performance on both synthetic and real datasets. More experiment setup details are provided in Sec. D.3.
Figure 1: Approximation error of PEFT methods on deep S4 models for synthetic experiments.
Method # Params (%) Accuracy
Frozen 0.00 73.9
LoRA (Proj) 16.00 77.6
LoRA (S4+Proj) 15.52 77.6
SDLoRA 11.17 78.0
Full Fine-Tuning 100.00 77.6
Table 4: Accuracy comparison between SDLoRA and LoRA on deep S4 models for CIFAR-10 (Krizhevsky et al., 2009).
Synthetic Dataset.
For the synthetic dataset, we employ a regression setting to validate our theoretical results. (Experiment Setting) We randomly initialize two models: a one-layer deep S4 model as the target and a four-layer deep S4 model as the frozen model. The task is to update the frozen model to match the functionality of the target model. We generate an input sequence ${\bm{X}}$ of length $200$ and dimension $64$, with values uniformly drawn from integers between 0 and 9. This input is then processed through the target model to obtain the corresponding outputs. These input-output pairs are used to train the frozen model over 500 iterations using the Mean Squared Error (MSE) loss.
(Results) Figure 1 displays the MSE, averaged across all tokens, plotted against the trainable parameters of different methods. We observe that by using only $\approx 28\%$ of the total parameters of the frozen S4 model, SDLoRA closely approximates the performance of the target S4 model, achieving results comparable to full fine-tuning, thereby substantiating our theorem. Meanwhile, we observe that SDLoRA outperforms both the approach of applying LoRA solely to linear projection matrices and the approach of applying LoRA to both the S4 module and linear projection matrices. In this latter approach, the diagonal vectors of state matrices ${\bm{A}}^{(d)}$, input transition vectors ${\bm{B}}^{(d)}$ and output mapping vectors ${\bm{C}}^{(d)}$ are naively concatenated across $D$ channels into three $D\times H$ matrices before low-rank updates are applied. In Sec. D.3, we also evaluate an extension of SDLoRA that performs sparse tuning on the linear projection matrices by updating only the columns corresponding to the channels selected by Alg.1, instead of applying LoRA. This extension shows promising results.
CIFAR-10.
We also conduct experiments on the CIFAR-10 dataset (Krizhevsky et al., 2009). We employ an eight-layer deep S4 model with a hidden state dimension of 16 and a model dimension of 64. Since pretrained deep S4 models are not available, we simulate a pretrained scenario by fully updating the model for 50 epochs first, then subsequently evaluating the PEFT methods over an additional 5 epochs. The results, as reported in Table 4, indicate that SDLoRA outperforms LoRA with fewer trainable parameters.
5.4 Empirical Evaluation on Mamba
Lastly, we conduct experiments on pretrained Mamba models. We consider four datasets, using Mamba-130M for GLUE and DART, and Mamba-1.4B for SAMSum and Spider. We evaluate three configurations each for LoRA and SDLoRA, applying LoRA to distinct parameter subsets and varying SDLoRA’s state freeze ratios while maintaining a 99% channel freeze ratio. In this experiment, we allow channels and states to learn directly from the datasets without manually setting any to zero. We then select a LoRA-rank such that all configurations have a similar number of trainable parameters for a fair comparison. Residual connections and biases are frozen in this experiment. All reported values represent averages across three simulations, with learning rates independently selected for each simulation. For more details, please see Sec. D.4. The experimental results are reported in Table 5. The results demonstrate that SDLoRA outperforms LoRA alone for fine-tuning the SSM even when 99% of the channels are frozen. This result underscores the efficacy of SDLoRA.
Model Mamba-130M Mamba-1.4B
Method Params (%) GLUE DART Params (%) SAMSum Spider
Avg. Score ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) R1 ($\uparrow$) R2 ($\uparrow$) RL ($\uparrow$) Acc. ($\uparrow$)
Val Test Val Test Val Test Val Test Val Test Val Test Val Test
LoRA .3178 80.71 78.74 50.44 41.27 70.00 65.84 .1594 51.59 50.56 27.66 26.49 42.87 42.22 82.08 61.19
.3600 80.79 79.39 51.03 42.02 70.16 66.18 .1810 51.61 51.03 28.15 26.81 43.18 42.36 83.52 62.64
.3883 80.39 79.49 50.70 41.55 69.83 65.98 .1947 51.48 50.90 27.90 26.63 43.26 42.41 82.98 59.25
SDLoRA .3492 80.93 79.75 51.45 42.37 70.45 66.60 .1760 51.63 50.90 27.97 26.86 43.32 42.52 84.36 62.57
.3498 81.05 79.16 51.47 43.85 70.46 66.38 .1761 51.61 50.76 28.02 26.65 43.38 42.29 84.48 59.96
.3509 80.67 78.73 51.54 42.56 70.45 66.45 .1764 51.74 50.86 28.08 26.54 43.39 42.19 84.19 61.25
Table 5: Performance comparison between SDLoRA and LoRA on pretrained Mamba models. Bold numbers indicate the best performance for each task. Underlined numbers indicate that the model outperforms all models fine-tuned via the alternative method for the same task (e.g., SDLoRA outperforms all LoRA methods, or vice versa). On Mamba-130M, we compare the performance of SDLoRA and LoRA on GLUE (Wang et al., 2019) and DART (Nan et al., 2021) benchmarks. On Mamba-1.4B, we compare performance of SDLoRA and LoRA on SAMSum (Gliwa et al., 2019) and Spider (Yu et al., 2018) benchmarks. R1, R2, and RL represent ROUGE-1, ROUGE-2, and ROUGE-L, respectively.
6 Conclusion & Discussion
In this paper, we present the first study on the performance of PEFT methods applied to SSM-based models. Our evaluation of existing PEFT methods provides valuable insights and guidelines for future researchers to parameter-efficiently fine-tune SSM-based models to other domains. Moreover, we take a first step in establishing a theoretical framework for studying PEFT methods on SSM-based models. Furthermore, we introduce SDLoRA, the first PEFT method specifically tailored for SSM-based models, which outperforms existing methods.
While our work offers numerous valuable insights, some limitations exist. Theoretically, our guarantees for SDLoRA are limited to linear activations and require full fine-tuning of the last layer. However, our experiments demonstrate that SDLoRA does not suffer from these limitations in practice. Removing such restrictions for SDLoRA in theory or developing new PEFT methods under more general theoretical cases is an interesting future direction. Additionally, our theory only demonstrates that updating a subset of channels and states is sufficient, without providing guidance on optimal selection. Our channel and state selection, based on a warmup stage and parameter magnitude, may not be optimal. Further investigation into the impact of channel/state selection and development of improved dimension selection algorithms presents an interesting avenue for future work. Lastly, our work primarily focuses on SSM-based models. Studying PEFT methods on SSM-Transformer hybrid models (Lieber et al., 2024; Park et al., 2024), is an interesting future direction.
Acknowledgement
The work of Kangwook Lee is supported in part by NSF CAREER Award CCF-2339978, Amazon Research Award, and a grant from FuriosaAI.
Reproducibility Statement
The code for all experiments reported in this paper is publicly accessible. For the purpose of reproducibility, the code can be found at the following GitHub repository: https://github.com/furiosa-ai/ssm-peft.
Appendix
[…]
Appendix A In-depth Introduction of Baselines
In this section, we provide a more detailed description of the baseline methods.
LoRA (Hu et al., 2021).
LoRA (Low-Rank Adaptation) focuses on fine-tuning large models by freezing most of the pretrained parameters and injecting trainable low-rank matrices into each layer of the Transformer’s architecture. The intuition behind using low-rank matrices comes from linear algebra, where a large matrix can be closely approximated by the product of two smaller matrices. The number of trainable parameters can be controlled with the rank of the low-rank matrices. LoRA also uses a scaling parameter (LoRA alpha) for the weight matrices to control the balance of the original model weights and LoRA weights during training. After fine-tuning, LoRA weights can be merged with the original model weights, introducing no additional inference overhead.
Prompt Tuning (Lester et al., 2021).
Prompt tuning freezes all model weights and prepends a trainable soft prompt to the input prompt. The soft prompt consists of trainable virtual tokens, which are continuous. At inference time, prompt tuning introduces an inference overhead based on the number of virtual tokens used.
Prefix-Tuning (Li & Liang, 2021).
Prefix-tuning also prepends trainable tokens to the input like prompt tuning but injects separate prefixes in every layer. For each Transformer layer, prefix-tuning prepends trainable embeddings to the attention’s ${\bm{K}}$ and ${\bm{V}}$ matrix. The authors have found that directly training these prefixes can lead to unstable training, so they propose to over-parameterize them with a large MLP to increase training stability. After training, the MLP can be dropped. Like prompt tuning, prefix-tuning introduces an inference overhead, scaling linearly with the number of trainable embeddings.
BitFit (Zaken et al., 2022).
BitFit is a simple but effective PEFT method that freezes all model weights except the bias terms, consequently greatly reducing the number of trainable parameters. As no additional parameters are added, no inference overhead occurs.
Appendix B Details of Datasets
In this paper, we consider five datasets across three domains: (i) Natural Language Understanding (NLU), represented by GLUE (Wang et al., 2019); (ii) Natural Language Generation (NLG), including SAMSum (Gliwa et al., 2019), Spider (Yu et al., 2018) and DART (Nan et al., 2021); and (iii) Computer Vision (CV), represented by CIFAR-10 (Krizhevsky et al., 2009).
GLUE (Wang et al., 2019).
The GLUE (General Language Understanding Evaluation) benchmark is a collection of datasets used for training, evaluating, and analyzing natural language understanding models across a range of diverse tasks. The benchmark includes nine sentence- or sentence-pair language understanding tasks that require various features of understanding, such as sentiment analysis, linguistic acceptability, semantic textual similarity, and question answering. We use seven datasets from the GLUE benchmark (RTE, MRPC, CoLA, SST-2, QNLI, QQP, MNLI) where the model has to choose between two or three (for MNLI) different choices for the respective task. Except for CoLA, we evaluate all used datasets with the accuracy metric. For CoLA, Matthews correlation is employed.
SAMSum (Gliwa et al., 2019).
SAMSum is a dataset for dialogue summarization research, comprising approximately 16,000 synthetic text conversations with accompanying summaries. Created by English-fluent linguists, these exchanges simulate real-world digital communications across various topics and styles. The conversations range from informal to formal, incorporating elements like slang and emoticons to reflect authentic messaging patterns. Each dialogue is paired with a concise, third-person summary, capturing its essential content. This structure makes SAMSum particularly useful for developing and evaluating automated summarization systems capable of processing conversational text.
Spider (Yu et al., 2018).
Spider is a large-scale, complex, and cross-domain semantic parsing and text-to-SQL dataset. It contains about 10,000 annotated SQL queries, distributed across 200+ databases, each with multiple tables. We follow Scholak et al. (2021) and use about 7,000 examples for training and about 1,000 examples for validation, where we ignore sequences longer than 1536 tokens. The dataset consists of English question and SQL query pairs, which cover a wide range of SQL operations including SELECT, WHERE, COUNT, GROUP BY, ORDER BY, JOIN, and more. Given an English question and an SQL database scheme, the task for the model is to translate the English question into an appropriate SQL statement. Evaluation is performed via accuracy where the output is considered as correct if the model’s predicted SQL query and the included GT SQL query give the same result when executed on the database. The dataset additionally categorizes each query into easy (25%), medium (40%), hard (20%), and extra hard (15%) based on the complexity of the required SQL statement. For evaluation, we report the execution accuracy of all categories.
DART (Nan et al., 2021).
The DART (DAta Record to Text) benchmark is a large-scale, structured dataset designed for RDF-to-text (Resource Description Framework-to-text) generation with 80,000+ instances. The DART benchmark is composed of a collection of structured data triples and corresponding text summaries which are organized into different categories. The task of the DART benchmark is to generate natural language summaries that correctly represent the given structured data inputs. DART is typically evaluated with METEOR and BLEU.
CIFAR-10 (Krizhevsky et al., 2009).
The CIFAR-10 (Canadian Institute For Advanced Research) dataset is a collection of images that are commonly used to train machine learning and computer vision algorithms. It is one of the most widely used datasets for image classification. The CIFAR-10 dataset contains 60,000 (50,000 for training, 10,000 for validation) 32$\times$32 color images in 10 different classes. The 10 different classes are: airplane, car, bird, cat, deer, dog, frog, horse, ship, and truck. There are 6,000 images of each class. For training, we center crop each image to 24$\times$24 pixels and flatten each image to a string, with a total of 24$\times$24$\times$3 words, where each word is a number between 0-255 representing the respective pixel value. Although CIFAR-10 is a dataset for computer vision, previous work (Dinh et al., 2022) showed that Transformers can be adapted to the vision domain from the language domain, and we tested this ability on the state-space model.
Data Size (Train) Size (Val) Size (Test) Max. seq. len. #Epochs Mamba Size Metrics
GLUE RTE 1992 498 277 291 10 130M Accuracy
MRPC 2934 734 408 105 10 130M Accuracy
CoLA 6840 1711 1043 47 10 130M Matthews corr.
SST-2 53879 13470 872 68 10 130M Accuracy
QNLI 83794 20949 5463 602 10 130M Accuracy
QQP 291076 72770 40430 316 3 130M Accuracy
MNLI 314161 78541 19647 425 3 130M Accuracy
Spider 5543 1375 1034 1412 10 1.4B, 2.8B Accuracy
SAMSum 14732 818 819 1174 10 1.4B ROUGE
DART 62659 2768 5097 491 10 130M METEOR, BLEU
CIFAR-10 40000 10000 10000 1730 5 130M Accuracy
Table 6: Datasets and models for our experiments. For each dataset, we report the number of training, validation, and test samples, maximum sequence length, training epochs, model size, and evaluation metric used.
The dataset characteristics, including our train, validation and test set sizes, sequence lengths, and number of epochs, are summarized in Table 6.
Appendix C Details of Sec. 4: Benchmarking PEFT Methods on SSM-based Models
In this section, we provide a comprehensive experimental setup, proofs and further discussion of theoretical results, and more detailed experimental outcomes.
C.1 Experiment Setup
[…]
C.2 Extended Results on Benchmarking Existing PEFT Methods
We present comprehensive fine-tuning results for the GLUE benchmark (Wang et al., 2019), DART dataset (Nan et al., 2021), SAMSum dataset (Gliwa et al., 2019) and Spider dataset (Yu et al., 2018) in Table 7, Table 8, Table 9 and Table 10, respectively. These experimental results encompass various LoRA implementations (on different weight matrices and modules) and provide more fine-grained results across all subtasks.
Layer Method # Params (%) RTE MRPC CoLA SST-2 QNLI QQP MNLI Avg.
Pretrained $0.0$ 0.469 3140.4693140.469\,3140.469 314 0.678 9220.6789220.678\,9220.678 922 0.000 $00\,0000$ 0.524 0830.5240830.524\,0830.524 083 0.505 2170.5052170.505\,2170.505 217 0.368 1180.3681180.368\,1180.368 118 0.322 5940.3225940.322\,5940.322 594 0.409 7500.4097500.409\,7500.409 750
All All Full $100.0$ 0.711 191 356 182 098 40.71119135618209840.711\,191\,356\,182\,098\,40.711 191 356 182 098 4 0.806 372 523 307 800 30.80637252330780030.806\,372\,523\,307\,800\,30.806 372 523 307 800 3 0.631 936 013 698 577 90.63193601369857790.631\,936\,013\,698\,577\,90.631 936 013 698 577 9 0.922 018 349 170 684 80.92201834917068480.922\,018\,349\,170\,684\,80.922 018 349 170 684 8 0.874 061 882 495 880 10.87406188249588010.874\,061\,882\,495\,880\,10.874 061 882 495 880 1 0.878 703 951 835 632 30.87870395183563230.878\,703\,951\,835\,632\,30.878 703 951 835 632 3 0.807 655 096 054 077 10.80765509605407710.807\,655\,096\,054\,077\,10.807 655 096 054 077 1 0.804 562 738 963 535 80.80456273896353580.804\,562\,738\,963\,535\,80.804 562 738 963 535 8
LoRA 1.922 527 283 438 617 31.92252728343861731.922\,527\,283\,438\,617\,31.922 527 283 438 617 3 0.699 275 374 412 536 60.69927537441253660.699\,275\,374\,412\,536\,60.699 275 374 412 536 6 0.808 823 525 905 609 10.80882352590560910.808\,823\,525\,905\,609\,10.808 823 525 905 609 1 0.614 111 959 934 234 60.61411195993423460.614\,111\,959\,934\,234\,60.614 111 959 934 234 6 0.918 577 969 074 249 30.91857796907424930.918\,577\,969\,074\,249\,30.918 577 969 074 249 3 0.884 312 629 699 7070.8843126296997070.884\,312\,629\,699\,7070.884 312 629 699 707 0.876 131 594 181 060 80.87613159418106080.876\,131\,594\,181\,060\,80.876 131 594 181 060 8 0.811 370 670 795 440 70.81137067079544070.811\,370\,670\,795\,440\,70.811 370 670 795 440 7 0.801 800 532 000 405 50.80180053200040550.801\,800\,532\,000\,405\,50.801 800 532 000 405 5
Prompt Prompt Tuning 16 tokens 0.009 514 691 278 001 4380.0095146912780014380.009\,514\,691\,278\,001\,4380.009 514 691 278 001 438 0.559 566 795 825 958 30.55956679582595830.559\,566\,795\,825\,958\,30.559 566 795 825 958 3 0.715 686 261 653 900 10.71568626165390010.715\,686\,261\,653\,900\,10.715 686 261 653 900 1 0.119 891 807 436 943 050.119891807436943050.119\,891\,807\,436\,943\,050.119 891 807 436 943 05 0.894 495 427 608 490.894495427608490.894\,495\,427\,608\,490.894 495 427 608 49 0.767 893 075 942 993 20.76789307594299320.767\,893\,075\,942\,993\,20.767 893 075 942 993 2 0.795 819 938 182 830 80.79581993818283080.795\,819\,938\,182\,830\,80.795 819 938 182 830 8 0.614 863 812 923 431 40.61486381292343140.614\,863\,812\,923\,431\,40.614 863 812 923 431 4 0.638 316 731 367 792 40.63831673136779240.638\,316\,731\,367\,792\,40.638 316 731 367 792 4
Prefix-Tuning 1 token (no MLP) 0.028 538 643 106 431 3040.0285386431064313040.028\,538\,643\,106\,431\,3040.028 538 643 106 431 304 0.675 090 253 353 118 90.67509025335311890.675\,090\,253\,353\,118\,90.675 090 253 353 118 9 0.757 352 948 188 781 70.75735294818878170.757\,352\,948\,188\,781\,70.757 352 948 188 781 7 0.434 039 652 347 564 70.43403965234756470.434\,039\,652\,347\,564\,70.434 039 652 347 564 7 0.915 137 588 977 813 70.91513758897781370.915\,137\,588\,977\,813\,70.915 137 588 977 813 7 0.834 157 049 655 914 30.83415704965591430.834\,157\,049\,655\,914\,30.834 157 049 655 914 3 0.831 066 012 382 507 30.83106601238250730.831\,066\,012\,382\,507\,30.831 066 012 382 507 3 0.356 390 297 412 872 30.35639029741287230.356\,390\,297\,412\,872\,30.356 390 297 412 872 3 0.686 176 257 474 081 80.68617625747408180.686\,176\,257\,474\,081\,80.686 176 257 474 081 8
[…]
linear projection matrices All LoRA 1.017 230 519 270 278 41.01723051927027841.017\,230\,519\,270\,278\,41.017 230 519 270 278 4 0.700 361 013 412 475 60.70036101341247560.700\,361\,013\,412\,475\,60.700 361 013 412 475 6 0.823 529 422 283 172 60.82352942228317260.823\,529\,422\,283\,172\,60.823 529 422 283 172 6 0.577 130 556 106 567 40.57713055610656740.577\,130\,556\,106\,567\,40.577 130 556 106 567 4 0.933 486 223 220 825 20.93348622322082520.933\,486\,223\,220\,825\,20.933 486 223 220 825 2 0.887 241 423 130 035 40.88724142313003540.887\,241\,423\,130\,035\,40.887 241 423 130 035 4 0.887 261 927 127 838 10.88726192712783810.887\,261\,927\,127\,838\,10.887 261 927 127 838 1 0.825 316 846 370 6970.8253168463706970.825\,316\,846\,370\,6970.825 316 846 370 697 0.804 903 915 950 230 10.80490391595023010.804\,903\,915\,950\,230\,10.804 903 915 950 230 1
$\bm{W}_{\text{in},x}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 0.703 971 147 537 231 40.70397114753723140.703\,971\,147\,537\,231\,40.703 971 147 537 231 4 0.821 078 419 685 363 80.82107841968536380.821\,078\,419\,685\,363\,80.821 078 419 685 363 8 0.573 514 342 308 044 40.57351434230804440.573\,514\,342\,308\,044\,40.573 514 342 308 044 4 0.917 431 175 708 770 80.91743117570877080.917\,431\,175\,708\,770\,80.917 431 175 708 770 8 0.883 031 308 650 970 50.88303130865097050.883\,031\,308\,650\,970\,50.883 031 308 650 970 5 0.877 145 707 607 269 30.87714570760726930.877\,145\,707\,607\,269\,30.877 145 707 607 269 3 0.812 185 049 057 006 80.81218504905700680.812\,185\,049\,057\,006\,80.812 185 049 057 006 8 0.798 336 735 793 522 40.79833673579352240.798\,336\,735\,793\,522\,40.798 336 735 793 522 4
$\bm{W}_{\text{in},z}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 0.700 361 013 412 475 60.70036101341247560.700\,361\,013\,412\,475\,60.700 361 013 412 475 6 0.823 529 422 283 172 60.82352942228317260.823\,529\,422\,283\,172\,60.823 529 422 283 172 6 0.581 120 252 609 252 90.58112025260925290.581\,120\,252\,609\,252\,90.581 120 252 609 252 9 0.924 311 935 901 641 80.92431193590164180.924\,311\,935\,901\,641\,80.924 311 935 901 641 8 0.873 146 593 570 709 20.87314659357070920.873\,146\,593\,570\,709\,20.873 146 593 570 709 2 0.873 212 933 540 344 20.87321293354034420.873\,212\,933\,540\,344\,20.873 212 933 540 344 2 0.803 990 423 679 351 80.80399042367935180.803\,990\,423\,679\,351\,80.803 990 423 679 351 8 0.797 096 082 142 421 20.79709608214242120.797\,096\,082\,142\,421\,20.797 096 082 142 421 2
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.680 460 965 049 587 40.68046096504958740.680\,460\,965\,049\,587\,40.680 460 965 049 587 4 0.703 971 147 537 231 40.70397114753723140.703\,971\,147\,537\,231\,40.703 971 147 537 231 4 0.843 137 264 251 7090.8431372642517090.843\,137\,264\,251\,7090.843 137 264 251 709 0.624 137 103 557 586 70.62413710355758670.624\,137\,103\,557\,586\,70.624 137 103 557 586 7 0.925 458 729 267 120 40.92545872926712040.925\,458\,729\,267\,120\,40.925 458 729 267 120 4 0.886 326 193 809 509 30.88632619380950930.886\,326\,193\,809\,509\,30.886 326 193 809 509 3 0.883 353 948 593 139 60.88335394859313960.883\,353\,948\,593\,139\,60.883 353 948 593 139 6 0.817 427 575 588 226 30.81742757558822630.817\,427\,575\,588\,226\,30.817 427 575 588 226 3 0.811 973 137 514 931 80.81197313751493180.811\,973\,137\,514\,931\,80.811 973 137 514 931 8
$\bm{W}_{\text{out}}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 0.703 971 147 537 231 40.70397114753723140.703\,971\,147\,537\,231\,40.703 971 147 537 231 4 0.828 431 367 874 145 50.82843136787414550.828\,431\,367\,874\,145\,50.828 431 367 874 145 5 0.605 746 030 807 495 10.60574603080749510.605\,746\,030\,807\,495\,10.605 746 030 807 495 1 0.924 311 935 901 641 80.92431193590164180.924\,311\,935\,901\,641\,80.924 311 935 901 641 8 0.884 312 629 699 7070.8843126296997070.884\,312\,629\,699\,7070.884 312 629 699 707 0.876 675 724 983 215 30.87667572498321530.876\,675\,724\,983\,215\,30.876 675 724 983 215 3 0.814 831 793 308 258 10.81483179330825810.814\,831\,793\,308\,258\,10.814 831 793 308 258 1 0.805 468 661 444 527 80.80546866144452780.805\,468\,661\,444\,527\,80.805 468 661 444 527 8
S6 All Full 4.310 565 285 913 9444.3105652859139444.310\,565\,285\,913\,9444.310 565 285 913 944 0.696 750 879 287 719 70.69675087928771970.696\,750\,879\,287\,719\,70.696 750 879 287 719 7 0.789 215 683 937 072 80.78921568393707280.789\,215\,683\,937\,072\,80.789 215 683 937 072 8 0.590 752 780 437 469 50.59075278043746950.590\,752\,780\,437\,469\,50.590 752 780 437 469 5 0.915 137 588 977 813 70.91513758897781370.915\,137\,588\,977\,813\,70.915 137 588 977 813 7 0.880 651 652 812 957 80.88065165281295780.880\,651\,652\,812\,957\,80.880 651 652 812 957 8 0.875 290 632 247 924 80.87529063224792480.875\,290\,632\,247\,924\,80.875 290 632 247 924 8 0.804 906 606 674 194 30.80490660667419430.804\,906\,606\,674\,194\,30.804 906 606 674 194 3 0.793 243 689 196 450 40.79324368919645040.793\,243\,689\,196\,450\,40.793 243 689 196 450 4
LoRA 0.923 912 723 244 536 30.92391272324453630.923\,912\,723\,244\,536\,30.923 912 723 244 536 3 0.660 649 836 063 3850.6606498360633850.660\,649\,836\,063\,3850.660 649 836 063 385 0.786 764 681 339 263 90.78676468133926390.786\,764\,681\,339\,263\,90.786 764 681 339 263 9 0.578 182 518 482 208 30.57818251848220830.578\,182\,518\,482\,208\,30.578 182 518 482 208 3 0.908 256 888 389 587 40.90825688838958740.908\,256\,888\,389\,587\,40.908 256 888 389 587 4 0.877 722 859 382 629 40.87772285938262940.877\,722\,859\,382\,629\,40.877 722 859 382 629 4 0.868 859 767 913 818 40.86885976791381840.868\,859\,767\,913\,818\,40.868 859 767 913 818 4 0.798 137 128 353 118 90.79813712835311890.798\,137\,128\,353\,118\,90.798 137 128 353 118 9 0.782 653 382 846 287 30.78265338284628730.782\,653\,382\,846\,287\,30.782 653 382 846 287 3
${\bm{A}}$ Full 0.456 748 639 567 040 370.456748639567040370.456\,748\,639\,567\,040\,370.456 748 639 567 040 37 0.682 310 461 997 985 80.68231046199798580.682\,310\,461\,997\,985\,80.682 310 461 997 985 8 0.821 078 419 685 363 80.82107841968536380.821\,078\,419\,685\,363\,80.821 078 419 685 363 8 0.542 106 330 394 744 90.54210633039474490.542\,106\,330\,394\,744\,90.542 106 330 394 744 9 0.909 403 681 755 065 90.90940368175506590.909\,403\,681\,755\,065\,90.909 403 681 755 065 9 0.863 811 075 687 408 40.86381107568740840.863\,811\,075\,687\,408\,40.863 811 075 687 408 4 0.878 898 799 419 403 10.87889879941940310.878\,898\,799\,419\,403\,10.878 898 799 419 403 1 0.793 505 370 616 912 80.79350537061691280.793\,505\,370\,616\,912\,80.793 505 370 616 912 8 0.784 444 877 079 5550.7844448770795550.784\,444\,877\,079\,5550.784 444 877 079 555
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.283 743 197 835 2022.2837431978352022.283\,743\,197\,835\,2022.283 743 197 835 202 0.696 750 879 287 719 70.69675087928771970.696\,750\,879\,287\,719\,70.696 750 879 287 719 7 0.769 607 841 968 536 40.76960784196853640.769\,607\,841\,968\,536\,40.769 607 841 968 536 4 0.558 246 970 176 696 80.55824697017669680.558\,246\,970\,176\,696\,80.558 246 970 176 696 8 0.913 990 855 216 980.913990855216980.913\,990\,855\,216\,980.913 990 855 216 98 0.854 475 557 804 107 70.85447555780410770.854\,475\,557\,804\,107\,70.854 475 557 804 107 7 0.850 157 082 080 841 10.85015708208084110.850\,157\,082\,080\,841\,10.850 157 082 080 841 1 0.767 954 409 122 4670.7679544091224670.767\,954\,409\,122\,4670.767 954 409 122 467 0.773 026 227 951 049 80.77302622795104980.773\,026\,227\,951\,049\,80.773 026 227 951 049 8
LoRA 0.692 192 750 875 576 50.69219275087557650.692\,192\,750\,875\,576\,50.692 192 750 875 576 5 0.678 700 387 477 874 80.67870038747787480.678\,700\,387\,477\,874\,80.678 700 387 477 874 8 0.789 215 683 937 072 80.78921568393707280.789\,215\,683\,937\,072\,80.789 215 683 937 072 8 0.488 315 135 240 554 80.48831513524055480.488\,315\,135\,240\,554\,80.488 315 135 240 554 8 0.913 990 855 216 980.913990855216980.913\,990\,855\,216\,980.913 990 855 216 98 0.868 753 433 227 539 10.86875343322753910.868\,753\,433\,227\,539\,10.868 753 433 227 539 1 0.858 443 200 588 226 30.85844320058822630.858\,443\,200\,588\,226\,30.858 443 200 588 226 3 0.786 114 215 850 830 10.78611421585083010.786\,114\,215\,850\,830\,10.786 114 215 850 830 1 0.769 076 130 219 868 20.76907613021986820.769\,076\,130\,219\,868\,20.769 076 130 219 868 2
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.398 792 708 674 0611.3987927086740611.398\,792\,708\,674\,0611.398 792 708 674 061 0.660 649 836 063 3850.6606498360633850.660\,649\,836\,063\,3850.660 649 836 063 385 0.752 451 002 597 808 80.75245100259780880.752\,451\,002\,597\,808\,80.752 451 002 597 808 8 0.566 903 471 946 716 30.56690347194671630.566\,903\,471\,946\,716\,30.566 903 471 946 716 3 0.910 550 475 120 544 40.91055047512054440.910\,550\,475\,120\,544\,40.910 550 475 120 544 4 0.862 479 388 713 836 70.86247938871383670.862\,479\,388\,713\,836\,70.862 479 388 713 836 7 0.871 308 445 930 4810.8713084459304810.871\,308\,445\,930\,4810.871 308 445 930 481 0.784 801 721 572 8760.7848017215728760.784\,801\,721\,572\,8760.784 801 721 572 876 0.772 734 905 992 235 50.77273490599223550.772\,734\,905\,992\,235\,50.772 734 905 992 235 5
LoRA 0.234 957 666 087 183 550.234957666087183550.234\,957\,666\,087\,183\,550.234 957 666 087 183 55 0.671 480 119 228 3630.6714801192283630.671\,480\,119\,228\,3630.671 480 119 228 363 0.799 019 634 723 663 30.79901963472366330.799\,019\,634\,723\,663\,30.799 019 634 723 663 3 0.550 651 848 316 192 60.55065184831619260.550\,651\,848\,316\,192\,60.550 651 848 316 192 6 0.909 403 681 755 065 90.90940368175506590.909\,403\,681\,755\,065\,90.909 403 681 755 065 9 0.527 375 936 508 178 70.52737593650817870.527\,375\,936\,508\,178\,70.527 375 936 508 178 7 0.865 866 959 095 001 20.86586695909500120.865\,866\,959\,095\,001\,20.865 866 959 095 001 2 0.786 634 087 562 5610.7866340875625610.786\,634\,087\,562\,5610.786 634 087 562 561 0.730 061 752 455 575 10.73006175245557510.730\,061\,752\,455\,575\,10.730 061 752 455 575 1
Conv1d Full 0.142 733 949 864 700 120.142733949864700120.142\,733\,949\,864\,700\,120.142 733 949 864 700 12 0.682 310 461 997 985 80.68231046199798580.682\,310\,461\,997\,985\,80.682 310 461 997 985 8 0.784 313 738 346 099 90.78431373834609990.784\,313\,738\,346\,099\,90.784 313 738 346 099 9 0.578 959 524 631 500 20.57895952463150020.578\,959\,524\,631\,500\,20.578 959 524 631 500 2 0.910 550 475 120 544 40.91055047512054440.910\,550\,475\,120\,544\,40.910 550 475 120 544 4 0.859 600 961 208 343 50.85960096120834350.859\,600\,961\,208\,343\,50.859 600 961 208 343 5 0.859 980 225 563 049 30.85998022556304930.859\,980\,225\,563\,049\,30.859 980 225 563 049 3 0.779 508 292 675 018 30.77950829267501830.779\,508\,292\,675\,018\,30.779 508 292 675 018 3 0.779 317 668 506 077 30.77931766850607730.779\,317\,668\,506\,077\,30.779 317 668 506 077 3
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.043 414 909 750 512 950.043414909750512950.043\,414\,909\,750\,512\,950.043 414 909 750 512 95 0.653 429 627 418 518 10.65342962741851810.653\,429\,627\,418\,518\,10.653 429 627 418 518 1 0.791 666 686 534 881 60.79166668653488160.791\,666\,686\,534\,881\,60.791 666 686 534 881 6 0.403 467 684 984 207 150.403467684984207150.403\,467\,684\,984\,207\,150.403 467 684 984 207 15 0.910 550 475 120 544 40.91055047512054440.910\,550\,475\,120\,544\,40.910 550 475 120 544 4 0.839 069 962 501 525 90.83906996250152590.839\,069\,962\,501\,525\,90.839 069 962 501 525 9 0.859 824 895 858 764 60.85982489585876460.859\,824\,895\,858\,764\,60.859 824 895 858 764 6 0.670 331 358 909 606 90.67033135890960690.670\,331\,358\,909\,606\,90.670 331 358 909 606 9 0.732 620 098 761 149 90.73262009876114990.732\,620\,098\,761\,149\,90.732 620 098 761 149 9
Table 7: Full experimental results on the GLUE (Wang et al., 2019) benchmark. We report accuracy ($\uparrow$) for RTE, MRPC, SST-2, QNLI, QQP, and MNLI tasks. CoLA performance is measured using Matthews Correlation Coefficient ($\uparrow$). Mamba-130M is employed in this experiment. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) METEOR BLEU
All All Full $100.0$ 71.008 300 781 2571.0083007812571.008\,300\,781\,2571.008 300 781 25 51.802 164 316 177 3751.8021643161773751.802\,164\,316\,177\,3751.802 164 316 177 37
LoRA 1.922 527 283 438 617 31.92252728343861731.922\,527\,283\,438\,617\,31.922 527 283 438 617 3 70.968 002 080 917 3670.9680020809173670.968\,002\,080\,917\,3670.968 002 080 917 36 49.521 303 176 879 8849.5213031768798849.521\,303\,176\,879\,8849.521 303 176 879 88
Prompt Prompt Tuning 64 tokens 0.038 047 904 689 998 750.038047904689998750.038\,047\,904\,689\,998\,750.038 047 904 689 998 75 66.186 875 104 904 1766.1868751049041766.186\,875\,104\,904\,1766.186 875 104 904 17 39.826 393 127 441 40639.82639312744140639.826\,393\,127\,441\,40639.826 393 127 441 406
Prefix-Tuning 64 tokens 22.688 043 993 029 53522.68804399302953522.688\,043\,993\,029\,53522.688 043 993 029 535 66.589 879 989 624 0266.5898799896240266.589\,879\,989\,624\,0266.589 879 989 624 02 42.462 074 756 622 31442.46207475662231442.462\,074\,756\,622\,31442.462 074 756 622 314
[…]
linear projection matrices All LoRA 1.017 230 519 270 278 41.01723051927027841.017\,230\,519\,270\,278\,41.017 230 519 270 278 4 71.176 528 930 664 0671.1765289306640671.176\,528\,930\,664\,0671.176 528 930 664 06 49.160 829 186 439 51449.16082918643951449.160\,829\,186\,439\,51449.160 829 186 439 514
$\bm{W}_{\text{in},x}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 70.250 225 067 138 6770.2502250671386770.250\,225\,067\,138\,6770.250 225 067 138 67 48.860 117 793 083 1948.8601177930831948.860\,117\,793\,083\,1948.860 117 793 083 19
$\bm{W}_{\text{in},z}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 70.425 111 055 374 1570.4251110553741570.425\,111\,055\,374\,1570.425 111 055 374 15 49.059 996 008 872 98649.05999600887298649.059\,996\,008\,872\,98649.059 996 008 872 986
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.680 460 965 049 587 40.68046096504958740.680\,460\,965\,049\,587\,40.680 460 965 049 587 4 70.937 132 835 388 1870.9371328353881870.937\,132\,835\,388\,1870.937 132 835 388 18 49.452 367 424 964 90549.45236742496490549.452\,367\,424\,964\,90549.452 367 424 964 905
$\bm{W}_{\text{out}}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 70.727 145 671 844 4870.7271456718444870.727\,145\,671\,844\,4870.727 145 671 844 48 46.977 826 952 934 26546.97782695293426546.977\,826\,952\,934\,26546.977 826 952 934 265
S6 All Full 4.310 565 285 913 9444.3105652859139444.310\,565\,285\,913\,9444.310 565 285 913 944 70.348 870 754 241 9470.3488707542419470.348\,870\,754\,241\,9470.348 870 754 241 94 48.673 865 199 089 0548.6738651990890548.673\,865\,199\,089\,0548.673 865 199 089 05
LoRA 0.923 912 723 244 536 30.92391272324453630.923\,912\,723\,244\,536\,30.923 912 723 244 536 3 69.898 468 255 996 769.898468255996769.898\,468\,255\,996\,769.898 468 255 996 7 50.779 956 579 208 37450.77995657920837450.779\,956\,579\,208\,37450.779 956 579 208 374
${\bm{A}}$ Full 0.456 748 639 567 040 370.456748639567040370.456\,748\,639\,567\,040\,370.456 748 639 567 040 37 69.333 219 528 198 2469.3332195281982469.333\,219\,528\,198\,2469.333 219 528 198 24 48.095 324 635 505 67648.09532463550567648.095\,324\,635\,505\,67648.095 324 635 505 676
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.283 743 197 835 2022.2837431978352022.283\,743\,197\,835\,2022.283 743 197 835 202 70.056 617 259 979 2570.0566172599792570.056\,617\,259\,979\,2570.056 617 259 979 25 49.983 921 647 071 8449.9839216470718449.983\,921\,647\,071\,8449.983 921 647 071 84
LoRA 0.692 192 750 875 576 50.69219275087557650.692\,192\,750\,875\,576\,50.692 192 750 875 576 5 68.773 013 353 347 7868.7730133533477868.773\,013\,353\,347\,7868.773 013 353 347 78 47.991 544 008 255 00547.99154400825500547.991\,544\,008\,255\,00547.991 544 008 255 005
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.398 792 708 674 0611.3987927086740611.398\,792\,708\,674\,0611.398 792 708 674 061 69.576 507 806 777 9569.5765078067779569.576\,507\,806\,777\,9569.576 507 806 777 95 47.242 343 425 750 7347.2423434257507347.242\,343\,425\,750\,7347.242 343 425 750 73
LoRA 0.234 957 666 087 183 550.234957666087183550.234\,957\,666\,087\,183\,550.234 957 666 087 183 55 68.856 698 274 612 4368.8566982746124368.856\,698\,274\,612\,4368.856 698 274 612 43 47.047 722 339 630 1347.0477223396301347.047\,722\,339\,630\,1347.047 722 339 630 13
Conv1d Full 0.142 733 949 864 700 120.142733949864700120.142\,733\,949\,864\,700\,120.142 733 949 864 700 12 68.621 623 516 082 7668.6216235160827668.621\,623\,516\,082\,7668.621 623 516 082 76 47.934 231 162 071 2347.9342311620712347.934\,231\,162\,071\,2347.934 231 162 071 23
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.043 414 909 750 512 950.043414909750512950.043\,414\,909\,750\,512\,950.043 414 909 750 512 95 67.027 038 335 800 1767.0270383358001767.027\,038\,335\,800\,1767.027 038 335 800 17 44.228 667 020 797 7344.2286670207977344.228\,667\,020\,797\,7344.228 667 020 797 73
Table 8: Full experimental results on the DART (Nan et al., 2021) benchmark. We report METEOR ()($\uparrow$)( ↑ ) and BLEU ()($\uparrow$)( ↑ ) scores. Mamba-130M is utilized in this experiment. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) R1 R2 RL
All All Full $100.0$ $51.2$ $27.3$ $42.9$
LoRA 0.972 818 142 718 422 30.97281814271842230.972\,818\,142\,718\,422\,30.972 818 142 718 422 3 50.844 812 3950.8448123950.844\,812\,3950.844 812 39 26.647 731 6626.6477316626.647\,731\,6626.647 731 66 42.692 086 142.692086142.692\,086\,142.692 086 1
Prompt Prompt Tuning 64 tokens 0.009 551 198 1530.0095511981530.009\,551\,198\,1530.009 551 198 153 $50.1$ $25.6$ $41.6$
Prefix-Tuning 64 tokens 12.807 268 9112.8072689112.807\,268\,9112.807 268 91 $50.6$ $26.5$ $42.1$
[…]
linear projection matrices All LoRA 0.513 166 979 724 261 80.51316697972426180.513\,166\,979\,724\,261\,80.513 166 979 724 261 8 50.820 308 9250.8203089250.820\,308\,9250.820 308 92 26.873 826 9826.8738269826.873\,826\,9826.873 826 98 42.773 035 1742.7730351742.773\,035\,1742.773 035 17
$\bm{W}_{\text{in},x}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 49.831 70349.83170349.831\,70349.831 703 25.436 69325.43669325.436\,69325.436 693 41.155 55341.15555341.155\,55341.155 553
$\bm{W}_{\text{in},z}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 50.021 547 0850.0215470850.021\,547\,0850.021 547 08 26.050 487 1626.0504871626.050\,487\,1626.050 487 16 41.673 398 0241.6733980241.673\,398\,0241.673 398 02
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.342 697 523 326 188 860.342697523326188860.342\,697\,523\,326\,188\,860.342 697 523 326 188 86 50.872 570 2850.8725702850.872\,570\,2850.872 570 28 26.967 412 2326.9674122326.967\,412\,2326.967 412 23 42.283 281 6842.2832816842.283\,281\,6842.283 281 68
$\bm{W}_{\text{out}}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 49.860 50149.86050149.860\,50149.860 501 25.443 37825.44337825.443\,37825.443 378 41.459 15541.45915541.459\,15541.459 155
S6 All Full 4.456 059 545 468 7924.4560595454687924.456\,059\,545\,468\,7924.456 059 545 468 792 51.134 622 151.134622151.134\,622\,151.134 622 1 26.891 985 5426.8919855426.891\,985\,5426.891 985 54 42.242 348 1942.2423481942.242\,348\,1942.242 348 19
LoRA 0.464 394 215 127 723 30.46439421512772330.464\,394\,215\,127\,723\,30.464 394 215 127 723 3 50.520 539 2850.5205392850.520\,539\,2850.520 539 28 26.363 494 9926.3634949926.363\,494\,9926.363 494 99 42.184 033 9942.1840339942.184\,033\,9942.184 033 99
${\bm{A}}$ Full 0.229 250 651 856 915 340.229250651856915340.229\,250\,651\,856\,915\,340.229 250 651 856 915 34 50.090 897 0850.0908970850.090\,897\,0850.090 897 08 25.944 900 5125.9449005125.944\,900\,5125.944 900 51 41.724 771 2641.7247712641.724\,771\,2641.724 771 26
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.292 506 518 569 153 82.29250651856915382.292\,506\,518\,569\,153\,82.292 506 518 569 153 8 50.457 620 6250.4576206250.457\,620\,6250.457 620 62 25.973 382 5925.9733825925.973\,382\,5925.973 382 59 41.786 766 0541.7867660541.786\,766\,0541.786 766 05
LoRA 0.347 144 240 365 408 70.34714424036540870.347\,144\,240\,365\,408\,70.347 144 240 365 408 7 50.416 523 2250.4165232250.416\,523\,2250.416 523 22 25.998 947 0225.9989470225.998\,947\,0225.998 947 02 41.848 915 8241.8489158241.848\,915\,8241.848 915 82
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.848 333 380 596 381.848333380596381.848\,333\,380\,596\,381.848 333 380 596 38 50.262 951 8550.2629518550.262\,951\,8550.262 951 85 25.663 086 7725.6630867725.663\,086\,7725.663 086 77 41.611 900 9341.6119009341.611\,900\,9341.611 900 93
LoRA 0.118 067 802 522 655 750.118067802522655750.118\,067\,802\,522\,655\,750.118 067 802 522 655 75 50.183 439 2550.1834392550.183\,439\,2550.183 439 25 25.424 981 1225.4249811225.424\,981\,1225.424 981 12 41.260 328 8941.2603288941.260\,328\,8941.260 328 89
Conv1d Full 0.071 640 828 705 286 060.071640828705286060.071\,640\,828\,705\,286\,060.071 640 828 705 286 06 50.085 21750.08521750.085\,21750.085 217 25.710 83025.71083025.710\,83025.710 830 41.920 88541.92088541.920\,88541.920 885
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.021 641 500 338 055 160.021641500338055160.021\,641\,500\,338\,055\,160.021 641 500 338 055 16 49.580 90849.58090849.580\,90849.580 908 24.796 23024.79623024.796\,23024.796 230 41.105 80741.10580741.105\,80741.105 807
Table 9: Full experimental results on the SAMSum (Gliwa et al., 2019) benchmark. R1, R2, and RL represent ROUGE-1 ($\uparrow$), ROUGE-2 ($\uparrow$), and ROUGE-L ($\uparrow$), respectively. Mamba-1.4B is utilized in this experiment. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) All Easy Medium Hard Extra
All All Full $100.0$ 66.150 867 938 995 3666.1508679389953666.150\,867\,938\,995\,3666.150 867 938 995 36 84.274 190 664 291 3884.2741906642913884.274\,190\,664\,291\,3884.274 190 664 291 38 69.506 728 649 139 469.506728649139469.506\,728\,649\,139\,469.506 728 649 139 4 53.448 277 711 868 28653.44827771186828653.448\,277\,711\,868\,28653.448 277 711 868 286 43.373 495 340 347 2943.3734953403472943.373\,495\,340\,347\,2943.373 495 340 347 29
LoRA 0.972 818 142 718 422 30.97281814271842230.972\,818\,142\,718\,422\,30.972 818 142 718 422 3 56.382 977 962 493 956.382977962493956.382\,977\,962\,493\,956.382 977 962 493 9 76.209 676 265 716 5576.2096762657165576.209\,676\,265\,716\,5576.209 676 265 716 55 56.950 670 480 728 1556.9506704807281556.950\,670\,480\,728\,1556.950 670 480 728 15 47.701 150 178 909 347.701150178909347.701\,150\,178\,909\,347.701 150 178 909 3 34.337 350 726 127 62534.33735072612762534.337\,350\,726\,127\,62534.337 350 726 127 625
Prompt Prompt Tuning 64 tokens 0.009 551 198 1530.0095511981530.009\,551\,198\,1530.009 551 198 153 43.617 022 0443.6170220443.617\,022\,0443.617 022 04 65.322 577 9565.3225779565.322\,577\,9565.322 577 95 42.376 682 1642.3766821642.376\,682\,1642.376 682 16 33.333 334 $3$.3333343333.333\,334\,$3$.333 334 33 25.301 206 1125.3012061125.301\,206\,1125.301 206 11
Prefix-Tuning 64 tokens 12.807 268 9112.8072689112.807\,268\,9112.807 268 91 39.651 837 9439.6518379439.651\,837\,9439.651 837 94 65.725 809 3465.7258093465.725\,809\,3465.725 809 34 38.565 021 7538.5650217538.565\,021\,7538.565 021 75 31.034 481 5331.0344815331.034\,481\,5331.034 481 53 15.060 241 5215.0602415215.060\,241\,5215.060 241 52
[…]
linear projection matrices All LoRA 0.513 166 979 724 261 80.51316697972426180.513\,166\,979\,724\,261\,80.513 166 979 724 261 8 54.738 879 203 796 3954.7388792037963954.738\,879\,203\,796\,3954.738 879 203 796 39 $75.0$ 55.605 381 727 218 6355.6053817272186355.605\,381\,727\,218\,6355.605 381 727 218 63 45.977 011 322 975 1645.9770113229751645.977\,011\,322\,975\,1645.977 011 322 975 16 31.325 301 527 976 9931.3253015279769931.325\,301\,527\,976\,9931.325 301 527 976 99
$\bm{W}_{\text{in},x}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 60.831 719 636 917 11460.83171963691711460.831\,719\,636\,917\,11460.831 719 636 917 114 76.612 901 687 622 0776.6129016876220776.612\,901\,687\,622\,0776.612 901 687 622 07 63.452 917 337 417 663.452917337417663.452\,917\,337\,417\,663.452 917 337 417 6 52.873 563 766 479 4952.8735637664794952.873\,563\,766\,479\,4952.873 563 766 479 49 38.554 215 431 213 3838.5542154312133838.554\,215\,431\,213\,3838.554 215 431 213 38
$\bm{W}_{\text{in},z}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 46.324 950 456 619 2646.3249504566192646.324\,950\,456\,619\,2646.324 950 456 619 26 68.548 387 289 047 2468.5483872890472468.548\,387\,289\,047\,2468.548 387 289 047 24 45.739 910 006 523 1345.7399100065231345.739\,910\,006\,523\,1345.739 910 006 523 13 36.781 609 058 380 1336.7816090583801336.781\,609\,058\,380\,1336.781 609 058 380 13 24.698 795 378 208 1624.6987953782081624.698\,795\,378\,208\,1624.698 795 378 208 16
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.342 697 523 326 188 860.342697523326188860.342\,697\,523\,326\,188\,860.342 697 523 326 188 86 57.543 522 119 522 09557.54352211952209557.543\,522\,119\,522\,09557.543 522 119 522 095 77.419 352 531 433 177.419352531433177.419\,352\,531\,433\,177.419 352 531 433 1 58.744 394 779 205 3258.7443947792053258.744\,394\,779\,205\,3258.744 394 779 205 32 45.402 297 377 586 36545.40229737758636545.402\,297\,377\,586\,36545.402 297 377 586 365 37.349 396 944 046 0237.3493969440460237.349\,396\,944\,046\,0237.349 396 944 046 02
$\bm{W}_{\text{out}}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 61.798 840 761 184 6961.7988407611846961.798\,840\,761\,184\,6961.798 840 761 184 69 81.854 838 132 858 2881.8548381328582881.854\,838\,132\,858\,2881.854 838 132 858 28 65.246 635 675 430 365.246635675430365.246\,635\,675\,430\,365.246 635 675 430 3 45.402 297 377 586 36545.40229737758636545.402\,297\,377\,586\,36545.402 297 377 586 365 39.759 036 898 612 97639.75903689861297639.759\,036\,898\,612\,97639.759 036 898 612 976
S6 All Full 4.456 059 545 468 7924.4560595454687924.456\,059\,545\,468\,7924.456 059 545 468 792 56.673 115 491 867 06556.67311549186706556.673\,115\,491\,867\,06556.673 115 491 867 065 76.612 901 687 622 0776.6129016876220776.612\,901\,687\,622\,0776.612 901 687 622 07 57.847 535 610 198 97557.84753561019897557.847\,535\,610\,198\,97557.847 535 610 198 975 45.977 011 322 975 1645.9770113229751645.977\,011\,322\,975\,1645.977 011 322 975 16 34.939 759 969 711 30434.93975996971130434.939\,759\,969\,711\,30434.939 759 969 711 304
LoRA 0.464 394 215 127 723 30.46439421512772330.464\,394\,215\,127\,723\,30.464 394 215 127 723 3 56.286 269 426 345 82556.28626942634582556.286\,269\,426\,345\,82556.286 269 426 345 825 $75.0$ 56.502 240 896 224 97656.50224089622497656.502\,240\,896\,224\,97656.502 240 896 224 976 50.574 713 945 388 79450.57471394538879450.574\,713\,945\,388\,79450.574 713 945 388 794 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
${\bm{A}}$ Full 0.229 250 651 856 915 340.229250651856915340.229\,250\,651\,856\,915\,340.229 250 651 856 915 34 51.063 829 660 415 6551.0638296604156551.063\,829\,660\,415\,6551.063 829 660 415 65 71.370 965 242 385 8671.3709652423858671.370\,965\,242\,385\,8671.370 965 242 385 86 52.466 368 675 231 93452.46636867523193452.466\,368\,675\,231\,93452.466 368 675 231 934 42.528 736 591 339 1142.5287365913391142.528\,736\,591\,339\,1142.528 736 591 339 11 25.903 615 355 491 63825.90361535549163825.903\,615\,355\,491\,63825.903 615 355 491 638
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.292 506 518 569 153 82.29250651856915382.292\,506\,518\,569\,153\,82.292 506 518 569 153 8 47.195 357 084 274 2947.1953570842742947.195\,357\,084\,274\,2947.195 357 084 274 29 72.177 422 046 661 3872.1774220466613872.177\,422\,046\,661\,3872.177 422 046 661 38 46.860 986 948 013 30646.86098694801330646.860\,986\,948\,013\,30646.860 986 948 013 306 35.632 184 147 834 7835.6321841478347835.632\,184\,147\,834\,7835.632 184 147 834 78 22.891 566 157 341 00322.89156615734100322.891\,566\,157\,341\,00322.891 566 157 341 003
LoRA 0.347 144 240 365 408 70.34714424036540870.347\,144\,240\,365\,408\,70.347 144 240 365 408 7 55.029 010 772 705 0855.0290107727050855.029\,010\,772\,705\,0855.029 010 772 705 08 73.790 323 734 283 4573.7903237342834573.790\,323\,734\,283\,4573.790 323 734 283 45 56.726 455 688 476 5656.7264556884765656.726\,455\,688\,476\,5656.726 455 688 476 56 44.252 872 467 041 01644.25287246704101644.252\,872\,467\,041\,01644.252 872 467 041 016 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.848 333 380 596 381.848333380596381.848\,333\,380\,596\,381.848 333 380 596 38 56.769 824 028 015 1456.7698240280151456.769\,824\,028\,015\,1456.769 824 028 015 14 77.016 127 109 527 5977.0161271095275977.016\,127\,109\,527\,5977.016 127 109 527 59 59.417 039 155 960 0859.4170391559600859.417\,039\,155\,960\,0859.417 039 155 960 08 43.678 161 501 884 4643.6781615018844643.678\,161\,501\,884\,4643.678 161 501 884 46 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
LoRA 0.118 067 802 522 655 750.118067802522655750.118\,067\,802\,522\,655\,750.118 067 802 522 655 75 58.027 076 721 191 40658.02707672119140658.027\,076\,721\,191\,40658.027 076 721 191 406 78.629 034 757 614 1478.6290347576141478.629\,034\,757\,614\,1478.629 034 757 614 14 59.417 039 155 960 0859.4170391559600859.417\,039\,155\,960\,0859.417 039 155 960 08 48.850 575 089 454 6548.8505750894546548.850\,575\,089\,454\,6548.850 575 089 454 65 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
Conv1d Full 0.071 640 828 705 286 060.071640828705286060.071\,640\,828\,705\,286\,060.071 640 828 705 286 06 53.191 488 981 246 9553.1914889812469553.191\,488\,981\,246\,9553.191 488 981 246 95 74.596 774 578 094 4874.5967745780944874.596\,774\,578\,094\,4874.596 774 578 094 48 52.914 798 259 735 1152.9147982597351152.914\,798\,259\,735\,1152.914 798 259 735 11 43.678 161 501 884 4643.6781615018844643.678\,161\,501\,884\,4643.678 161 501 884 46 31.927 710 771 560 6731.9277107715606731.927\,710\,771\,560\,6731.927 710 771 560 67
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.021 641 500 338 055 160.021641500338055160.021\,641\,500\,338\,055\,160.021 641 500 338 055 16 49.613 153 934 478 7649.6131539344787649.613\,153\,934\,478\,7649.613 153 934 478 76 70.564 514 398 574 8370.5645143985748370.564\,514\,398\,574\,8370.564 514 398 574 83 50.448 429 584 503 17450.44842958450317450.448\,429\,584\,503\,17450.448 429 584 503 174 40.229 883 790 016 17440.22988379001617440.229\,883\,790\,016\,17440.229 883 790 016 174 25.903 615 355 491 63825.90361535549163825.903\,615\,355\,491\,63825.903 615 355 491 638
(a) Comprehensive experimental results on Spider using Mamba-1.4B.
Layer Method # Params (%) All Easy Medium Hard Extra
All All Full $100.0$ 71.760 153 770 446 7871.7601537704467871.760\,153\,770\,446\,7871.760 153 770 446 78 $87.5$ 73.542 600 870 132 4573.5426008701324573.542\,600\,870\,132\,4573.542 600 870 132 45 63.793 104 887 008 6763.7931048870086763.793\,104\,887\,008\,6763.793 104 887 008 67 51.807 230 710 983 27651.80723071098327651.807\,230\,710\,983\,27651.807 230 710 983 276
LoRA 0.804 873 122 224 144 40.80487312222414440.804\,873\,122\,224\,144\,40.804 873 122 224 144 4 70.889 747 142 791 7570.8897471427917570.889\,747\,142\,791\,7570.889 747 142 791 75 90.725 809 335 708 6290.7258093357086290.725\,809\,335\,708\,6290.725 809 335 708 62 73.991 030 454 635 6273.9910304546356273.991\,030\,454\,635\,6273.991 030 454 635 62 58.620 691 299 438 4858.6206912994384858.620\,691\,299\,438\,4858.620 691 299 438 48 45.783 132 314 682 0145.7831323146820145.783\,132\,314\,682\,0145.783 132 314 682 01
Prompt Prompt Tuning 64 tokens 0.005 917 985 961 427 6770.0059179859614276770.005\,917\,985\,961\,427\,6770.005 917 985 961 427 677 50.676 983 594 894 4150.6769835948944150.676\,983\,594\,894\,4150.676 983 594 894 41 75.403 225 421 905 5275.4032254219055275.403\,225\,421\,905\,5275.403 225 421 905 52 53.811 657 428 741 45553.81165742874145553.811\,657\,428\,741\,45553.811 657 428 741 455 37.356 323 003 768 9237.3563230037689237.356\,323\,003\,768\,9237.356 323 003 768 92 19.277 107 715 606 6919.2771077156066919.277\,107\,715\,606\,6919.277 107 715 606 69
Prefix-Tuning 1 token 10.820 948 510 794 74110.82094851079474110.820\,948\,510\,794\,74110.820 948 510 794 741 45.067 697 763 442 9945.0676977634429945.067\,697\,763\,442\,9945.067 697 763 442 99 $75.0$ 45.067 265 629 768 3745.0672656297683745.067\,265\,629\,768\,3745.067 265 629 768 37 32.183 909 416 198 7332.1839094161987332.183\,909\,416\,198\,7332.183 909 416 198 73 13.855 421 543 121 33813.85542154312133813.855\,421\,543\,121\,33813.855 421 543 121 338
[…]
linear projection matrices All LoRA 0.424 312 127 242 072 460.424312127242072460.424\,312\,127\,242\,072\,460.424 312 127 242 072 46 58.220 505 714 416 50458.22050571441650458.220\,505\,714\,416\,50458.220 505 714 416 504 74.596 774 578 094 4874.5967745780944874.596\,774\,578\,094\,4874.596 774 578 094 48 58.295 965 194 702 1558.2959651947021558.295\,965\,194\,702\,1558.295 965 194 702 15 51.724 135 875 701 90451.72413587570190451.724\,135\,875\,701\,90451.724 135 875 701 904 40.361 446 142 196 65540.36144614219665540.361\,446\,142\,196\,65540.361 446 142 196 655
$\bm{W}_{\text{in},x}$ LoRA 0.141 838 601 338 417 10.14183860133841710.141\,838\,601\,338\,417\,10.141 838 601 338 417 1 66.731 142 997 741 766.731142997741766.731\,142\,997\,741\,766.731 142 997 741 7 87.903 225 421 905 5287.9032254219055287.903\,225\,421\,905\,5287.903 225 421 905 52 67.713 004 350 662 2367.7130043506622367.713\,004\,350\,662\,2367.713 004 350 662 23 56.896 549 463 272 09556.89654946327209556.896\,549\,463\,272\,09556.896 549 463 272 095 42.771 083 116 531 3742.7710831165313742.771\,083\,116\,531\,3742.771 083 116 531 37
$\bm{W}_{\text{in},z}$ LoRA 0.141 838 601 338 417 10.14183860133841710.141\,838\,601\,338\,417\,10.141 838 601 338 417 1 65.377 175 807 952 8865.3771758079528865.377\,175\,807\,952\,8865.377 175 807 952 88 86.693 549 156 188 9686.6935491561889686.693\,549\,156\,188\,9686.693 549 156 188 96 68.834 078 311 920 1768.8340783119201768.834\,078\,311\,920\,1768.834 078 311 920 17 54.597 699 642 181 454.597699642181454.597\,699\,642\,181\,454.597 699 642 181 4 35.542 169 213 294 9835.5421692132949835.542\,169\,213\,294\,9835.542 169 213 294 98
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.283 275 408 799 057 940.283275408799057940.283\,275\,408\,799\,057\,940.283 275 408 799 057 94 65.183 752 775 192 2665.1837527751922665.183\,752\,775\,192\,2665.183 752 775 192 26 89.112 901 687 622 0789.1129016876220789.112\,901\,687\,622\,0789.112 901 687 622 07 67.264 574 766 159 0667.2645747661590667.264\,574\,766\,159\,0667.264 574 766 159 06 51.724 135 875 701 90451.72413587570190451.724\,135\,875\,701\,90451.724 135 875 701 904 37.951 806 187 629 737.951806187629737.951\,806\,187\,629\,737.951 806 187 629 7
$\bm{W}_{\text{out}}$ LoRA 0.141 838 601 338 417 10.14183860133841710.141\,838\,601\,338\,417\,10.141 838 601 338 417 1 67.021 274 566 650 3967.0212745666503967.021\,274\,566\,650\,3967.021 274 566 650 39 87.096 774 578 094 4887.0967745780944887.096\,774\,578\,094\,4887.096 774 578 094 48 69.058 293 104 171 7569.0582931041717569.058\,293\,104\,171\,7569.058 293 104 171 75 52.873 563 766 479 4952.8735637664794952.873\,563\,766\,479\,4952.873 563 766 479 49 46.385 541 558 265 68646.38554155826568646.385\,541\,558\,265\,68646.385 541 558 265 686
S6 All Full 4.438 752 155 800 24.43875215580024.438\,752\,155\,800\,24.438 752 155 800 2 65.667 313 337 326 0565.6673133373260565.667\,313\,337\,326\,0565.667 313 337 326 05 81.854 838 132 858 2881.8548381328582881.854\,838\,132\,858\,2881.854 838 132 858 28 68.834 078 311 920 1768.8340783119201768.834\,078\,311\,920\,1768.834 078 311 920 17 58.045 977 354 049 6858.0459773540496858.045\,977\,354\,049\,6858.045 977 354 049 68 40.963 855 385 780 33440.96385538578033440.963\,855\,385\,780\,33440.963 855 385 780 334
LoRA 0.383 804 949 263 572 40.38380494926357240.383\,804\,949\,263\,572\,40.383 804 949 263 572 4 63.926 500 082 015 9963.9265000820159963.926\,500\,082\,015\,9963.926 500 082 015 99 86.290 323 734 283 4586.2903237342834586.290\,323\,734\,283\,4586.290 323 734 283 45 68.161 433 935 165 468.161433935165468.161\,433\,935\,165\,468.161 433 935 165 4 49.425 286 054 611 20649.42528605461120649.425\,286\,054\,611\,20649.425 286 054 611 206 34.337 350 726 127 62534.33735072612762534.337\,350\,726\,127\,62534.337 350 726 127 625
${\bm{A}}$ Full 0.189 386 758 647 475 230.189386758647475230.189\,386\,758\,647\,475\,230.189 386 758 647 475 23 56.576 400 995 254 5256.5764009952545256.576\,400\,995\,254\,5256.576 400 995 254 52 77.016 127 109 527 5977.0161271095275977.016\,127\,109\,527\,5977.016 127 109 527 59 58.071 750 402 450 5658.0717504024505658.071\,750\,402\,450\,5658.071 750 402 450 56 45.977 011 322 975 1645.9770113229751645.977\,011\,322\,975\,1645.977 011 322 975 16 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.272 641 103 769 702 72.27264110376970272.272\,641\,103\,769\,702\,72.272 641 103 769 702 7 58.800 774 812 698 36458.80077481269836458.800\,774\,812\,698\,36458.800 774 812 698 364 79.032 260 179 519 6579.0322601795196579.032\,260\,179\,519\,6579.032 260 179 519 65 60.986 548 662 185 6760.9865486621856760.986\,548\,662\,185\,6760.986 548 662 185 67 50.574 713 945 388 79450.57471394538879450.574\,713\,945\,388\,79450.574 713 945 388 794 31.325 301 527 976 9931.3253015279769931.325\,301\,527\,976\,9931.325 301 527 976 99
LoRA 0.286 806 195 766 252 20.28680619576625220.286\,806\,195\,766\,252\,20.286 806 195 766 252 2 60.251 450 538 635 25460.25145053863525460.251\,450\,538\,635\,25460.251 450 538 635 254 82.661 288 976 669 3182.6612889766693182.661\,288\,976\,669\,3182.661 288 976 669 31 63.004 481 792 449 9563.0044817924499563.004\,481\,792\,449\,9563.004 481 792 449 95 46.551 725 268 363 9546.5517252683639546.551\,725\,268\,363\,9546.551 725 268 363 95 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.905 704 258 890 219 31.90570425889021931.905\,704\,258\,890\,219\,31.905 704 258 890 219 3 62.185 686 826 705 9362.1856868267059362.185\,686\,826\,705\,9362.185 686 826 705 93 82.258 063 554 763 882.258063554763882.258\,063\,554\,763\,882.258 063 554 763 8 65.695 065 259 933 4765.6950652599334765.695\,065\,259\,933\,4765.695 065 259 933 47 51.724 135 875 701 90451.72413587570190451.724\,135\,875\,701\,90451.724 135 875 701 904 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
LoRA 0.097 557 280 258 231 170.097557280258231170.097\,557\,280\,258\,231\,170.097 557 280 258 231 17 62.185 686 826 705 9362.1856868267059362.185\,686\,826\,705\,9362.185 686 826 705 93 80.241 936 445 236 280.241936445236280.241\,936\,445\,236\,280.241 936 445 236 2 66.591 930 389 404 366.591930389404366.591\,930\,389\,404\,366.591 930 389 404 3 49.425 286 054 611 20649.42528605461120649.425\,286\,054\,611\,20649.425 286 054 611 206 36.746 987 700 462 3436.7469877004623436.746\,987\,700\,462\,3436.746 987 700 462 34
Conv1d Full 0.059 183 362 077 336 010.059183362077336010.059\,183\,362\,077\,336\,010.059 183 362 077 336 01 62.475 824 356 079 162.475824356079162.475\,824\,356\,079\,162.475 824 356 079 1 81.854 838 132 858 2881.8548381328582881.854\,838\,132\,858\,2881.854 838 132 858 28 66.143 494 844 436 6566.1434948444366566.143\,494\,844\,436\,6566.143 494 844 436 65 51.149 427 890 777 5951.1494278907775951.149\,427\,890\,777\,5951.149 427 890 777 59 35.542 169 213 294 9835.5421692132949835.542\,169\,213\,294\,9835.542 169 213 294 98
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.017 847 482 626 446 640.017847482626446640.017\,847\,482\,626\,446\,640.017 847 482 626 446 64 50.967 115 163 803 150.967115163803150.967\,115\,163\,803\,150.967 115 163 803 1 70.967 739 820 480 3570.9677398204803570.967\,739\,820\,480\,3570.967 739 820 480 35 51.121 073 961 257 93551.12107396125793551.121\,073\,961\,257\,93551.121 073 961 257 935 42.528 736 591 339 1142.5287365913391142.528\,736\,591\,339\,1142.528 736 591 339 11 29.518 070 816 993 71329.51807081699371329.518\,070\,816\,993\,71329.518 070 816 993 713
(b) Comprehensive experimental results on Spider using Mamba-2.8B.
Table 10: Full experimental results on Spider (Yu et al., 2018) dataset. We report the accuracy ($\uparrow$) for Spider and its subsets. We consider two models in our experiments: Mamba-1.4B and Mamba-2.8B. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) Accuracy
Pretrained $0.0$ 0.081 5000.0815000.081\,5000.081 500
All All Full $100.0$ 59.960 001 707 077 02659.96000170707702659.960\,001\,707\,077\,02659.960 001 707 077 026
LoRA 1.922 527 283 438 617 31.92252728343861731.922\,527\,283\,438\,617\,31.922 527 283 438 617 3 60.350 000 858 306 88560.35000085830688560.350\,000\,858\,306\,88560.350 000 858 306 885
[…]
linear projection matrices All LoRA 1.017 230 519 270 278 41.01723051927027841.017\,230\,519\,270\,278\,41.017 230 519 270 278 4 62.790 000 438 690 18662.79000043869018662.790\,000\,438\,690\,18662.790 000 438 690 186
$\bm{W}_{\text{in},x}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 53.490 000 963 211 0653.4900009632110653.490\,000\,963\,211\,0653.490 000 963 211 06
$\bm{W}_{\text{in},z}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 58.149 999 380 111 69458.14999938011169458.149\,999\,380\,111\,69458.149 999 380 111 694
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.680 460 965 049 587 40.68046096504958740.680\,460\,965\,049\,587\,40.680 460 965 049 587 4 61.040 002 107 620 2461.0400021076202461.040\,002\,107\,620\,2461.040 002 107 620 24
$\bm{W}_{\text{out}}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 52.039 998 769 760 1352.0399987697601352.039\,998\,769\,760\,1352.039 998 769 760 13
S6 All Full 4.310 565 285 913 9444.3105652859139444.310\,565\,285\,913\,9444.310 565 285 913 944 55.510 002 374 649 0555.5100023746490555.510\,002\,374\,649\,0555.510 002 374 649 05
LoRA 0.923 912 723 244 536 30.92391272324453630.923\,912\,723\,244\,536\,30.923 912 723 244 536 3 43.959 999 084 472 65643.95999908447265643.959\,999\,084\,472\,65643.959 999 084 472 656
${\bm{A}}$ Full 0.456 748 639 567 040 370.456748639567040370.456\,748\,639\,567\,040\,370.456 748 639 567 040 37 61.210 000 514 984 1361.2100005149841361.210\,000\,514\,984\,1361.210 000 514 984 13
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.283 743 197 835 2022.2837431978352022.283\,743\,197\,835\,2022.283 743 197 835 202 49.509 999 155 998 2349.5099991559982349.509\,999\,155\,998\,2349.509 999 155 998 23
LoRA 0.692 192 750 875 576 50.69219275087557650.692\,192\,750\,875\,576\,50.692 192 750 875 576 5 52.270 001 173 019 4152.2700011730194152.270\,001\,173\,019\,4152.270 001 173 019 41
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.398 792 708 674 0611.3987927086740611.398\,792\,708\,674\,0611.398 792 708 674 061 34.540 000 557 899 47534.54000055789947534.540\,000\,557\,899\,47534.540 000 557 899 475
LoRA 0.234 957 666 087 183 550.234957666087183550.234\,957\,666\,087\,183\,550.234 957 666 087 183 55 56.489 998 102 188 1156.4899981021881156.489\,998\,102\,188\,1156.489 998 102 188 11
Conv1d Full 0.142 733 949 864 700 120.142733949864700120.142\,733\,949\,864\,700\,120.142 733 949 864 700 12 55.650 001 764 297 48555.65000176429748555.650\,001\,764\,297\,48555.650 001 764 297 485
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.043 414 909 750 512 950.043414909750512950.043\,414\,909\,750\,512\,950.043 414 909 750 512 95 58.090 001 344 680 78658.09000134468078658.090\,001\,344\,680\,78658.090 001 344 680 786
Table 11: Full experimenal results on the CIFAR-10 (Krizhevsky et al., 2009) dataset. We report accuracy ($\uparrow$). Mama-130M is utilized in this experiment. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. ${\bm{\beta}}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
C.3 Limitations of Applying Prompt-based Methods on SSMs
We provide the formal version of Proposition 1 and its corresponding proof here. We start by introducing the necessary notations. Denote the space of S4 mechanisms with $D$ channels as ${\mathcal{F}}_{\text{S4},D}$. Let ${\bm{H}}_{0}=({\bm{h}}_{0}^{(1)},{\bm{h}}_{0}^{(2)},\dots,{\bm{h}}_{0}^{(D)})\in{\mathbb{R}}^{H\times D}$ represent the initial hidden state, and ${\bm{X}}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$ denote the input sequence. The output of the S4 mechanism is represented as $f({\bm{X}};{\bm{H}}_{0})$. Furthermore, for $d$-th channel, let state transition matrix $\overline{{\bm{A}}}^{(d)}=\operatorname{diag}{(a_{1}^{(d)},\cdots,a_{H}^{(d)})}$ and input transition vector $\overline{{\bm{B}}}^{(d)}=(b_{1},\cdots,b_{H})^{\top}$, where $d=1,\ldots,D$. For any vector ${\bm{v}}\in{\mathbb{R}}^{n}$, we use ${\bm{v}}_{i:j}\in{\mathbb{R}}^{j-i}$ to denote the subvector of ${\bm{v}}$ containing elements from $i\in{\mathbb{N}}^{+}$ to $j\in{\mathbb{N}}^{+}$, where $i<j$. Similarly, for any matrix ${\bm{M}}\in{\mathbb{R}}^{m\times n}$, we use ${\bm{M}}_{i_{1}:j_{1},i_{2}:j_{2}}$ to denote the submatrix containing rows $i_{1}\in{\mathbb{N}}^{+}$ to $j_{1}\in{\mathbb{N}}^{+}$ and columns $i_{2}\in{\mathbb{N}}^{+}$ to $j_{2}\in{\mathbb{N}}^{+}$, where $i_{1}<j_{1}$, $i_{2}<j_{2}$.
Proposition 5 (Formal Version of Proposition 1).
Let $f\in{\mathcal{F}}_{\text{S4},D}$ be an S4 mechanism. Consider prefix-tuning that prepends a sequence ${\bm{P}}=({\bm{p}}_{1},\ldots,{\bm{p}}_{M})\in{\mathbb{R}}^{D\times M}$ to the input sequence ${\bm{X}}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$. For any prefix ${\bm{P}}\in{\mathbb{R}}^{D\times M}$, there exists an initial hidden state ${\bm{H}}_{0}^{\star}\in{\mathbb{R}}^{H\times D}$ such that the output of S4 after prefix-tuning and that after initial state tuning are identical, i.e., $f({\bm{X}};{\bm{H}}_{0}^{\star})\equiv f([{\bm{P}},{\bm{X}}];{\bm{H}}_{0})_{1:D,M+1:M+N}$ for all ${\bm{X}}\in{\mathbb{R}}^{D\times N}$.
Furthermore, assume that $\prod_{0\leq i<j\leq H}(a_{j}^{(d)}-a_{i}^{(d)})\neq 0$ and $\prod_{k=1}^{H}b_{k}^{(d)}\neq 0$ for all channels $d=1,\ldots,D$. Then the converse (i.e., for any ${\bm{H}}_{0}\in{\mathbb{R}}^{H\times D}$, there exists a ${\bm{P}}^{\star}\in{\mathbb{R}}^{D\times M}$ such that $f([{\bm{P}}^{\star},{\bm{X}}];{\bm{H}}_{0})_{1:D,M+1:M+N}\equiv f({\bm{X}};{\bm{H}}_{0}^{\star})$ for all ${\bm{X}}\in{\mathbb{R}}^{D\times N}$) holds if and only if $M\geq H$.
Proof of Proposition 5.
Given that operations in S4 are independent across all channels, we can, without loss of generality, consider the case where the number of channels $D=1$. Consequently, we can simplify our notation: the initial hidden states ${\bm{H}}_{0}\in{\mathbb{R}}^{H\times D}$ become ${\bm{h}}_{0}\in{\mathbb{R}}^{H}$, the input sequence ${\bm{X}}\in{\mathbb{R}}^{D\times N}$ becomes ${\bm{x}}\in{\mathbb{R}}^{N}$, and the prefix ${\bm{P}}\in{\mathbb{R}}^{D\times M}$ becomes ${\bm{p}}\in{\mathbb{R}}^{M}$. We omit the superscript $(d)$ denoting the channel index. To differentiate between the hidden states and output of prefix-tuned S4 (i.e., $f([{\bm{P}},{\bm{X}}];{\bm{H}}_{0})_{1:D,M+1:M+N}$) and initial state tuned S4 (i.e., $f({\bm{X}};{\bm{H}}_{0}^{\star})$), we introduce superscripts “PT” and “IST” respectively. The “PT” superscript denotes hidden states and output of S4 after prefix-tuning, while “IST” indicates those after initial state tuning.
We divide the proposition into two statements:
  1. 1.
    For any prefix ${\bm{p}}\in{\mathbb{R}}^{M}$, there exists an initial hidden state ${\bm{h}}_{0}^{\star}\in{\mathbb{R}}^{H}$ such that the output of S4 after prefix-tuning and that after initial state tuning are identical, i.e., $f({\bm{x}};{\bm{h}}_{0}^{\star})\equiv f([{\bm{p}},{\bm{x}}];{\bm{h}}_{0})_{M+1:N+M}$ for all ${\bm{x}}\in{\mathbb{R}}^{N}$.
  2. 2.
    Furthermore, assume that $\prod_{0\leq i<j\leq H}(a_{j}-a_{i})\neq 0$ and $\prod_{k=1}^{H}b_{k}\neq 0$. Then the converse (i.e., for any ${\bm{h}}_{0}\in{\mathbb{R}}^{H}$, there exists a ${\bm{p}}^{\star}\in{\mathbb{R}}^{M}$ such that $f([{\bm{p}}^{\star},{\bm{x}}];{\bm{h}}_{0})_{M+1:N+M}\equiv f({\bm{x}};{\bm{h}}_{0}^{\star})$ for all ${\bm{x}}\in{\mathbb{R}}^{N}$) holds if and only if $M\geq H$.
We will first prove the first statement and then proceed to prove the second statement.
Statement 1. The recurrent computation formulation of S4 in (2) implies that for each position $i$, the output $y_{i}$ depends solely on the previous hidden state $h_{i-1}$ and the current input $x_{i}$. Thus, to demonstrate that $f({\bm{x}};{\bm{h}}_{0}^{\star})\equiv f([{\bm{p}},{\bm{x}}];{\bm{h}}_{0})_{M+1:N+M}$ for all ${\bm{x}}\in{\mathbb{R}}^{N}$, it suffices to show that the hidden state for predicting output $y^{\text{IST}}_{1}$ equals that for predicting output $y^{\text{PT}}_{M+1}$, where $y^{\text{IST}}_{1}$ and $y^{\text{PT}}_{M+1}$ are outputs corresponding to the input $x_{1}$ for initial state tuning and prefix-tuning, respectively. In other words, it is sufficient to show that the initial state of initial-state-tuned model ${\bm{h}}^{\text{IST}}_{0}={\bm{h}}_{0}^{\star}$ is equal to the $(M+1)$-th hidden state of prefix-tuned model ${\bm{h}}^{\text{PT}}_{M+1}=\sum_{m=1}^{M}\overline{{\bm{A}}}^{M-m}\overline{{\bm{B}}}p_{m}$. When this equality holds, the subsequent hidden states and outputs for both versions of S4 will be identical, as the input sequence from that point onward is the same. Therefore, We prove the first statement by letting
${\bm{h}}_{0}^{\star}=\sum_{m=1}^{M}\overline{{\bm{A}}}^{M-m}\overline{{\bm{B}}}p_{m}.$ (11)
Statement 2. We aim to investigate the conditions under which there exists a ${\bm{h}}_{0}^{\star}\in{\mathbb{R}}^{H}$ such that for any ${\bm{p}}\in{\mathbb{R}}^{M}$, $f([{\bm{p}}^{\star},{\bm{x}}];{\bm{h}}_{0})_{M+1:N+M}\neq f({\bm{x}};{\bm{h}}_{0}^{\star})$. This is equivalent to demonstrating the existence of ${\bm{h}}_{0}^{\star}\in{\mathbb{R}}^{H}$ such that
${\bm{h}}_{0}^{\star}\neq\sum_{m=1}^{M}\overline{{\bm{A}}}^{M-m}\overline{{\bm{B}}}p_{m},\quad\text{for all }{\bm{p}}\in{\mathbb{R}}^{M}.$ (12)
This condition can be further reformulated as
${\mathbb{R}}^{H}\setminus\text{span}(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})\neq\emptyset,$ (13)
which is equivalent to
$\text{span}(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})\subsetneq{\mathbb{R}}^{H}.$ (14)
To determine when this condition holds, we analyze three distinct cases: (i) $M<H$, (ii) $M=H$, and (iii) $M>H$.
(Case 1: When $M<H$). In this scenario, it is obvious that (14) holds. The existence of such a ${\bm{h}}_{0}^{\star}$ is guaranteed because the dimension of the span is at most $M$, which is strictly less than $H$. This choice of ${\bm{h}}_{0}^{\star}$ ensures that it cannot be represented as a linear combination of the vectors in the span, thereby establishing the inequality.
(Case 2: When $M=H$). In this scenario, $\text{span}(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})={\mathbb{R}}^{H}$ if and only if $(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})$ are linearly independent. Note that
$\det(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})=\det(\overline{{\bm{A}}}^{M},\overline{{\bm{A}}}^{M-1},\ldots,{\bm{1}})\prod_{k=1}^{H}b_{k},$ (15)
where
[…]
Combining (15) and (17) yields
$\det(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})=(-1)^{\frac{H(H-1)}{2}}\prod_{0\leq i<j\leq H}^{H}(a_{j}-a_{i})\prod_{k=1}^{H}b_{k}.$ (18)
Therefore, if and only if $\prod_{1\leq i<j\leq H}(a_{j}-a_{i})\neq 0$ and $\prod_{k=1}^{H}b_{k}\neq 0$, we have
$\det(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})\neq 0,$ (19)
which is both necessary and sufficient for the linear independence of $(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})$, and consequently, for the condition in (14) to be satisfied.
(Case 3: When $M>H$). The analysis presented in case 2 extends naturally to this scenario.
The combination of the three cases above completes the proof of statement 2. ∎
C.4 Optimal Application of LoRA in SSM-based Models
Several studies (Hu et al., 2023; He et al., 2021) present findings on Transformers, indicating that applying LoRA to linear projection matrices yields performance comparable to or marginally superior to that of attention layers. In contrast, our experimental results on SSMs reveal that applying LoRA to linear projection matrices is more effective than applying it to S6 (see Table 3). To elucidate this phenomenon, we examine the influence of updating linear projection matrices on the model’s output.
Notations.
For the feasibility of the analysis, we consider a simplified SSM-based architecture which only consists of the input projection matrix $\bm{W}_{\text{in}}\in{\mathbb{R}}^{D\times D}$ and the S6 module parameterized by diagonal state transition matrices $\{{\bm{A}}^{(d)}\}_{d=1}^{D}$ with ${\bm{A}}^{(d)}\in{\mathbb{R}}^{H\times H}$, the weight matrices $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}\in{\mathbb{R}}^{H\times D}$ for computing input-dependent input transition vectors ${\bm{B}}_{n}\in{\mathbb{R}}^{H}$ and output mapping vectors ${\bm{C}}_{n}\in{\mathbb{R}}^{H}$, the down and up projection matrices ${\bm{W}}_{\bm{\Delta},\downarrow}\in{\mathbb{R}}^{D\times R},{\bm{W}}_{\bm{\Delta},\uparrow}\in{\mathbb{R}}^{R\times D}$ (where $R$ is the rank) for low-rank weight matrices for computing the input-depdenent step size $\bm{\Delta}_{n}=(\Delta^{(1)}_{n},\ldots,\Delta^{(D)}_{n})\in{\mathbb{R}}_{D}$, for $n=1,\ldots,N$. Define $\bm{W}_{\text{S6}}=[\bm{W}_{{\bm{B}}}^{\top},\bm{W}_{{\bm{C}}}^{\top},{\bm{W}}_{\bm{\Delta},\uparrow}^{\top}]^{\top}\in{\mathbb{R}}^{(2H+R)\times D}$. In the Mamba implementation, $\bm{W}_{\text{S6}}$ is implemented as the weight matrix of a single linear layer, referred to as x_proj in the codebase. Therefore, the parameters of the S6 can be formulated as
${\bm{\theta}}(\cdot;\{{\bm{A}}\}_{d=1}^{D},\bm{W}_{\text{S6}},{\bm{W}}_{\bm{\Delta},\downarrow},\bm{W}_{\text{in}})=\{\overline{{\bm{A}}}_{n},\overline{{\bm{B}}}_{n},{\bm{C}}_{n}\}_{n=1}^{N}.$ (20)
Consider input sequence ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$. Let ${\bm{Z}}=({\bm{z}}_{1},\ldots,{\bm{z}}_{N})\in{\mathbb{R}}^{D\times N}$ denote the intermediate output after the input projection. The intermediate output at position $n\in\{1,\ldots,N\}$ is
${\bm{z}}_{n}=\bm{W}_{\text{in}}\bm{x}_{n}.$ (21)
Note that
${\bm{B}}_{n}=\bm{W}_{{\bm{B}}}{\bm{z}}_{n},\quad{\bm{C}}_{n}=\bm{W}_{{\bm{C}}}{\bm{z}}_{n},\quad\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\bm{\Delta},\downarrow}{\bm{z}}_{n}+{\bm{\beta}}_{\bm{\Delta}}),$ (22)
and after discretization, we have
$\overline{{\bm{A}}}_{n}^{(d)}=\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)}),\quad\overline{{\bm{B}}}_{n}=\Delta_{n}^{(d)}{\bm{B}}_{n}=\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}{\bm{z}}_{n}.$ (23)
Combining (21), (22) and (23) yields
${\bm{\theta}}({\bm{X}};\{{\bm{A}}\}_{d=1}^{D},\bm{W}_{\text{S6}},{\bm{W}}_{\bm{\Delta},\downarrow},\bm{W}_{\text{in}})=\{\overline{{\bm{A}}}_{n},\overline{{\bm{B}}}_{n},{\bm{C}}_{n}\}_{n=1}^{N},\text{ where}$ (24)
$\overline{{\bm{A}}}_{n}^{(d)}=\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)}),\quad\overline{{\bm{B}}}_{n}=\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}\bm{W}_{\text{in}}\bm{x}_{n},\quad{\bm{C}}_{n}=\bm{W}_{{\bm{C}}}\bm{W}_{\text{in}}\bm{x}_{n},$ (25)
$\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}\bm{W}_{\text{in}}\bm{x}_{n}+{\bm{\beta}}_{\bm{\Delta}}).$ (26)
Theoretical Analysis.
In the following theorem, we demonstrate that applying LoRA exclusively to $\bm{W}_{\text{in}}$ is equivalent to applying it to $\bm{W}_{\text{S6}}$.
Lemma 6 (Detailed Version of Lemma 2).
Consider a model consists of an S6 module augmented with a linear input projection $\bm{W}_{\text{in}}\in{\mathbb{R}}^{D\times D}$. For any fine-tuned model where only $\bm{W}_{\text{S6}}$ is updated to $\overline{\bm{W}}_{\text{S6}}$, there exists $\widehat{{\bm{W}}}_{\text{in}}$ such that updating only $\bm{W}_{\text{in}}$ to $\widehat{{\bm{W}}}_{\text{in}}$ yields:
${\bm{\theta}}({\bm{X}};\{{\bm{A}}^{(d)}\}_{d=1}^{D},\overline{\bm{W}}_{\text{S6}},{\bm{W}}_{\bm{\Delta},\downarrow},\bm{W}_{\text{in}})={\bm{\theta}}({\bm{X}};\{{\bm{A}}^{(d)}\}_{d=1}^{D},\bm{W}_{\text{S6}},{\bm{W}}_{\bm{\Delta},\downarrow},\widehat{{\bm{W}}}_{\text{in}})$ (27)
Proof of Lemma 6.
In this proof, we use $\overline{\cdot}$ to denote the corresponding notations for the model with only $\bm{W}_{\text{S6}}$ updated, and use $\widehat{\cdot}$ to denote the corresponding notations for the model with only $\bm{W}_{\text{in}}$ updated.
To demonstrate (27), it is sufficient, according to (24), to find $\widehat{{\bm{W}}}_{\text{in}}$ that satisfies the following equations:
$\overline{{\bm{W}}}_{{\bm{C}}}\bm{W}_{\text{in}}=\bm{W}_{{\bm{C}}}\widehat{{\bm{W}}}_{\text{in}}$ (28)
$\overline{{\bm{W}}}_{\bm{\Delta},\uparrow}\bm{W}_{\text{in}}={\bm{W}}_{\bm{\Delta},\uparrow}\widehat{{\bm{W}}}_{\text{in}}$ (29)
$\overline{{\bm{W}}}_{{\bm{B}}}\bm{W}_{\text{in}}=\bm{W}_{{\bm{B}}}\widehat{{\bm{W}}}_{\text{in}}.$ (30)
Since $\bm{W}_{\text{S6}}=\begin{bmatrix}\bm{W}_{{\bm{B}}}\\ \bm{W}_{{\bm{C}}}\\ {\bm{W}}_{\bm{\Delta},\uparrow}\end{bmatrix}$, the three conditions (28) can be written as
$\overline{\bm{W}}_{\text{S6}}\bm{W}_{\text{in}}=\bm{W}_{\text{S6}}\widehat{{\bm{W}}}_{\text{in}}.$ (31)
By applying Singular Value Decomposition (SVD) to $\bm{W}_{\text{S6}}$ and $\left(\bm{W}_{\text{S6}}-\overline{\bm{W}}_{\text{S6}}\right)\bm{W}_{\text{in}}$, we obtain:
$\bm{W}_{\text{S6}}={\bm{U}}\begin{bmatrix}{\bm{\Sigma}}&{\bm{O}}_{(2H+R)\times(D-2H-R)}\end{bmatrix}{\bm{V}}^{\top},$ (32)
$\left(\bm{W}_{\text{S6}}-\overline{\bm{W}}_{\text{S6}}\right)\bm{W}_{\text{in}}={\bm{U}}^{\prime}\begin{bmatrix}{\bm{\Sigma}}^{\prime}&{\bm{O}}_{(2H+R)\times(D-2H-R)}\end{bmatrix}{\bm{V}}^{\prime\top},$ (33)
where ${\bm{U}},{\bm{U}}^{\prime}\in{\mathbb{R}}^{(2H+R)\times(2H+R)}$, ${\bm{\Sigma}},{\bm{\Sigma}}^{\prime}\in{\mathbb{R}}^{(2H+R)\times(2H+R)}$, and ${\bm{V}},{\bm{V}}^{\prime}\in{\mathbb{R}}^{D\times D}$. The diagonal elements of ${\bm{\Sigma}}$ and ${\bm{\Sigma}}^{\prime}$ are in decreasing order.
We let
$\bm{\widehat{W}}_{\textrm{in}}={\bm{V}}\begin{bmatrix}{\bm{\Sigma}}^{-1}{\bm{U}}^{\top}\overline{\bm{W}}_{\text{S6}}\bm{W}_{\text{in}}\\ {\bm{Q}}\end{bmatrix},$ (34)
where ${\bm{Q}}\in{\mathbb{R}}^{(D-2H-R)\times D}$ is an arbitrary matrix to be determined later. Plugging (32) and(34) back to $\bm{W}_{\text{S6}}\widehat{{\bm{W}}}_{\text{in}}$ and simplifying results in
$\bm{W}_{\text{S6}}\widehat{{\bm{W}}}_{\text{in}}$ (35)
$={\bm{U}}\begin{bmatrix}{\bm{\Sigma}}&{\bm{O}}_{(2H+R)\times(D-2H-R)}\end{bmatrix}{\bm{V}}^{\top}{\bm{V}}\begin{bmatrix}{\bm{\Sigma}}^{-1}{\bm{U}}^{\top}\overline{\bm{W}}_{\text{S6}}\bm{W}_{\text{in}}\\ {\bm{Q}}\end{bmatrix}$ ((32) & (34)) (36)
$=\overline{\bm{W}}_{\text{S6}}\bm{W}_{\text{in}},$ (Simplifying) (37)
which demonstrates that (31) is satisfied and completes the proof. ∎
Appendix D Details of Sec. 5: Selective Dimension Tuning
D.1 Understanding the Roles of State Matrix A, Input Transition Vector B, and Output Mapping Vector C for a Single Channel in S4 Modules
Problem Setting.
Inspired by Zeng & Lee (2024)’s theoretical analysis of LoRA’s expressive power, we adopt a similar framework to explore the expressive potential of various parameters in the S4 model. Specifically, we assume a target model that performs well on the intended task and a frozen model, which may be either pretrained or randomly initialized. Our goal is to identify a parameter-efficient method to update the frozen model so that it becomes functionally equivalent to the target model. In alignment with Zeng & Lee (2024), we assume that the frozen model’s capacity is equal to or exceeds that of the target model. This assumption is based on two main considerations: (i) analytical tractability, which necessitates that the frozen model must have the potential to match the functionality of the target model, and (ii) a practical rationale, given that the models typically used in practice are often overparameterized. Assume that both the target model and the frozen model are S4, with the target model having a hidden state dimension $H_{\star}$ and the frozen model having a hidden state dimension $H\geq H_{\star}$. Meanwhile, suppose that all the hidden dimensions of both models are valid, meaning that none of the parameter elements are zero. The target model, frozen model, and the updated model after tuning the parameters on the frozen model can be formulated using discretized parameters $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$ as follows:
(Target model) $f^{\star}({\bm{x}})_{n}=\sum_{m=1}^{n}{\bm{C}}_{\star}\overline{{\bm{A}}}_{\star}^{m-n}\overline{{\bm{B}}}_{\star}x_{m},\text{ where }\operatorname{diag}(\overline{{\bm{A}}}_{\star}),\overline{{\bm{B}}}_{\star},{\bm{C}}_{\star}\in{\mathbb{R}}^{H_{\star}},$ (38)
(Frozen model) $f_{0}({\bm{x}})_{n}=\sum_{m=1}^{n}{\bm{C}}\overline{{\bm{A}}}^{m-n}\overline{{\bm{B}}}x_{m},\text{ where }\operatorname{diag}(\overline{{\bm{A}}}),\overline{{\bm{B}}},{\bm{C}}\in{\mathbb{R}}^{H},$ (39)
(Updated model) $\hat{f}({\bm{x}})_{n}=\sum_{m=1}^{n}\widehat{{\bm{C}}}\widehat{\overline{{\bm{A}}}}^{m-n}\widehat{\overline{{\bm{B}}}}x_{m},\text{ where }\operatorname{diag}(\widehat{\overline{{\bm{A}}}}),\widehat{\overline{{\bm{B}}}},\widehat{{\bm{C}}}\in{\mathbb{R}}^{H}.$ (40)
Parameter Efficiency Analysis on S4.
Let ${\mathcal{P}}^{H}$ denote the set of all $H\times H$ permutation matrices. Given this formulation, we present our first analysis of parameter efficiency for the S4 model in the following lemma. This analysis is based on the parameters after necessary discretization $(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})$.
Lemma 0 (Essential Discretized Parameter Set for S4).
Consider the parameters after discretization, i.e., $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\otimes\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
Proof of Lemma 3.
The key idea of this proof is straightforward. To facilitate the analysis and update the frozen model to be equivalent to the target model, we first equalize the number of hidden state dimensions between the two models. This is achieved by expanding the target model’s ${\bm{A}}_{\star}$, ${\bm{B}}_{\star}$, and ${\bm{C}}_{\star}$ to match the $H$ hidden state dimensions of the frozen model, padding the additional $H-H^{\star}$ dimensions with zeros.
Define $\otimes$ as the element-wise product. We can express the target model as:
$f^{\star}(\bm{x})_{n}$ $=\sum_{m=1}^{n}\begin{bmatrix}{\bm{C}}_{\star}&{\bm{0}}^{\top}\end{bmatrix}\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}^{n-m}\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}x_{m}$ (42)
$=\sum_{m=1}^{n}\operatorname{diag}\left(\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}\right)^{n-m}\left(\begin{bmatrix}{\bm{C}}_{\star}^{\top}\\ {\bm{0}}\end{bmatrix}\otimes\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}\right)x_{m}$ (43)
Consider any permutation matrix ${\bm{P}}\in{\mathcal{P}}^{H}$. Applying ${\bm{P}}$ to permute the frozen model leaves the model functionally unchanged:
$f_{0}(\bm{x})_{n}$ $=\sum_{m=1}^{n}{\bm{C}}\overline{{\bm{A}}}^{n-m}\overline{{\bm{B}}}x_{m}=\sum_{m=1}^{n}{\bm{C}}{\bm{P}}\left({\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right)^{n-m}{\bm{P}}^{\top}\overline{{\bm{B}}}x_{m}$ (44)
$=\sum_{m=1}^{n}\operatorname{diag}\left({\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right)^{n-m}\left(({\bm{P}}^{\top}{\bm{C}}^{\top})\otimes({\bm{P}}^{\top}\overline{{\bm{B}}})\right)x_{m}$ (45)
Therefore, to make the updated model equivalent to the target model, we need to update ${\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}$ to align with $\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}$, and $({\bm{P}}^{\top}{\bm{C}}^{\top})\otimes({\bm{P}}^{\top}\overline{{\bm{B}}})$ to align with $\begin{bmatrix}{\bm{C}}_{\star}^{\top}\\ {\bm{0}}\end{bmatrix}\otimes\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}$. If they are already matching or partially matched for certain entries, no updates are required for those entries; only the unmatched entries need to be updated. Then, the required trainable parameters for this permutation matrix ${\bm{P}}$ are:
$\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\otimes\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}+\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}+\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}.$
Optimizing the permutation matrix ${\bm{P}}\in{\mathcal{P}}^{H}$ yields the desired results. ∎
This lemma highlights the significance of identifying essential hidden state dimensions. The term $\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\otimes\overline{{\bm{B}}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}$ underscores the importance of excluding redundant dimensions. This can be achieved by either directly removing these dimensions from the state matrix $\overline{{\bm{A}}}$, or by updating $\overline{{\bm{B}}}$ or ${\bm{C}}$ to ensure that only the selected hidden state dimensions are utilized during the input transition or output mapping phases. Once redundant dimensions are filtered out, tuning only the essential dimensions is sufficient to align the updated model with the target model.
Furthermore, based on the lemma, the roles of the input transition vector $\overline{{\bm{B}}}$ and ${\bm{C}}^{\top}$ are nearly identical, as they consistently appear together as the combined term $\overline{{\bm{B}}}\otimes{\bm{C}}^{\top}$, which is also discussed in Gupta et al. (2022). Consequently, one could opt to tune either $\overline{{\bm{B}}}$ or ${\bm{C}}$ exclusively or alternatively, split the indices into two groups, tuning $\overline{{\bm{B}}}$ for the first group and ${\bm{C}}$ for the second. Both vectors indicate how information from different hidden state dimensions is integrated, whereas $\overline{{\bm{A}}}$ plays a distinct role, determining how the hidden states are stored.
In practice, instead of directly using the discretized parameters $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$, S4 is implemented using the continuous parameters ${\bm{A}},{\bm{B}},{\bm{C}}$ with step size $\Delta$. To provide further practical guidance on parameter tuning, the following two lemmas analyze the parameter efficiency of continuous parameters under different discretization methods: Two exemplary methods of discretization are bilinear and zero-order hold (ZOH):
[…]
Lemma 7 (Essential Continuous Parameter Set for S4 with Bilinear Discritization).
Consider the parameters before discretization, i.e., ${\bm{A}},{\bm{B}},{\bm{C}}$, and they are discretized via bilinear discretization. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[\Delta{\bm{P}}^{\top}(\operatorname{diag}({\bm{I}}+\Delta/2{\bm{A}})\otimes{\bm{B}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}{\bm{A}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-{\bm{A}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}({\bm{B}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-{\bm{B}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
Proof of Lemma 7.
Combining Lemma 3 and the Bilinear discretization method in (47) yields the desired results. ∎
Lemma 8 (Essential Continuous Parameter Set for S4 with ZOH Discritization).
Consider the parameters before discretization, i.e., ${\bm{A}},{\bm{B}},{\bm{C}}$, and they are discretized via ZOH discretization. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[\Delta{\bm{P}}^{\top}(\operatorname{diag}(\exp(\Delta{\bm{A}})-{\bm{I}})\otimes{\bm{B}}\otimes{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}{\bm{A}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-{\bm{A}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}({\bm{B}}\otimes{\bm{C}}^{\top})\right]_{1:H_{\star}}-{\bm{B}}_{\star}\otimes{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
Proof of Lemma 8.
Combining Lemma 3 and the ZOH discretization method in (47) yields the desired results. ∎
The insights provided by Lemma 7 and Lemma 8 are the same as those provided by Lemma 3. The analysis here supports the second step of SDLoRA presented in Sec. 5.
D.2 Extension to Deep S4 Models
Our previous analysis focused on single-channel S4 models. We now expand our investigation to more complex scenarios involving deep S4 models for both target and frozen architectures, incorporating $D$ channels and varying layer depths. In this section, we consider two PEFT methods: (i) Selective Dimension Tuning (SDT) and (ii) SDLoRA. The key distinction between SDT and SDLoRA lies in their treatment of linear projection matrices. SDT exclusively updates the columns of weight matrices corresponding to the updatable channels identified through Alg. 1. In contrast, SDLoRA employs LoRA to modify these matrices. It is worth noting that the linear projection matrix updates in SDT are inherently low-rank, making it a specialized case of SDLoRA. Our analysis starts with SDT, and it automatically applies to SDLoRA.
In this analysis, we assume that each input token $x_{t}$ belongs to ${\mathcal{X}}$, a bounded subset of ${\mathbb{R}}^{D}$, and that the length of the input sequence is finite. Let the frozen model have $L$ layers, and the target model have $L^{\star}$ layers, where $L\geq L^{\star}$. Similar to the technique used in Zeng & Lee (2024) and Giannou et al. (2023). The basic idea of updating the frozen model to match the functionality of the target model is to utilize every $\lceil L/L^{\star}\rceil$ layers of the frozen model to approximate every layer of the target model. We start introducing this proof idea from the simplest case where $L^{\star}=1,L=D$. In this scenario, we can simply choose one different channel to tune and maintain all other channels at zero at every layer. The outputs from the various channels of the deep S4 layers are then combined through a residual connection. This proof idea inspires us to perform channel selection and make use of the residual connections, which is the first and third step of SDLoRA presented in Sec. 5. Building on this idea, we present the following results for when the target model has only $L^{\star}=1$ layer, and $L=D=2$.
Lemma 9.
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target one-layer deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$. Then this can be achieved by selectively fine-tuning at most $\lceil D/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
Proof of Lemma 9.
In this proof, we start by considering the case where $L=D$. In this case, we update a single distinct channel for each layer while setting the other channels to zero. Essentially, we modify the frozen model so that each layer corresponds to and functions as an individual channel in the target model. To be more specific, we fully update the first channel in the first layer to match the first channel of the target model, second channel in the second layer to match the second channel of the target model, so on and so forth.
For the $l$-th layer of the frozen model , we append subscript $l$ to all parameters of the deep S4 layer as introduced in (4). For the $d$-th channel, corresponding notations are denoted with a superscript $(d)$. We define the $t$-th intermediate output token of the $l$-th deep S4 layer as ${\bm{z}}_{l,t}\in{\mathbb{R}}^{D}$. Additionally, the updated S4 module in layer $l$ is denoted as $\widehat{\operatorname{S4}}_{l}$, with $\widehat{\operatorname{S4}}_{l,t}$ referring specifically to the sub-function that outputs the $t$-th token. Therefore, for the $t$-th intermediate output token of the $l$-th deep S4 layer of the updated model can be written as
$\begin{split}{\bm{z}}_{l,t}&=\widehat{{\bm{W}}}_{l}\cdot\widehat{\operatorname{S4}}_{l,t}({\bm{z}}_{l-1,1},\ldots,{\bm{z}}_{l-1,t})+\widehat{{\bm{\beta}}}_{l}+\widehat{{\bm{u}}}_{l}\otimes{\bm{z}}_{l-1,t}\\ &=\widehat{{\bm{W}}}_{l}\cdot\begin{bmatrix}\widehat{\operatorname{S4}}_{l,t}^{(1)}(z_{l-1,1}^{(1)},\ldots,z_{l-1,t}^{(1)})\\ \vdots\\ \widehat{\operatorname{S4}}_{l,t}^{(D)}(z_{l-1,1}^{(D)},\ldots,z_{l-1,t}^{(D)})\\ \end{bmatrix}+\widehat{{\bm{\beta}}}_{l}+\widehat{{\bm{u}}}_{l}\otimes{\bm{z}}_{l-1,t},\end{split}$ (50)
where $\widehat{{\bm{W}}}_{l}\in{\mathbb{R}}^{D\times D},\widehat{{\bm{\beta}}}_{l}\in{\mathbb{R}}^{D}$ are the updated weight and biases of the $l$-th layer of the frozen model, and $\widehat{{\bm{u}}}_{l}\in{\mathbb{R}}^{D}$ is the updated residual connection weight of the frozen model.
For layers $l<L=D$.
We follow the steps provided in Sec. 5 to update the $l$-th layer of the frozen model such that it functionally equivalent to the $l$-th channel of the target model. For the reader’s convinence, we restate our strategies here:
  • (Channel Selection) Select $D^{\prime}\leq D$ ($D^{\prime}=1$ here) important channels for making predictions. Any channel $d$ that is not utilized will have their corresponding ${\bm{C}}^{(d)}$ set to zero, eliminating the need to update parameters for ${\bm{A}}^{(d)}$ and the $d$-th column of ${\bm{W}}$. To be more specific, we let ${\bm{C}}^{(d)}={\bm{0}}$ for all $d\neq l$ in this scenario.
  • (Hidden State Selection) Within the selected channels, select $H^{\prime}\leq H$ important hidden states. For any hidden state that is not used within a selected channel $d$, the corresponding element in ${\bm{C}}^{(d)}$ will be set to zero, thus eliminating the need to tune the corresponding element in ${\bm{A}}^{(d)}$. To be more specific, we can achieve $\widehat{\operatorname{S4}}_{l,t}^{(l)}(\cdot)=\operatorname{S4}_{\star,t}^{(l)}(\cdot)$ by Lemma 3.
  • (Residual and Bias Tuning) Regardless of other selections, SDLoRA consistently tunes the coefficients of residual connections and biases in linear projections, as these components contain a negligible number of parameters. In this scenario, we let $\widehat{{\bm{\beta}}}_{l}={\bm{0}},\leavevmode\nobreak\ \widehat{{\bm{u}}}_{l}=\begin{bmatrix}\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{l-1\text{ elements}}}&0&\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{D-l\text{ elements}}}\end{bmatrix}^{\top}.$
This construction yields
${\bm{z}}_{l,t}=\begin{bmatrix}z_{l-1,t}^{(1)}&\ldots&z_{l-1,t}^{(l-1)}&\operatorname{S4}_{\star,t}^{(l)}(z_{l,1}^{(l)},\ldots,z_{l,t}^{(l)})&z_{l-1,t}^{(l+1)}&\ldots&z_{l-1,t}^{(D)}\end{bmatrix}^{\top}.$ (51)
Consequently, only the $l$-th channel is active in the $l$-th layer, while all other layers function as identity mappings, propagating the output of the preceding layer without modification.
For layer $l=L=D$.
Based on the setup of the first $L-1$ layers, we have
${\bm{z}}_{L-1,t}=\begin{bmatrix}\operatorname{S4}_{\star,t}^{(1)}(x^{(1)})&\cdots&\operatorname{S4}_{\star,t}^{(L-1)}(x^{(L-1)})&x^{(L)}\end{bmatrix}^{\top}.$ (52)
For the last layer, we let
$\widehat{{\bm{W}}}_{L}={\bm{W}}_{\star},\quad\widehat{{\bm{\beta}}}_{L}={\bm{\beta}}_{\star},\quad\widehat{{\bm{u}}}_{L}={\bm{0}},$ (53)
[…]
It is easy to verify that the output of the updated frozen model is identical to the output of the target model, i.e.,
$\bm{y}_{t}={\bm{z}}_{L,t}={\bm{W}}_{\star}\begin{bmatrix}\operatorname{S4}_{\star,t}^{(1)}(x^{(1)})&\cdots&\operatorname{S4}_{\star,t}^{(L-1)}(x^{(L-1)})&\operatorname{S4}_{\star,t}^{(L)}(x^{(L)})\end{bmatrix}^{\top}+{\bm{\beta}}_{\star}.$ (55)
Thus far, we have demonstrated that the statement holds when $L=D$. This analysis can be readily extended to cases where $L\neq D$ by tuning $\lceil D/L\rceil$ channels at each layer. For example, when $L=D/2$, we can tune two channels per layer using a construction similar to the one described above. This generalization completes the proof. ∎
Theorem 10 (Expressive Power of SDLoRA on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer.
Proof of Theorem 10.
We update every $\lceil D/L\rceil$ layers of the frozen model to approximate each layer of the target model. By applying Lemma 9 iteratively to each set of $\lceil D/L\rceil$ layers, we obtain the desired result. ∎
For reader’s convience, we restate the following statement presented in the main body again here.
Theorem 0 (Expressive Power of SDLoRA on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDLoRA, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states on SSM modules, applying rank-$\lceil\frac{L}{L^{\star}}\rceil$ updates on linear projection matrices and updating residual connections and biases at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
Proof of Theorem 4.
Since SDT is a special case of SDLoRA, Theorem 10 directly implies the desired statement. ∎
SDLoRA for Mamba.
In the Mamba model, the output mapping vector ${\bm{C}}$ is input-dependent, making it unsuitable for direction modification. Therefore, we focus our channel and hidden state selection solely on ${\bm{A}}$. For any channels or hidden states that are not selected, we set the corresponding elements of ${\bm{A}}$ to minimal values, effectively setting the associated entries in $\overline{{\bm{A}}}$ to zero. For channels and states that are updatable, we update the corresponding entries for ${\bm{A}}$. However, since ${\bm{B}}^{(d)}$ and ${\bm{C}}^{(d)}$ cannot be directly updated, we modify the corresponding weight matrices that compute these vectors. Specifically, for updatable channels, we update the corresponding columns in ${\bm{W}}_{{\bm{B}}}$ and ${\bm{W}}_{{\bm{C}}}$; for updatable states, we adjust the corresponding rows in these weight matrices.
D.3 Experiments on Deep S4 Models
Synthetic.
For selecting channels and hidden states, we initiate with a warmup learning rate between $21e-$ 2 and $31e-$ 3 and conduct 20 warmup iterations. Learning rates are adjusted between $25e-$ 2, $21e-$ 2, $35e-$ 3, and $31e-$ 3. We apply LoRA with ranks of 2 and 4 to the SSM and with ranks of 4, 8, and 16 to the linear projection matrices. Non-zero states are selected from the sets {4, 8}, and non-zero channels from {8, 16}.
We additionally consider SDT (Selective Dimension Tuning), which is introduced in Sec. D.2, and the results are visualized in Fig. 2. We observe that SDT even outperforms SDLoRA in this synthetic experiments, demonstrating highly promising performance. Unfortunetly, we fail to make it work on pretraned Mamba, and identify it as one of the promising future directions.
Figure 2: Performance comparison between various methods. SDT (Selective Dimension Tuning) is compared to SDLoRA. Unlike SDLoRA, which applies LoRA to linear projection matrices, SDT performs sparse tuning on linear projection matrices by updating only the columns corresponding to channels selected via Alg. 1. Notably, SDT achieves superior performance, matching full fine-tuning results while using only 25% of the parameters, and even surpassing full fine-tuning with more parameters. Extending SDT to real datasets is considered a promising future direction for SDLoRA.
CIFAR-10 (Krizhevsky et al., 2009).
We adhere to the preprocessing steps for CIFAR-10 as outlined by Gu et al. (2022a). The LoRA ranks for linear projection matrices are tuned among {1, 2, 4, 8, 16}, and for the S4 component, ranks are set from {1, 2, 4}. Non-zero states are chosen from {8, 12, 16}, and non-zero channels from {48, 64}. A warmup phase includes 1 epoch with a learning rate of $21e-$ 2. For linear projection matrices, LoRA ranks are explored at {2, 4, 8, 16}, and for the SSM, ranks at {2, 4, 8}. All state dimensions are updated, and channel dimensions considered for updates are {4, 8, 16, 32}.
D.4 Experiments on Pretrained Mamba
Here, we provide more experiment details. Unless otherwise stated, our experiment setting is identical to Sec. C.1. For LoRA, we consider three different LoRA configurations at each layer, involving the following matrices which comprise most of the parameters: $\bm{W}_{\text{out}}$ (output linear projection), $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$ (weight matrices for computing input-dependent ${\bm{B}}_{n},{\bm{C}}_{n}$), and ${\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\bm{\Delta},\uparrow}$ (down and up projection matrices of LoRA adapters for computing $\bm{\Delta}$). The three LoRA application methods are: (i) $\bm{W}_{\text{out}}$, $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\bm{\Delta},\uparrow}$; (ii) $\bm{W}_{\text{out}},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$ and ${\bm{W}}_{\bm{\Delta},\downarrow}$; and (iii) $\bm{W}_{\text{out}}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$. For SDLoRA, we set the channel freeze ratio at 99% across all scenarios. We select the state freeze ratio $\alpha$ from the set ${75\%,90\%,95\%}$ and apply LoRA exclusively to $\bm{W}_{\text{out}}$ to maintain a comparable number of trainable parameters. Residual connections and bias are frozen in this experiment. For the warmup, we employ 500 data batches to fully train the SSM modules prior to dimension selection, except for the RTE task in GLUE, where we use 250 batches due to its limited dataset size. Note that the parameters are reverted back after the warmup stage.