### v3

# Kevin Galim    Wonjun Kang    Yuchen Zeng    Hyung Il Koo    Kangwook Lee

Abstract
Deep State Space Models (SSMs), such as Mamba (Gu & Dao, 2024), have become powerful tools for language modeling, offering high performance and linear scalability with sequence length. However, the application of parameter-efficient fine-tuning (PEFT) methods to SSM-based models remains largely underexplored. We start by investigating two fundamental questions on existing PEFT methods: (i) How do they perform on SSM-based models? (ii) Which parameters should they target for optimal results? Our analysis shows that LoRA and its variants consistently outperform all other PEFT methods. While LoRA is effective for linear projection matrices, it fails on SSM modules—yet still outperforms other methods applicable to SSMs, indicating their limitations. This underscores the need for a specialized SSM tuning approach. To address this, we propose Sparse Dimension Tuning (SDT), a PEFT method tailored for SSM modules. Combining SDT for SSMs with LoRA for linear projection matrices, we achieve state-of-the-art performance across extensive experiments.
Machine Learning, ICML
[…]
1 Introduction
In the past few years, Large Language Models (LLMs) such as ChatGPT (Achiam et al., 2023; Brown et al., 2020) have achieved groundbreaking performance and are now widely used in daily life. While many models rely on the Transformer architecture (Vaswani et al., 2017), its quadratic time complexity due to the attention mechanism poses challenges for long sequences. To address this, alternative architectures such as linear attention (Katharopoulos et al., 2020), RWKV (Peng et al., 2023), RetNet (Sun et al., 2023), and Mamba (Gu & Dao, 2024) have been developed, offering subquadratic time complexity. Efficient attention alternatives often rely on State Space Models (SSMs) or their variants (Gu et al., 2021, 2022b, 2022a; Gu & Dao, 2024), which are akin to linear RNNs, maintaining hidden states of fixed size for sequential processing. S4 (Gu et al., 2022b, a) overcomes RNNs’ parallel training limitations by constraining parameter structures, enabling a convolutional form for efficient parallel computation. S6 (Gu & Dao, 2024) improves this with input-dependent parameters, enabling selective focus on relevant information per token. Building on S6 with linear projection matrices (analogous to the Feed-Forward Networks in Transformer layers), Mamba-I (Gu & Dao, 2024) emerged as a prominent SSM-based model. Mamba-I was later extended to Mamba-II (Dao & Gu, 2024), with both models achieving Transformer-level performance in language modeling and gaining widespread recognition.
As SSMs gain popularity, performing parameter-efficient fine-tuning (PEFT) on pretrained models for downstream tasks is crucial, since full fine-tuning is costly and inefficient. Numerous PEFT methods (Houlsby et al., 2019; Hu et al., 2021; He et al., 2021; Li & Liang, 2021; Lester et al., 2021; Zaken et al., 2022; Liu et al., 2021, 2022; Houlsby et al., 2019) have been developed, achieving notable success on Transformer models. The most popular PEFT methods fall into three categories: (i) input-injection methods, which add sequences to the model’s main input (Lester et al., 2021) or prepend tokens to the intermediate inputs at each layer (Li & Liang, 2021); (ii) architecture-enhancement methods, which adjust the model architecture. For example, Houlsby et al. (2019) added layers between Transformer layers, while Additional-scan (Yoshimura et al., 2025) expands state dimensions in the SSM module; (iii) weight-tuning methods, which directly modify existing model weights. Notable weight-tuning approaches include BitFit (Zaken et al., 2022), which updates only bias terms, and LoRA (Hu et al., 2021), which modifies weight matrices through low-rank updates, along with its variants such as DoRA (Liu et al., 2024) and LoRA+ (Hayou et al., 2024). For simplicity, we denote LoRA and its variants as LoRA.
Figure 1: A visual guide to PEFT methods in SSM-based models: benchmarking and innovation. We compare various existing PEFT approaches on SSM-based models, demonstrating that LoRA applied to linear projection matrices outperforms all other methods. However, extending LoRA to SSM modules fails to yield further improvements. To address this, we propose Sparse Dimension Tuning (SDT), which achieves state-of-the-art performance on SSM-based models when combined with LoRA for linear projection matrices.
Despite the success that existing PEFT methods have achieved in adapting Transformer-based models, their efficacy in adapting SSM-based models remains largely underexplored, leaving many interesting questions open.
  1. 1.
    Do existing popular PEFT methods remain effective for SSM-based models?
  2. 2.
    If applicable, what is the optimal way to integrate these methods into SSM-based models, and which parameters should be updated?
  3. 3.
    If not, can we design specialized variants tailored to SSMs that yield superior performance?
Our main contributions to address these questions are:
  • Comprehensive Benchmarking of PEFT Methods. We benchmark six widely used PEFT methods across three categories on diverse tasks, including natural language understanding, generation, and computer vision. We evaluate these methods on both SSM-based models (i.e., Mamba) and a hybrid model (i.e., Jamba (Lieber et al., 2025)), which consists of both Transformer layers and Mamba layers. Our results show that LoRA consistently outperforms all other PEFT methods on both SSM-based and hybrid models. However, its effectiveness is limited to linear projection matrices, as further tuning of SSM modules does not improve performance. Notably, other methods applicable to SSM modules perform worse than LoRA, further underscoring the need for a specialized approach to tuning SSM modules.
  • Introducing Sparse Dimension Tuning (SDT) for SSM Modules. To develop an effective method for tuning SSM modules, we conduct a theoretical analysis to understand the roles of different parameters. This analysis motivates the Sparse Dimension Tuning and Pruning (SDT-P) method, which improves efficiency by freezing and pruning certain channel and state dimensions while training only the remaining ones. We establish theoretical guarantees for its effectiveness in SSM-based models when combined with LoRA applied to linear projection matrices. We then simplify SDT-P into Sparse Dimension Tuning (SDT) by omitting explicit pruning, as pruned dimensions can be considered equivalent to training dimensions set to zero. SDT selectively updates channels and fine-tunes specific dimensions within them, as illustrated in Fig. 1.
  • Demonstrating Effectiveness of SDT. Through extensive experiments, we demonstrate that integrating SDT into SSM-based models, combined with applying LoRA to their linear projection matrices, achieves state-of-the-art fine-tuning performance.
The roadmap of our paper is illustrated in Fig. 1. Our code is available at https://github.com/furiosa-ai/ssm-peft.
2 Related Works
Concurrent Works of PEFT on SSMs.
Several concurrent studies (Halloran et al., 2024; Yoshimura et al., 2025; Kang et al., 2025) have investigated PEFT methods for SSM-based models. Halloran et al. (2024) studied both in-context learning and parameter-efficient fine-tuning, with an orthogonal focus on analyzing Mamba’s stability under mixed-precision training using Lyapunov exponents. Kang et al. (2025) introduced state-based PEFT methods and proposed State-offset Tuning, solely focusing fine-tuning Mamba’s S6 blocks. Yoshimura et al. (2025) benchmarked multiple PEFT approaches—including established methods and a new method called Additional-scan (which adds a trainable state dimension to the SSM module), plus partial tuning (fine-tuning only a subset of parameters)—and introduced MambaPEFT through PEFT search strategies. While Yoshimura et al. (2025) solely focused on Mamba-I, providing an in-depth study of that particular architecture, our work investigates a broader class of SSM-based models including deep S4, Mamba-I, Jamba in the main body, as well as Mamba-II presented in Sec. C.2 and E.2, aiming to offer general insights on how to effectively tune SSMs rather than focusing on a single variant.
Sparse Tuning.
Several studies have explored sparse parameter selection in fine-tuning (Song et al., 2024) and skill localization (Panigrahi et al., 2023). Song et al. (2024) showed that sparse tuning is an effective PEFT method, linking the low intrinsic dimensionality of pre-trained models to the proportion of parameters needing updates. They propose selecting optimal fine-tuning parameters based on gradient magnitudes. We enable sparse tuning for SSM by applying sparsity across entire dimensions (channel and state) rather than specific neurons. Panigrahi et al. (2023) focused on identifying neurons responsible for specific downstream tasks by fully fine-tuning the model and computing neuron masks to minimize task loss. While effective for skill localization, this method is computationally expensive and not optimized for parameter-efficient fine-tuning.
In Sec. A, we provide a more detailed discussion of related work on SSMs and PEFT.
3 Preliminaries
3.1 State Space Models
Discrete-Time SSMs.
[…]
To handle discrete inputs, the continuous parameters $({\bm{A}},{\bm{B}})$ are discretized into $(\overline{{\bm{A}}},\overline{{\bm{B}}})$ using a learnable step size $\Delta\in{\mathbb{R}}$. A common discretization rule, the zero-order hold, defines $\overline{{\bm{A}}}=\exp(\Delta{\bm{A}}),\overline{{\bm{B}}}=(\Delta{\bm{A}})^{-1}(\exp(\Delta{\bm{A}})-{\bm{I}})\cdot\Delta{\bm{B}}$. The discrete-time SSM, given in (2), enables efficient inference via long convolution described in (3). For multi-channel inputs $\bm{x},\bm{y}\in{\mathbb{R}}^{D}$, separate SSMs are used per channel, with a superscript $(d)$ indicating channel-specific parameters when needed.
Structured State Space Sequence Model (S4).
[…]
$\bm{y}_{t}=\operatorname{ReLU}({\bm{W}}\cdot\operatorname{S4}_{t}(\bm{x}_{1},\ldots,\bm{x}_{t})+\bm{\beta}+{\bm{u}}\odot\bm{x}_{t}),$ (4)
where ${\bm{W}}\in{\mathbb{R}}^{D\times D}$ and $\bm{\beta}\in{\mathbb{R}}^{D}$ represent the linear projection matrix and bias, respectively, and ${\bm{u}}\in{\mathbb{R}}^{D}$ is the coefficient of the residual connection. Trainable parameters include SSM parameters $({\bm{A}}^{(d)},{\bm{B}}^{(d)},{\bm{C}}^{(d)},\Delta^{(d)})$ across $D$ channels with ${\bm{A}}^{(d)}$ being diagonal, as well as linear layer $({\bm{W}},\bm{\beta})$ and residual connection ${\bm{u}}$.
Selective State Space Models (S6).
All SSMs mentioned above exhibit linear time invariance (LTI), meaning their dynamics remain constant over time. A key limitation of LTI SSMs is their fixed dynamics, hindering selective context extraction and input-dependent state transitions. S6 (Gu & Dao, 2024) addresses this by making parameters input-dependent. At each time step $t$, given input $\bm{x}_{t}\in{\mathbb{R}}^{D}$, S6 introduces input-dependent step sizes $\bm{\Delta}_{t}=(\Delta^{(1)}_{t},\ldots,\Delta^{(D)}_{t})^{\top}\in{\mathbb{R}}^{D}$, input transition vectors ${\bm{B}}_{t}\in{\mathbb{R}}^{H\times 1}$ and output mapping vectors ${\bm{C}}_{t}\in{\mathbb{R}}^{1\times H}$ via linear projection:
$\bm{\Delta}_{t}=\operatorname{softplus}(\bm{W}_{\bm{\Delta}}\bm{x}_{t}+\bm{\beta}_{\bm{\Delta}}),\quad{\bm{B}}_{t}={\bm{W}}_{{\bm{B}}}\bm{x}_{t},\quad{\bm{C}}_{t}={\bm{W}}_{{\bm{C}}}\bm{x}_{t},$
where the diagonal state matrices ${\bm{A}}^{(1)},\ldots,{\bm{A}}^{(D)}$ remain input-independent. The weight $\bm{W}_{\bm{\Delta}}\in{\mathbb{R}}^{D\times D}$ is factorized as $\bm{W}_{\bm{\Delta}}={\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}$, with ${\bm{W}}_{\bm{\Delta},\downarrow}\in{\mathbb{R}}^{D\times R}$, ${\bm{W}}_{\bm{\Delta},\uparrow}\in{\mathbb{R}}^{R\times D}$ to reduce computation (Wang et al., 2021, 2023a). Trainable parameters in S6 include ${\bm{A}}^{(d)}$ across $D$ channels, ${\bm{W}}_{\bm{\Delta},\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow}$ and $\bm{\beta}_{\bm{\Delta}}$ for computing $\bm{\Delta}_{t}$, and ${\bm{W}}_{{\bm{B}}},{\bm{W}}_{{\bm{C}}}\in{\mathbb{R}}^{H\times D}$ for computing ${\bm{B}}_{t},{\bm{C}}_{t}$. Discretization follows: $\overline{{\bm{A}}}_{t}^{(d)}=\exp(\Delta_{t}^{(d)}{\bm{A}}^{(d)}),\overline{{\bm{B}}}_{t}^{(d)}=\Delta_{t}^{(d)}{\bm{B}}_{t}$. Unlike S4, where ${\bm{B}}^{(d)}$ varies per channel, S6’s variation on $\overline{{\bm{B}}}^{(d)}$ stems from the scalar $\Delta_{t}^{(d)}$. Additionally, S6 shares ${\bm{C}}_{t}$ for all channels at each time step $t$, while S4 assigns a distinct ${\bm{C}}^{(d)}$ to each channel.
Mamba & Jamba.
Similar to the Transformer block, which consists of attention and linear layers, the Mamba-I block proposed by Gu & Dao (2024) features an S6 module, a point-wise 1D causal convolution layer (Conv1d) for token mixing, linear layers — including input ($\bm{W}_{\text{in}}$) and output ($\bm{W}_{\text{out}}$) projection layers and a gated MLP. Mamba-II (Dao & Gu, 2024) further simplifies the state matrix ${\bm{A}}$ to be a scalar. Building on Mamba-I, Jamba (Lieber et al., 2025) introduces a hybrid architecture that integrates both Transformer blocks and Mamba blocks, leveraging the strengths of both to enhance performance. This paper focuses on Mamba-I (referred as Mamba in this paper) and Jamba, deferring Mamba-II discussions to the appendix.
3.2 Parameter-Efficient Fine-Tuning
Input-Injection Methods.
Input-injection methods, such as prompt tuning (Lester et al., 2021) and prefix-tuning (Li & Liang, 2021), enhance the model’s input by injecting specialized sequences. Prompt tuning prepends a set of trainable embeddings ${\bm{P}}\in{\mathbb{R}}^{D\times M}$ to the original input ${\bm{X}}\in{\mathbb{R}}^{D\times N}$, forming the concatenated sequence $\widetilde{{\bm{X}}}=[{\bm{P}};{\bm{X}}]$. Prefix-tuning (Li & Liang, 2021) instead injects learnable vectors into the key and value matrices of each attention layer. For a Transformer layer, it prepends prefix states ${\bm{P}}^{K},{\bm{P}}^{V}\in{\mathbb{R}}^{L\times D}$ to the original projections:
$\widetilde{{\bm{K}}}=[{\bm{P}}^{K};{\bm{K}}],\quad\widetilde{{\bm{V}}}=[{\bm{P}}^{V};{\bm{V}}],$ (6)
where ${\bm{K}}$ and ${\bm{V}}$ are the key and value matrices derived from the input. We note that prefix-tuning is functionally equivalent to prepending soft tokens to the input at each attention layer and discarding the corresponding outputs associated with the prepended tokens. This view simplifies adaptation to SSMs, which lack explicit key and query projections. Yoshimura et al. (2025) also adopt this implementation, though they refer to it as affix-tuning.
Architecture-Enhancement Methods.
These methods modify the model’s internal structure to introduce tunable components. In the context of SSMs, one example is Additional-scan (Yoshimura et al., 2025), which expands the state dimensions within the SSM block and fine-tunes only the added parameters, leaving the original weights untouched.
Weight-Tuning Methods.
Notable weight-tuning methods include LoRA (Hu et al., 2021) and its variants (Liu et al., 2024; Hayou et al., 2024), as well as BitFit (Zaken et al., 2022). LoRA (Hu et al., 2021) fine-tunes a model by introducing low-rank updates to its weight matrices. Given a weight matrix ${\bm{W}}_{0}\in{\mathbb{R}}^{D\times D}$, LoRA updates it as follows:
${\bm{W}}={\bm{W}}_{0}+{\bm{W}}_{\downarrow}{\bm{W}}_{\uparrow},$ (7)
with ${\bm{W}}_{\downarrow}\in{\mathbb{R}}^{D\times R}$, ${\bm{W}}_{\uparrow}\in{\mathbb{R}}^{R\times D}$, and $R\ll D$ being the rank. Only ${\bm{W}}_{\downarrow}$ and ${\bm{W}}_{\uparrow}$ are trained, reducing the number of trainable parameters from $D^{2}$ to $2RD$. Weight-Decomposed Low-Rank Adaptation (DoRA) (Liu et al., 2024) improves upon LoRA by decomposing the weight matrix into two components: magnitude (${\bm{m}}\in{\mathbb{R}}^{D}$) and direction (${\bm{W}}_{\downarrow}{\bm{W}}_{\uparrow}$), leading to the formulation
${\bm{W}}={\bm{m}}\frac{{\bm{W}}_{0}+{\bm{W}}_{\downarrow}{\bm{W}}_{\uparrow}}{\lVert{\bm{W}}_{0}+{\bm{W}}_{\downarrow}{\bm{W}}_{\uparrow}\rVert}.$ (8)
This additional parameter ${\bm{m}}$ enhances both training capacity and stability. LoRA+ (Hayou et al., 2024) modifies LoRA by applying different learning rates to ${\bm{W}}_{\downarrow}$ and ${\bm{W}}_{\uparrow}$, enabling more effective feature learning. In contrast, BitFit (Zaken et al., 2022) updates only the bias terms, offering a lightweight and highly parameter-efficient alternative.
4 Benchmarking PEFT Methods on SSM-based Models
In this section, we examine the effectiveness of popular PEFT methods when applied naively to SSM-based models, specifically Mamba and Jamba.
4.1 Experiment Setup
We evaluate PEFT methods across three categories: input-injection, architecture-enhancement, and weight-tuning. For input-injection methods, we use prompt tuning (Lester et al., 2021) and prefix-tuning (Li & Liang, 2021), where prefix-tuning employs an overparameterized MLP for stable optimization. For architecture-enhancement methods, we include Additional-scan (Yoshimura et al., 2025), which introduces and fine-tunes newly added state dimensions in SSM modules. For weight-tuning, we consider BitFit (Zaken et al., 2022) and LoRA, including LoRA (Hu et al., 2021) and DoRA (Liu et al., 2024), while LoRA$+$ (Hayou et al., 2024) is deferred to Sec. E.2. BitFit fine-tunes the bias terms of Conv1d and ${\bm{W}}_{\bm{\Delta},\uparrow}$.
We use six datasets spanning different domains: GLUE for natural language understanding (Wang et al., 2019), DART for RDF-to-text generation (Nan et al., 2021), SAMSum (Gliwa et al., 2019) for summarization, Spider for text-to-SQL generation (Yu et al., 2018), and two vision datasets—CIFAR-10 (Krizhevsky et al., 2009) and CelebA (Liu et al., 2015), with the vision datasets processed by cropping, resizing, and flattening pixel values into space-separated numerical sentences. Details are in Sec. B. Prefix-tuning requires significantly more parameters than other PEFT methods due to its per-layer MLP for projecting fixed sequences into soft tokens. For all methods—except prefix-tuning and the special case of LoRA and DoRA when applied to both linear projection layers—we limit trainable parameters to below 1% for Mamba and below 0.15% for Jamba. For Jamba, all PEFT methods are applied to Mamba layers, while Transformer layers remain frozen to isolate performance effects. See more details in Sec. C.1.
4.2 Results
Table 1 summarizes the benchmarking results. Detailed results for GLUE and Spider subtasks appear in Sec. C.2. We analyze the results from three key perspectives below.
Model Method Major Target Module GLUE DART SAMSum Spider CIFAR-10 CelebA
Avg. Score METEOR BLEU R1 R2 RL Acc. Acc. Acc.
Mamba Prompt Tuning Other 63.8 66.2 39.8 50.1 25.6 41.6 43.6 30.4 82.5
Prefix-Tuning SSM 68.6 66.6 42.5 50.6 26.5 42.1 39.7 41.0 86.5
BitFit Both 76.8 67.0 43.7 50.3 25.7 41.9 48.4 44.4 86.9
LoRA SSM 76.9 68.8 48.0 50.4 26.0 41.8 55.0 52.3 87.0
LinProj 81.2 70.9 49.5 50.9 27.0 42.3 57.5 61.0 87.0
Both 80.3 70.2 52.2 50.7 26.8 42.4 57.0 58.4 89.8
DoRA SSM 77.9 68.3 47.3 48.1 24.2 39.6 55.3 44.5 87.1
LinProj 81.1 70.7 51.6 51.0 26.9 42.8 60.7 57.6 86.7
Both 80.8 70.8 51.4 51.3 27.2 43.0 58.1 58.2 89.8
Additional-Scan SSM 62.4 60.6 15.8 37.6 17.5 30.9 26.9 32.2 86.0
Full Fine-Tuning Both 80.5 71.0 51.8 51.2 27.3 42.9 66.2 60.0 89.4
Jamba Prompt Tuning Other 73.3 54.1 6.3 54.7 31.8 46.8 74.9 40.9 85.6
Prefix-Tuning SSM 56.9 59.6 14.4 11.5 1.8 10.4 0.3 29.9 82.2
BitFit Other 75.2 59.2 14.8 54.7 31.9 47.0 73.7 45.6 86.3
LoRA LinProj 73.9 68.9 37.8 54.6 32.3 46.8 69.3 59.7 89.0
DoRA LinProj 71.4 68.1 28.8 55.2 32.2 47.3 70.9 58.6 89.0
Additional-Scan SSM 68.3 63.3 20.1 53.4 30.5 45.6 69.3 50.6 0.0
Table 1: Benchmarking popular Parameter-Efficient Fine-Tuning (PEFT) methods on Mamba (Gu & Dao, 2024) and Jamba (Lieber et al., 2025) across six real-world datasets. R1/R2/RL stand for ROUGE-1/2/L. We evaluate PEFT applied to different target modules: SSM module only, linear projection matrices (LinProj) only, both, or other components such as embedding layer. For both Mamba and Jamba, all methods use fewer than 1% and 0.15% of parameters, respectively, except when the target module for LoRA or DoRA is set to “Both” or when prefix-tuning is applied. Comprehensive hyperparameter tuning was performed for all methods. Bold values indicate the best performance for each model (Mamba and Jamba) separately, while underlined values denote the second-best performance for each task, excluding full fine-tuning. Key findings include: (i) among PEFT methods applied to SSM modules, LoRA outperforms others, (ii) for all PEFT methods, LoRA achieves the best performance, (iii) applying LoRA to linear projections yields results comparable to applying it to both linear projections and SSM modules, while outperforming its application solely to SSM modules, and (iv) input-injection methods (i.e., prompt tuning and prefix-tuning), are generally ineffective.
Superiority of LoRA.
The most prominent finding is that LoRA consistently outperforms other PEFT methods (e.g., prompt tuning, prefix-tuning, BitFit, additional-scan), regardless of the target module.
Finding:
Across all target modules, LoRA surpasses existing PEFT methods in performance.
Even when restricted to SSM modules, LoRA still outperforms all other PEFT baselines applied to the same target.
Limitations of Input-Injection Methods.
Input-injection methods like prefix-tuning are ineffective for SSM-based models (Table 1), as their expressiveness reduces to tuning only the initial hidden state (Proposition 1). Formal statement, proof and empirical verification are in Section C.3.
Optimal Application of LoRA in SSM-based Models.
Table 1 shows that LoRA outperforms all other PEFT methods in most scenarios. From our results, we explore the optimal layers for applying LoRA in SSM-based models: the SSM module, the linear projection matrices, or a combination of both. Note that S6 in Mamba and Jamba includes fine-grained parameters like x_proj ($\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$) and dt_proj (${\bm{W}}_{\bm{\Delta},\uparrow}$), which were already explored by Yoshimura et al. (2025) on Mamba. We defer a deeper discussion of them to Sec. C.4 and focus on the key question here: Is applying LoRA to SSM modules necessary for performance gains? By narrowing our scope, we aim to clarify LoRA’s impact across major components (e.g., SSM modules, linear projection matrices) rather than all specific parameters.
We evaluate LoRA’s performance on linear projections using $\bm{W}_{\text{in}}$, $\bm{W}_{\text{out}}$, and both combined. Since the performance of different combinations of linear projections is consistent across datasets (see Sec. C.4.), we only report the results for LoRA applied to $\bm{W}_{\text{in}}$ in Table 1. For SSM modules, we apply LoRA to weight matrices, including those for the input-dependent step size $\bm{\Delta}$. For state transition matrices ${\bm{A}}$, we treat their diagonal structures as vectors, concatenate them across channels to form a matrix, and apply LoRA. Table 1 summarizes results for the best-performing configurations (see Section C.2 for full results). Based on these results, we present the following finding:
Finding:
For LoRA: Tuning on SSMs is less effective than tuning linear projection matrices, with the latter performing comparably to tuning both.
Detailed experiments, including LoRA on different linear projection matrices and additional evaluations of Mamba-II, are presented in Sec. C.2. These experiments reinforce the finding that LoRA is highly effective for linear projections but less suitable for SSM modules.
To further elucidate this concept, we present the following lemma, which examines a simplified model architecture consisting of S6 with two linear input projection matrices at each layer. We demonstrate that fine-tuning one input projection matrix encompasses the expressivity of fine-tuning the parameters $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\uparrow}$. Consider an S6 model with two input projection matrices ${\bm{W}}_{\text{in,1}},{\bm{W}}_{\text{in,2}}\in{\mathbb{R}}^{D\times D}$: the first affects how internal parameters depend on the input, while the second governs the input passed directly into the S6 module. Under this setup, the output $y^{(d)}_{N}$ can be expressed as:
$&y^{(d)}_{N}=\underbrace{\overbrace{{\bm{C}}({\bm{W}}_{\text{in,1}}{\bm{x}}_{N})^{\top}}^{\text{Input-dependent }{\bm{C}}_{N}}\sum_{n=1}^{N}\left(\prod_{m=1}^{n}\overbrace{\overline{{\bm{A}}}({\bm{W}}_{\text{in,1}}{\bm{x}}_{m})}^{\text{Input-dependent }\overline{{\bm{A}}}_{m}}\right)\overbrace{\overline{{\bm{B}}}_{n}({\bm{W}}_{\text{in,1}}{\bm{x}}_{n})}^{\text{Input-dependent }\overline{{\bm{B}}}_{n}}}_{\text{Parameters depending on input after projection }{\bm{W}}_{\text{in,1}}}\underbrace{({\bm{W}}_{\text{in,2}}{\bm{x}}_{n})^{(d)}}_{\text{Input after projection}{\bm{W}}_{\text{in,2}}}.\end{aligned}$
When ${\bm{W}}_{\text{in,1}}={\bm{W}}_{\text{in,2}}$, this reduces to a standard architecture with a single input projection followed by an S6 layer. For simplicity, we let $\bm{\beta}_{\bm{\Delta}}={\bm{0}}.$ Then the full model is parameterized by ($(\{{\bm{A}}^{(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\text{in,1}},{\bm{W}}_{\text{in,2}})$. Assume none of the parameters are zero and $D>2H+R$, where $R$ is the rank of ${\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}$.
Lemma 1 (Expressivity of Fine-Tuning Projection Matrices).
Consider two models with the architecture described above. Let:
  • A target model $f^{\star}$ parameterized by ($(\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, $\bm{W}_{{\bm{B}}}^{\star}$, $\bm{W}_{{\bm{C}}}^{\star}$, ${\bm{W}}_{\bm{\Delta},\uparrow}^{\star}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, ${\bm{W}}_{\text{in,1}}^{\star}$, ${\bm{W}}_{\text{in,2}}^{\star})$;
  • A frozen model $f_{0}$ parameterized by ($(\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, ${\bm{W}}_{\text{in,1}}$, ${\bm{W}}_{\text{in,2}}^{\star})$.
The two models share $\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, and ${\bm{W}}_{\text{in,2}}^{\star}$, while differing in $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$, and ${\bm{W}}_{\text{in,1}}$. Then, there exists an updated projection matrix $\widehat{{\bm{W}}}_{\text{in,1}}$ such that the frozen model matches the output of the target model without updating $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$ for any input sequence, i.e.,
$&f(\cdot;\{{\bm{A}}^{\star(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow}^{\star},\widehat{{\bm{W}}_{\text{in,1}}},{\bm{W}}_{\text{in,2}}^{\star})\\ &=f^{\star}(\cdot;\{{\bm{A}}^{\star(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}}^{\star},\bm{W}_{{\bm{C}}}^{\star},{\bm{W}}_{\bm{\Delta},\uparrow}^{\star},{\bm{W}}_{\bm{\Delta},\downarrow}^{\star},{\bm{W}}_{\text{in,1}}^{\star},{\bm{W}}_{\text{in,2}}^{\star}).\end{aligned}$
We expand on this discussion in Sec. C.4, where we present both theoretical proofs and empirical validation. The lemma shows that tuning the linear projection matrix can match the expressive power of certain SSM parameters (i.e., $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\uparrow}$), aligning with our empirical observation that tuning only the linear projections already performs well. However, a key limitation of tuning only the linear projection matrices remains: such tuning lacks the expressive power to affect the state matrix ${\bm{A}}$, which is an essential parameter for sequence-to-sequence operations. Therefore, tuning the SSM modules is still necessary. Existing PEFT methods fall short in effectively tuning SSM modules: (i) alternative methods underperform compared to LoRA on SSM modules, and (ii) applying LoRA to SSM modules does not improve performance beyond applying it to linear projections alone. These findings highlight a gap in current PEFT techniques for SSM modules, leading to an importantca question: Is there a more effective strategy for fine-tuning SSM modules?
5 Sparse Dimension Tuning
This section aims to develop an algorithm for tuning SSM modules. In doing so, we start by first analyzing the roles of different parameters, as outlined in Lemma 2. This analysis motivates us to classify channels and state dimensions into three categories: (i) zero, (ii) trainable, and (iii) frozen, leading to the development of the Sparse Dimension Tuning and Pruning (SDT-P) method. We then establish theoretical guarantees for applying SDT-P to SSM modules and LoRA to linear projection matrices (Theorem 1). Finally, we simplify SDT-P into Sparse Dimension Tuning (SDT) by omitting pruning, as pruned parameters can be effectively considered as being trained to zero. This simplified version serves as the primary method used in our experiments.
5.1 Understanding Key Parameters in S4 Modules
Problem Setting.
Inspired by the work by Zeng & Lee (2024), we analyze the expressive power of S4 parameters using a similar framework. We assume a well-performing target model and a frozen model (pretrained or random) and aim to update the frozen model efficiently to match the target. Following Zeng & Lee (2024), we assume the frozen model has a capacity at least as large as the target model. This assumption ensures analytical tractability and is reasonable, as frozen models are typically overparameterized in practice. Both models are S4 with hidden dimensions $H_{\star}$ (target) and $H\geq H_{\star}$ (frozen). Assuming all hidden dimensions are active (i.e., all parameters are non-zero), we define their dynamics using discretized parameters $(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})$:
(Target model) $f^{\star}({\bm{x}})_{n}=\sum\nolimits_{m=1}^{n}{\bm{C}}_{\star}\overline{{\bm{A}}}_{\star}^{m-n}\overline{{\bm{B}}}_{\star}x_{m},$ (11)
(Frozen model) $f_{0}({\bm{x}})_{n}=\sum\nolimits_{m=1}^{n}{\bm{C}}_{0}\overline{{\bm{A}}}_{0}^{m-n}\overline{{\bm{B}}}_{0}x_{m},$ (12)
where $\operatorname{diag}(\overline{{\bm{A}}}_{\star}),\overline{{\bm{B}}}_{\star},{\bm{C}}_{\star}\in{\mathbb{R}}^{H_{\star}}$, $\operatorname{diag}(\overline{{\bm{A}}}_{0}),\overline{{\bm{B}}}_{0},{\bm{C}}_{0}\in{\mathbb{R}}^{H}$. This formulation shows that the S4 module remains unchanged even if the state dimensions are permuted.
Parameter Efficiency Analysis on S4.
We analyze the parameter efficiency of updating a frozen S4 module after discretizing its parameters $(\overline{{\bm{A}}}_{0},\overline{{\bm{B}}}_{0},{\bm{C}}_{0})$ to match the functionality of a target S4 module with discretized parameters $(\overline{{\bm{A}}}_{\star},\overline{{\bm{B}}}_{\star},{\bm{C}}_{\star})$. Based on this setup, we present the following result characterizing the minimum number of parameters that must be tuned for functional equivalence.
Lemma 2 (Minimal Parameter Adjustment for S4 Fine-Tuning).
Assume all hidden dimensions of the target model $f^{\star}$ are non-zero, i.e., all elements of $\operatorname{diag}(\overline{{\bm{A}}}_{\star})\odot\overline{{\bm{B}}}_{\star}\odot{\bm{C}}_{\star}$ are non-zero. To update frozen model $f_{0}$ such that it becomes functionally equivalent to the target model $f^{\star}$, the minimum number of tunable parameters is:
$&\min_{\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}}\overbrace{\left\lVert\left[\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}\odot{\bm{C}}^{\top}\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}\\ &+\overbrace{\left\lVert\left[\overline{{\bm{A}}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}+\left\lVert\left[\overline{{\bm{B}}}\odot{\bm{C}}^{\top}\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}}^{\text{ aligning remaining dimensions with target model}},\\ \end{aligned}$
subject to
$(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})\in\{({\bm{P}}^{\top}\overline{{\bm{A}}}_{0}{\bm{P}},{\bm{P}}^{\top}\overline{{\bm{B}}}_{0},{\bm{C}}_{0}{\bm{P}}):{\bm{P}}\text{ is a permutation matrix}\}.\end{aligned}$
Note that the search space consists of all possible S4 parameterizations that can be obtained by permuting the hidden dimensions of the frozen model. Proofs and further details are provided in Sec. D.1. This result highlights three distinct roles of the state dimensions. First, any dimensions that do not contribute to the target function (represented by the first term in (13)) are effectively zero and can be pruned. These correspond to state dimensions larger than those of the target model after permutation, indicating that redundant information can be directly removed to eliminate its impact. Second, among the remaining dimensions, alignment is necessary for those that do not already match the target. The state matrix ${\bm{A}}$ plays a crucial role in sequence modeling by capturing dependencies between tokens at different positions. To achieve functional equivalence (as represented by the second term in (13)), ${\bm{A}}$ must be aligned. Notably, dimensions that are already aligned with the target require no updates. These two insights motivate our Sparse Dimension Tuning and Pruning (SDT-P) method, which classifies hidden dimensions into three categories: (i) zero, (ii) frozen (already aligned), and (iii) trainable. Finally, the third term in (13) indicates that the expressive power of $\overline{{\bm{B}}}$ and ${\bm{C}}$ is essentially equivalent, meaning that tuning either one is equivalent to updating both.
5.2 Sparse Dimension Tuning and Pruning (SDT-P)
Building on Lemma 2, we introduce SDT-P, the precursor to Sparse Dimension Tuning (SDT). SDT-P updates parameters selectively based on the role of each state dimension. In the multi-channel case, we first categorize the channel dimensions into three groups: pruned, frozen, and trainable. Then, the state dimensions of each trainable channel are also categorized as pruned, frozen, or trainable. This hierarchical selection ensures that updates are applied only when necessary, while pruned dimensions are discarded and frozen dimensions remain unchanged.
Dimension Selection Algorithm.
To enable this structured tuning process, we first introduce our dimension selection algorithm. The algorithm starts with a warmup epoch, where the SSM modules are updated using a subset of the dataset for one epoch. After this warmup, we classify channel dimensions based on the magnitude of the state matrix ${\bm{A}}$: dimensions with small magnitude are pruned (set to zero), those with significant changes are marked as trainable, and the rest remain frozen. Next, we apply the same classification to state dimensions, but only within the trainable channels. The detailed pseudo-code is in Sec. D.4.
Parameter Update Scheme.
Once the channel and state dimensions are selected, we determine how to update the parameters. (S4) For S4, Gu et al. (2022a) showed that tuning ${\bm{C}}$ alone is as effective as tuning both $\overline{{\bm{B}}}$ and ${\bm{C}}$. Therefore, we always freeze $\overline{{\bm{B}}}$ and update only $\overline{{\bm{A}}}$ and ${\bm{C}}$. Specifically, an entry in $\overline{{\bm{A}}}$ or ${\bm{C}}$ is trainable if and only if both its channel and state dimensions are trainable. If either the channel or state dimension is pruned, the entry is pruned as well. All other entries remain frozen. (S6) For S6, where parameters are input-dependent, we update $\overline{{\bm{A}}},\bm{W}_{{\bm{B}}}$, and $\bm{W}_{{\bm{C}}}$ instead. Since $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ operate across channels, we categorize their updates based only on channel dimensions—we do not update individual state dimensions differently for each channel. Based on this categorization, we mark the corresponding columns of $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ as trainable, frozen, or pruned accordingly.
The dimension selection algorithm and parameter updates together form the SDT-P method for tuning SSM modules. Next, we provide theoretical guarantees for applying SDT-P to SSM modules and LoRA to linear projection matrices.
5.3 Expressive Power of SDT-P Combined with LoRA
Our analysis focuses on simplified SSM-based models, where each layer consists of an SSM module followed by linear projection matrices with residual connections. We refer to this structure as a deep SSM layer: i) a deep S4 layer consists of an S4 module followed by linear projections; ii) a deep S6 layer follows the same structure but replace S4 with S6. A deep S4 model is composed of deep S4 layers, while a deep S6 model consists of deep S6 layers. The detailed formulation of deep S4 layers is provided in Sec. 3, and a deep S6 layer follows the same structure with S4 replaced by S6. The following theorem highlights the expressive power of SDT-P on updating SSM modules, where each layer uses a single type of SSM module (S4 or S6) followed by linear projections.
Theorem 1 (Expressive Power of SDT-P with LoRA on Simplified SSM-based Models).
Assume all layers use linear activations. Let $f_{0}$ be a frozen deep S4 or S6 model with $L$ layers, each containing $H$ hidden states per channel. Let $f^{\star}$ be a smaller target model of the same type (S4 or S6), with no residual connections, $L^{\star}<L$ layers, and $H^{\star}<H$ hidden states per channel. Then, there exists a set of parameter updates to $f_{0}$ satisfying the following conditions such that for any finite-length input sequence ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{N})$ with ${\bm{x}}_{n}\in{\mathcal{X}}\subset{\mathbb{R}}^{D}$, where ${\mathcal{X}}$ is bounded, the resulting model $f$ satisfies $f({\bm{X}})=f^{\star}({\bm{X}})$:
  1. 1.
    (SDT-P on SSM) In each SSM module, update at most $\lceil DL^{\star}/L\rceil$ channels. Within each updated channel, fine-tune at most $H^{\star}$ hidden states and set the rest to zero.
  2. 2.
    (LoRA on Linear Projections) Apply rank-$\lceil L/L^{\star}\rceil$ updates to each linear projection matrix.
  3. 3.
    (Minimal Additional Updates) Update only the residual connections, per-layer biases, and the final-layer output projection matrix.
For proof and details, refer to Sec. D.2 and D.3. This theorem shows that a larger pretrained model can be fine-tuned into any smaller model of the same architecture by applying SDT-P to SSM modules and LoRA to linear projection matrices. Moreover, for less complex tasks, where the target model has fewer layers ($L^{\star}$) and hidden states ($H^{\star}$), the required number of trainable channels and hidden states also decreases. This aligns with the theoretical analysis of LoRA by Zeng & Lee (2024), which demonstrates that larger pre-trained models require fewer learnable parameters (i.e., a lower-rank update) during fine-tuning, especially for simpler tasks. While our theorem assumes linear activations, no residual connections in the target model, and full fine-tuning of the last-layer projection matrix, our findings have broader implications. As our experimental results in Sec. 6 will show, these insights generalize beyond these theoretical constraints.
Algorithm 1 Dimension Selection Algorithm of SDT
Input: A small subset of dataset ${\mathcal{D}}$, warmup epochs $E$, number of layers $L$, total channels $D$, total states $H$, channel freeze ratio $\alpha$, state freeze ratio $\beta$
/* Warmup epochs */
Perform full update on SSM modules using ${\mathcal{D}}$ for $E$ epochs  for $l=1$ to $L$ do
       /* Unfreeze dimensions */ Sort channels ${\mathbb{D}}$ based on changes of $\|\overline{{\bm{A}}}^{(d)}\|$  Freeze the bottom $\beta|{\mathbb{D}}|$ channels, denoted by ${\mathbb{D}}^{\prime}$for $d\in{\mathbb{D}}^{\prime}$ do
             Sort state dimensions by the changes in $\|\overline{{\bm{A}}}^{(d)}\|$  Freeze the bottom $\alpha|{\mathbb{H}}|$ state dimensions at the $d$-th channel
5.4 Sparse Dimension Tuning (SDT): A Pruning-Free Alternative
While SDT-P classifies channels and states into three categories, we simplify our approach by omitting pruning and categorizing parameters as either trainable or frozen. We refer to this simplified method as Sparse Dimension Tuning (SDT). This reduces the number of hyperparameters, as pruned parameters are effectively equivalent to being trained to zero. The resulting dimension selection approach is outlined in the pseudo-code (Alg. 1), which corresponds to the update scheme illustrated in Fig. 1. Experiments will show that this simplification remains effective.
Overhead Analysis.
We assess the computational overhead of applying SDT with LoRA (for linear projection matrices) versus LoRA alone with Table 2 summarizing the results. Although SDT involves an additional dimension selection stage, Table 2 shows that this incurs minimal extra cost. Furthermore, with the same parameter budget, SDT for SSM modules combined with LoRA on linear projections runs faster than LoRA alone, since LoRA introduces extra matrix multiplications between two low-rank matrices for the SSM modules, whereas SDT does not. In Sec. D.6, we detail the experimental settings and present a memory usage analysis showing that SDT also consumes less memory during fine-tuning for the same reason.
Stage Method Mamba-130M Mamba-1.4B Jamba-Mini-52B
Dim. Selection LoRA & SDT 16.5 $\pm$ 3.9 85.8 $\pm$ 5.3 163.9 $\pm$ 10.2
Training (per epoch) LoRA 410.0 $\pm$ 80.0 2060.0 $\pm$ 135.0 3427.5 $\pm$ 185.0
LoRA & SDT 330.0 $\pm$ 77.5 1697.5 $\pm$ 87.5 3065.0 $\pm$ 232.5
Table 2: PEFT combining SDT with LoRA is more efficient than LoRA alone when the same number of trainable parameters are used. Shown are dimension selection and per-epoch training times (s) for Mamba and Jamba models.
6 Experimental Studies of SDT
In this section, we evaluate the performance of SDT in tuning SSM modules, comparing it to LoRA, the best existing PEFT method for fine-tuning SSM modules, as shown in Sec. 4. Our experiments reveal the key result:
Finding: SDT outperforms LoRA on SSM modules.
6.1 Synthetic Experiments on Deep S4 Models
This experiment validates our theoretical guarantees under broader conditions, including residual connections and ReLU activations in both models, without fully fine-tuning the last-layer projection matrix. See Sec. E.1 for details.
(Experimental Setup) We employ a regression setting to validate our theoretical results. We randomly initialize two models: a one-layer deep S4 model as the target and a four-layer deep S4 model as the frozen model. LoRA is applied to linear projection matrices, while different methods are tested on the SSM module to assess their effectiveness. The goal is to update the frozen model to match the target model’s functionality.
Figure 2: SDT surpasses LoRA in tuning S4 within deep S4 models when LoRA is applied to linear projection matrices in synthetic experiments.
We generate an input sequence ${\bm{X}}$ of length $200$ and dimension $64$, with values uniformly drawn from integers between 0 and 9. This input is then processed through the target model to obtain the corresponding outputs. These input-output pairs are used to train the frozen model over 500 iterations using the Mean Squared Error (MSE) loss. (Results) Figure 2 shows the MSE, averaged across all tokens, plotted against the number of trainable parameters for different methods on SSM modules. SDT achieves significantly lower MSE than LoRA on SSM modules, demonstrating its effectiveness in updating SSM modules.
6.2 Real-World Experiments on Pretrained Models
Lastly, we conduct experiments to evaluate our approach on pretrained models, including Mamba and Jamba with different model sizes. We consider five datasets: GLUE, DART, SAMSum, Spider, and CelebA. For these experiments, we split the datasets into three parts: train, validation, and test, different from benchmarking experiments. We combine our proposed SDT with LoRA and evaluate it in three different settings against three pure LoRA settings. In SDT, 99% of channels are frozen, and we adjust state freeze ratios. For the pure LoRA settings, we apply LoRA to different parameter sets, selecting ranks to ensure all settings have a comparable parameter budget for fair comparison. Residual connections and biases are frozen and learning rates are independently selected via a small grid search over data subsets. See Sec. E.2 for further details.
Mamba.
The experimental results of Mamba are reported in Table 3, showing that applying SDT on SSM modules outperforms pure LoRA, even when 99% of the channels are frozen. This underscores the effectiveness of SDT on fine-tuning SSM modules.
LinProj S6 GLUE DART CelebA SAMSum Spider
Avg. BLEU MET. Acc. R1 R2 RL Acc.
LoRA LoRA 80.8 51.0 70.2 88.6 51.6 28.2 43.2 83.5
SDT 81.1 51.5 70.5 88.6 51.7 28.1 43.4 84.5
DoRA DoRA 80.1 51.2 70.4 88.4 51.8 28.0 43.4 83.8
SDT 78.2 51.5 70.8 88.6 52.1 28.3 43.7 85.1
Table 3: Performance comparison between SDT and LoRA on pretrained Mamba models. Bold numbers indicate the best performance for each task. We use Mamba-130M to compare the performance of SDT and LoRA on GLUE (Wang et al., 2019), DART (Nan et al., 2021), and CelebA (Liu et al., 2015) benchmarks. For all other datasets, we employ Mamba-1.4B. We report only the best setting out of three for each method. We observe that SDT outperforms LoRA on updating SSM modules on Mamba.
Jamba.
We extend our experiments to Jamba, applying all tested methods exclusively to its Mamba layers. Notably, the performance gain on Jamba is smaller compared to Mamba. This is because we freeze all Transformer layers to isolate the effect of Mamba layers for a fair evaluation. Additionally, since the Mamba layers in Jamba contain significantly fewer parameters than those in the Mamba model, fine-tuning them yields limited performance improvements. Nevertheless, results on GLUE (Table 4) validate the effectiveness of our method. See Table 22 for more results.
LinProj S6 RTE MRPC CoLA SST-2 QNLI QQP MNLI Avg.
DoRA DoRA 65.7 77.8 7.1 93.9 77.8 67.8 85.4 67.9
SDT 67.1 77.5 7.5 94.2 79.6 72.7 85.5 69.2
Table 4: Performance comparison between SDT and DoRA on pretrained Jamba models. Bold numbers indicate the best performance for each task. We use Jamba-Tiny-319M to compare the performance of SDT and DoRA on the GLUE (Wang et al., 2019) benchmark. We report only the best setting out of three for each method. We observe that SDT outperforms DoRA on updating SSM modules on Jamba.
7 Discussion
In this paper, we study PEFT methods applied to SSM-based models. Our evaluation of existing PEFT methods provides valuable insights and guidelines for future researchers to parameter-efficiently fine-tune SSM-based models for other domains. Moreover, we take an initial step in establishing a theoretical framework for studying PEFT methods on SSM-based models. Additionally, we introduce SDT, a PEFT method specifically tailored to SSM modules, demonstrating superior performance compared to existing approaches.
Limitations & Future Works.
The theoretical guarantees for SDT are restricted to linear activations and require full fine-tuning of the last layer. Nonetheless, our experiments show that SDT performs well in practice despite these constraints. Addressing these theoretical limitations or developing new PEFT methods applicable to broader scenarios is a promising future direction. Additionally, our theory shows that modifying a subset of channels and states is sufficient but does not guide optimal selection. Our approach, based on a warmup stage and parameter magnitude, might not be optimal. Future research could explore the impact of channel/state selection and improve dimension selection algorithms.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgment
The works is supported by NSF Award DMS-2023239, NSF CAREER Award CCF-2339978, Amazon Research Award, and a grant from FuriosaAI.
Appendix
[…]
Appendix A Additional Related Works
A.1 Additional Related Works on SSMs
Linear State-Space Layers (LSSL) represent one of the earliest SSM layers utilized in deep learning, functioning as continuous-time, recurrent, and convolutional models (Gu et al., 2021). LSSL employs HiPPO theory (Gu et al., 2020) to initialize the state matrix ${\bm{A}}$, enabling the capture of long dependencies. However, LSSL is computationally expensive, limiting its practical application. Gu et al. (2022b) introduced Structured State Space Models (S4), which optimize computation efficiency by employing a structured state matrix ${\bm{A}}$. Gupta et al. (2022) proposed DSS, which simplifies the model by using a diagonal matrix for ${\bm{A}}$ and empirically demonstrated that it suffices to achieve performance comparable to S4. Further, Gu et al. (2022a) provided a theoretical explanation for the effectiveness of the diagonal state matrix ${\bm{A}}$ in DSS and introduced S4D, which offers various initialization methods for ${\bm{A}}$. Subsequently, the diagonal structure of the state matrix ${\bm{A}}$ has been adopted in follow-up methods (Gu & Dao, 2024). Despite differences in optimization algorithms, we refer to S4 and its close variants, including DSS and S4D, collectively as S4. This terminology encompasses models that maintain the standard discrete-time SSM form with a diagonal state matrix.
Despite of the remarkable performance of SSMs on certain tasks of sequence modeling, SSMs still showed worse performance than Transformers on language modeling. Fu et al. (2022) transitioned from synthetic language modeling tasks to real language modeling tasks with SSMs. They proposed H3, which is inspired by Linear Attention (Katharopoulos et al., 2020), introducing both diagonal SSM and shift SSM. Recently, Mamba (Gu & Dao, 2024; Dao & Gu, 2024) escaped from linear time invariance (LTI) modeling by introducing input-dependent terms and achieved better performance than Transformers on language modeling. Furthermore, several hybrid models (Lieber et al., 2025; Park et al., 2024) tried to exploit the advantages of both SSMs and Transformers.
A.2 Additional Related Works on PEFT
In this section, we provide a more detailed description of the baseline methods.
LoRA (Hu et al., 2021).
LoRA (Low-Rank Adaptation) focuses on fine-tuning large models by freezing pretrained parameters and injecting trainable low-rank matrices into each layer of the Transformer architecture. The intuition behind using low-rank matrices comes from linear algebra, where a large matrix can be closely approximated by the product of two smaller matrices. The number of trainable parameters can be controlled with the rank of the low-rank matrices. LoRA also uses a scaling parameter (LoRA alpha) for the weight matrices to control the balance of the original model weights and LoRA weights during training. After fine-tuning, LoRA weights can be merged with the original model weights, introducing no additional inference overhead.
Prompt Tuning (Lester et al., 2021).
Prompt tuning freezes all model weights and prepends a trainable soft prompt to the input prompt. The soft prompt consists of trainable virtual tokens, which are continuous. At inference time, prompt tuning introduces an inference overhead based on the number of virtual tokens used.
Prefix-Tuning (Li & Liang, 2021).
Prefix-tuning also prepends trainable tokens to the input like prompt tuning but injects separate prefixes in every layer. For each Transformer layer, prefix-tuning prepends trainable embeddings to the attention’s ${\bm{K}}$ and ${\bm{V}}$ matrix. The authors have found that directly training these prefixes can lead to unstable training, so they propose to over-parameterize them with a large MLP to increase training stability. After training, the MLP can be dropped. Like prompt tuning, prefix-tuning introduces an inference overhead, scaling with the number of trainable embeddings.
BitFit (Zaken et al., 2022).
BitFit is a simple but effective PEFT method that freezes all model weights except the bias terms, consequently greatly reducing the number of trainable parameters. As no additional parameters are added, no inference overhead occurs.
Theoretical Understanding of PEFT.
Numerous efforts have been made to theoretically understand existing PEFT methods. For input-injection methods, Wang et al. (2023b), Petrov et al. (2024), and Oymak et al. (2023) have theoretically analyzed the effectiveness and limitations of prompt tuning and prefix-tuning for Transformer-based models. For LoRA, Zeng & Lee (2024) explored its expressive power by demonstrating that even a randomly initialized model can be adapted to match any smaller target model using LoRA. Some of our theoretical analysis draws upon the framework established by Zeng & Lee (2024). Jang et al. (2024) conducted a theoretical exploration of LoRA within the neural tangent kernel (NTK) regime.
Appendix B Details of Datasets
In this paper, we consider six datasets across three domains: (i) Natural Language Understanding (NLU), represented by GLUE (Wang et al., 2019); (ii) Natural Language Generation (NLG), including SAMSum (Gliwa et al., 2019), Spider (Yu et al., 2018) and DART (Nan et al., 2021); and (iii) Computer Vision (CV), represented by CIFAR-10 (Krizhevsky et al., 2009) and CelebA (Liu et al., 2015).
GLUE (Wang et al., 2019).
The GLUE (General Language Understanding Evaluation) benchmark is a collection of datasets used for training, evaluating, and analyzing natural language understanding models across a range of diverse tasks. The benchmark includes nine sentence- or sentence-pair language understanding tasks that require various features of understanding, such as sentiment analysis, linguistic acceptability, semantic textual similarity, and question answering. We use seven datasets from the GLUE benchmark (RTE, MRPC, CoLA, SST-2, QNLI, QQP, MNLI) where the model has to choose between two or three (for MNLI) different choices for the respective task. Except for CoLA, we evaluate all used datasets with the accuracy metric. For CoLA, Matthews correlation is employed.
SAMSum (Gliwa et al., 2019).
SAMSum is a dataset for dialogue summarization research, comprising approximately 16,000 synthetic text conversations with accompanying summaries. Created by English-fluent linguists, these exchanges simulate real-world digital communications across various topics and styles. The conversations range from informal to formal, incorporating elements like slang and emoticons to reflect authentic messaging patterns. Each dialogue is paired with a concise, third-person summary, capturing its essential content. This structure makes SAMSum particularly useful for developing and evaluating automated summarization systems capable of processing conversational text.
Spider (Yu et al., 2018).
Spider is a large-scale, complex, and cross-domain semantic parsing and text-to-SQL dataset. It contains about 10,000 annotated SQL queries, distributed across 200+ databases, each with multiple tables. We follow Scholak et al. (2021) and use about 7,000 examples for training and about 1,000 examples for validation, where we ignore sequences longer than 1536 tokens. The dataset consists of English question and SQL query pairs, which cover a wide range of SQL operations including SELECT, WHERE, COUNT, GROUP BY, ORDER BY, JOIN, and more. Given an English question and an SQL database scheme, the task for the model is to translate the English question into an appropriate SQL statement. Evaluation is performed via accuracy where the output is considered as correct if the model’s predicted SQL query and the included GT SQL query give the same result when executed on the database. The dataset additionally categorizes each query into easy (25%), medium (40%), hard (20%), and extra hard (15%) based on the complexity of the required SQL statement. For evaluation, we report the execution accuracy of all categories.
DART (Nan et al., 2021).
The DART (DAta Record to Text) benchmark is a large-scale, structured dataset designed for RDF-to-text (Resource Description Framework-to-text) generation with 80,000+ instances. The DART benchmark is composed of a collection of structured data triples and corresponding text summaries which are organized into different categories. The task of the DART benchmark is to generate natural language summaries that correctly represent the given structured data inputs. DART is typically evaluated with METEOR and BLEU.
CIFAR-10 (Krizhevsky et al., 2009).
The CIFAR-10 (Canadian Institute For Advanced Research) dataset is a collection of images that are commonly used to train machine learning and computer vision algorithms. It is one of the most widely used datasets for image classification. The CIFAR-10 dataset contains 60,000 (50,000 for training, 10,000 for validation) 32$\times$32 color images in 10 different classes. The 10 different classes are: airplane, car, bird, cat, deer, dog, frog, horse, ship, and truck. There are 6,000 images of each class. For training, we center crop each image to 24$\times$24 pixels and flatten each image to a string, with a total of 24$\times$24$\times$3 words, where each word is a number between 0-255 representing the respective pixel value. Although CIFAR-10 is a dataset for computer vision, previous work (Dinh et al., 2022) showed that Transformers can be adapted to the vision domain from the language domain. In our work, we extend this investigation to SSMs, examining their ability to perform on vision data.
CelebA (Liu et al., 2015).
The CelebA (CelebFaces Attributes) dataset is an extensive collection of more than 200,000 celebrity images, each tagged with 40 attributes. This dataset is notable for its diversity, volume, and comprehensive annotations, encompassing 10,177 distinct identities, 202,599 facial images, and annotations of five landmark points with 40 binary attributes per image. The dataset, which includes images with varied poses and complex backgrounds, is an essential resource for tasks in computer vision such as face recognition, attribute analysis, and detection, as well as facial landmark localization, and it offers significant utility in face editing and synthesis.
Dataset Size (Train) Size (Val) Size (Test) Max. seq. len. #Epochs Mamba Size Jamba Size Metrics
GLUE RTE 1992 498 277 291 10 130M 319M Accuracy
MRPC 2934 734 408 105 10 130M 319M Accuracy
CoLA 6840 1711 1043 47 10 130M 319M Matthews corr.
SST-2 53879 13470 872 68 10 130M 319M Accuracy
QNLI 83794 20949 5463 602 10 130M 319M Accuracy
QQP 291076 72770 40430 316 3 130M 319M Accuracy
MNLI 314161 78541 19647 425 3 130M 319M Accuracy
Spider 5543 1375 1034 1412 10 1.4B, 2.8B 52B Accuracy
SAMSum 14732 818 819 1174 10 1.4B 52B ROUGE
DART 62659 2768 5097 491 10 130M 52B METEOR, BLEU
CIFAR-10 40000 10000 10000 1730 5 130M 319M Accuracy
CelebA 162770 19867 19962 12614 3 130M 319M Accuracy
Table 5: Datasets and models for our experiments. For each dataset, we report the number of training, validation, and test samples, maximum sequence length, training epochs, model size, and evaluation metric used.
The dataset characteristics, including our train, validation and test set sizes, sequence lengths, and number of epochs, are summarized in Table 5.
Appendix C Details of Sec. 4: Benchmarking PEFT Methods on SSM-based Models
In this section, we provide a comprehensive experimental setup, proofs and further discussion of theoretical results, and more detailed experimental outcomes.
C.1 Experiment Setup
[…]
C.2 Extended Results on Benchmarking Existing PEFT Methods
Mamba-I.
We present comprehensive fine-tuning results for the GLUE benchmark (Wang et al., 2019), DART dataset (Nan et al., 2021), SAMSum dataset (Gliwa et al., 2019), Spider dataset (Yu et al., 2018), and CIFAR-10 (Krizhevsky et al., 2009) in Table 6, Table 7, Table 8, Table 9, and Table 10 respectively. These experimental results encompass various LoRA implementations (on different weight matrices and modules) and provide more fine-grained results across all subtasks.
Mamba-II.
Table 11 and Table 12 present the benchmark results of LoRA and full fine-tuning across different layers of Mamba-II. We follow the same experimental setup used for Mamba-I and demonstrate that, on Mamba-II, our conclusion holds: LoRA is more effective on linear projection layers than on SSM modules.
Jamba.
Table 13 presents the benchmark results of LoRA and full fine-tuning across different layers of Jamba. Our findings demonstrate that, on Jamba, LoRA is more effective on linear projection layers than on SSM modules, which aligns with our conclusion on Mamba.
Layer Method # Params (%) RTE MRPC CoLA SST-2 QNLI QQP MNLI Avg.
Pretrained $0.0$ 0.469 3140.4693140.469\,3140.469 314 0.678 9220.6789220.678\,9220.678 922 0.000 $00\,0000$ 0.524 0830.5240830.524\,0830.524 083 0.505 2170.5052170.505\,2170.505 217 0.368 1180.3681180.368\,1180.368 118 0.322 5940.3225940.322\,5940.322 594 0.409 7500.4097500.409\,7500.409 750
All All Full $100.0$ 0.711 191 356 182 098 40.71119135618209840.711\,191\,356\,182\,098\,40.711 191 356 182 098 4 0.806 372 523 307 800 30.80637252330780030.806\,372\,523\,307\,800\,30.806 372 523 307 800 3 0.631 936 013 698 577 90.63193601369857790.631\,936\,013\,698\,577\,90.631 936 013 698 577 9 0.922 018 349 170 684 80.92201834917068480.922\,018\,349\,170\,684\,80.922 018 349 170 684 8 0.874 061 882 495 880 10.87406188249588010.874\,061\,882\,495\,880\,10.874 061 882 495 880 1 0.878 703 951 835 632 30.87870395183563230.878\,703\,951\,835\,632\,30.878 703 951 835 632 3 0.807 655 096 054 077 10.80765509605407710.807\,655\,096\,054\,077\,10.807 655 096 054 077 1 0.804 562 738 963 535 80.80456273896353580.804\,562\,738\,963\,535\,80.804 562 738 963 535 8
LoRA 1.922 527 283 438 617 31.92252728343861731.922\,527\,283\,438\,617\,31.922 527 283 438 617 3 0.699 275 374 412 536 60.69927537441253660.699\,275\,374\,412\,536\,60.699 275 374 412 536 6 0.808 823 525 905 609 10.80882352590560910.808\,823\,525\,905\,609\,10.808 823 525 905 609 1 0.614 111 959 934 234 60.61411195993423460.614\,111\,959\,934\,234\,60.614 111 959 934 234 6 0.918 577 969 074 249 30.91857796907424930.918\,577\,969\,074\,249\,30.918 577 969 074 249 3 0.884 312 629 699 7070.8843126296997070.884\,312\,629\,699\,7070.884 312 629 699 707 0.876 131 594 181 060 80.87613159418106080.876\,131\,594\,181\,060\,80.876 131 594 181 060 8 0.811 370 670 795 440 70.81137067079544070.811\,370\,670\,795\,440\,70.811 370 670 795 440 7 0.801 800 532 000 405 50.80180053200040550.801\,800\,532\,000\,405\,50.801 800 532 000 405 5
Prompt Prompt Tuning 16 tokens 0.009 514 691 278 001 4380.0095146912780014380.009\,514\,691\,278\,001\,4380.009 514 691 278 001 438 0.559 566 795 825 958 30.55956679582595830.559\,566\,795\,825\,958\,30.559 566 795 825 958 3 0.715 686 261 653 900 10.71568626165390010.715\,686\,261\,653\,900\,10.715 686 261 653 900 1 0.119 891 807 436 943 050.119891807436943050.119\,891\,807\,436\,943\,050.119 891 807 436 943 05 0.894 495 427 608 490.894495427608490.894\,495\,427\,608\,490.894 495 427 608 49 0.767 893 075 942 993 20.76789307594299320.767\,893\,075\,942\,993\,20.767 893 075 942 993 2 0.795 819 938 182 830 80.79581993818283080.795\,819\,938\,182\,830\,80.795 819 938 182 830 8 0.614 863 812 923 431 40.61486381292343140.614\,863\,812\,923\,431\,40.614 863 812 923 431 4 0.638 316 731 367 792 40.63831673136779240.638\,316\,731\,367\,792\,40.638 316 731 367 792 4
Prefix-Tuning 1 token (no MLP) 0.028 538 643 106 431 3040.0285386431064313040.028\,538\,643\,106\,431\,3040.028 538 643 106 431 304 0.675 090 253 353 118 90.67509025335311890.675\,090\,253\,353\,118\,90.675 090 253 353 118 9 0.757 352 948 188 781 70.75735294818878170.757\,352\,948\,188\,781\,70.757 352 948 188 781 7 0.434 039 652 347 564 70.43403965234756470.434\,039\,652\,347\,564\,70.434 039 652 347 564 7 0.915 137 588 977 813 70.91513758897781370.915\,137\,588\,977\,813\,70.915 137 588 977 813 7 0.834 157 049 655 914 30.83415704965591430.834\,157\,049\,655\,914\,30.834 157 049 655 914 3 0.831 066 012 382 507 30.83106601238250730.831\,066\,012\,382\,507\,30.831 066 012 382 507 3 0.356 390 297 412 872 30.35639029741287230.356\,390\,297\,412\,872\,30.356 390 297 412 872 3 0.686 176 257 474 081 80.68617625747408180.686\,176\,257\,474\,081\,80.686 176 257 474 081 8
[…]
Linear Projection Matrices All LoRA 1.017 230 519 270 278 41.01723051927027841.017\,230\,519\,270\,278\,41.017 230 519 270 278 4 0.700 361 013 412 475 60.70036101341247560.700\,361\,013\,412\,475\,60.700 361 013 412 475 6 0.823 529 422 283 172 60.82352942228317260.823\,529\,422\,283\,172\,60.823 529 422 283 172 6 0.577 130 556 106 567 40.57713055610656740.577\,130\,556\,106\,567\,40.577 130 556 106 567 4 0.933 486 223 220 825 20.93348622322082520.933\,486\,223\,220\,825\,20.933 486 223 220 825 2 0.887 241 423 130 035 40.88724142313003540.887\,241\,423\,130\,035\,40.887 241 423 130 035 4 0.887 261 927 127 838 10.88726192712783810.887\,261\,927\,127\,838\,10.887 261 927 127 838 1 0.825 316 846 370 6970.8253168463706970.825\,316\,846\,370\,6970.825 316 846 370 697 0.804 903 915 950 230 10.80490391595023010.804\,903\,915\,950\,230\,10.804 903 915 950 230 1
$\bm{W}_{\text{in},x}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 0.703 971 147 537 231 40.70397114753723140.703\,971\,147\,537\,231\,40.703 971 147 537 231 4 0.821 078 419 685 363 80.82107841968536380.821\,078\,419\,685\,363\,80.821 078 419 685 363 8 0.573 514 342 308 044 40.57351434230804440.573\,514\,342\,308\,044\,40.573 514 342 308 044 4 0.917 431 175 708 770 80.91743117570877080.917\,431\,175\,708\,770\,80.917 431 175 708 770 8 0.883 031 308 650 970 50.88303130865097050.883\,031\,308\,650\,970\,50.883 031 308 650 970 5 0.877 145 707 607 269 30.87714570760726930.877\,145\,707\,607\,269\,30.877 145 707 607 269 3 0.812 185 049 057 006 80.81218504905700680.812\,185\,049\,057\,006\,80.812 185 049 057 006 8 0.798 336 735 793 522 40.79833673579352240.798\,336\,735\,793\,522\,40.798 336 735 793 522 4
$\bm{W}_{\text{in},z}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 0.700 361 013 412 475 60.70036101341247560.700\,361\,013\,412\,475\,60.700 361 013 412 475 6 0.823 529 422 283 172 60.82352942228317260.823\,529\,422\,283\,172\,60.823 529 422 283 172 6 0.581 120 252 609 252 90.58112025260925290.581\,120\,252\,609\,252\,90.581 120 252 609 252 9 0.924 311 935 901 641 80.92431193590164180.924\,311\,935\,901\,641\,80.924 311 935 901 641 8 0.873 146 593 570 709 20.87314659357070920.873\,146\,593\,570\,709\,20.873 146 593 570 709 2 0.873 212 933 540 344 20.87321293354034420.873\,212\,933\,540\,344\,20.873 212 933 540 344 2 0.803 990 423 679 351 80.80399042367935180.803\,990\,423\,679\,351\,80.803 990 423 679 351 8 0.797 096 082 142 421 20.79709608214242120.797\,096\,082\,142\,421\,20.797 096 082 142 421 2
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.680 460 965 049 587 40.68046096504958740.680\,460\,965\,049\,587\,40.680 460 965 049 587 4 0.703 971 147 537 231 40.70397114753723140.703\,971\,147\,537\,231\,40.703 971 147 537 231 4 0.843 137 264 251 7090.8431372642517090.843\,137\,264\,251\,7090.843 137 264 251 709 0.624 137 103 557 586 70.62413710355758670.624\,137\,103\,557\,586\,70.624 137 103 557 586 7 0.925 458 729 267 120 40.92545872926712040.925\,458\,729\,267\,120\,40.925 458 729 267 120 4 0.886 326 193 809 509 30.88632619380950930.886\,326\,193\,809\,509\,30.886 326 193 809 509 3 0.883 353 948 593 139 60.88335394859313960.883\,353\,948\,593\,139\,60.883 353 948 593 139 6 0.817 427 575 588 226 30.81742757558822630.817\,427\,575\,588\,226\,30.817 427 575 588 226 3 0.811 973 137 514 931 80.81197313751493180.811\,973\,137\,514\,931\,80.811 973 137 514 931 8
$\bm{W}_{\text{out}}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 0.703 971 147 537 231 40.70397114753723140.703\,971\,147\,537\,231\,40.703 971 147 537 231 4 0.828 431 367 874 145 50.82843136787414550.828\,431\,367\,874\,145\,50.828 431 367 874 145 5 0.605 746 030 807 495 10.60574603080749510.605\,746\,030\,807\,495\,10.605 746 030 807 495 1 0.924 311 935 901 641 80.92431193590164180.924\,311\,935\,901\,641\,80.924 311 935 901 641 8 0.884 312 629 699 7070.8843126296997070.884\,312\,629\,699\,7070.884 312 629 699 707 0.876 675 724 983 215 30.87667572498321530.876\,675\,724\,983\,215\,30.876 675 724 983 215 3 0.814 831 793 308 258 10.81483179330825810.814\,831\,793\,308\,258\,10.814 831 793 308 258 1 0.805 468 661 444 527 80.80546866144452780.805\,468\,661\,444\,527\,80.805 468 661 444 527 8
S6 All Full 4.310 565 285 913 9444.3105652859139444.310\,565\,285\,913\,9444.310 565 285 913 944 0.696 750 879 287 719 70.69675087928771970.696\,750\,879\,287\,719\,70.696 750 879 287 719 7 0.789 215 683 937 072 80.78921568393707280.789\,215\,683\,937\,072\,80.789 215 683 937 072 8 0.590 752 780 437 469 50.59075278043746950.590\,752\,780\,437\,469\,50.590 752 780 437 469 5 0.915 137 588 977 813 70.91513758897781370.915\,137\,588\,977\,813\,70.915 137 588 977 813 7 0.880 651 652 812 957 80.88065165281295780.880\,651\,652\,812\,957\,80.880 651 652 812 957 8 0.875 290 632 247 924 80.87529063224792480.875\,290\,632\,247\,924\,80.875 290 632 247 924 8 0.804 906 606 674 194 30.80490660667419430.804\,906\,606\,674\,194\,30.804 906 606 674 194 3 0.793 243 689 196 450 40.79324368919645040.793\,243\,689\,196\,450\,40.793 243 689 196 450 4
LoRA 0.923 912 723 244 536 30.92391272324453630.923\,912\,723\,244\,536\,30.923 912 723 244 536 3 0.660 649 836 063 3850.6606498360633850.660\,649\,836\,063\,3850.660 649 836 063 385 0.786 764 681 339 263 90.78676468133926390.786\,764\,681\,339\,263\,90.786 764 681 339 263 9 0.578 182 518 482 208 30.57818251848220830.578\,182\,518\,482\,208\,30.578 182 518 482 208 3 0.908 256 888 389 587 40.90825688838958740.908\,256\,888\,389\,587\,40.908 256 888 389 587 4 0.877 722 859 382 629 40.87772285938262940.877\,722\,859\,382\,629\,40.877 722 859 382 629 4 0.868 859 767 913 818 40.86885976791381840.868\,859\,767\,913\,818\,40.868 859 767 913 818 4 0.798 137 128 353 118 90.79813712835311890.798\,137\,128\,353\,118\,90.798 137 128 353 118 9 0.782 653 382 846 287 30.78265338284628730.782\,653\,382\,846\,287\,30.782 653 382 846 287 3
${\bm{A}}$ Full 0.456 748 639 567 040 370.456748639567040370.456\,748\,639\,567\,040\,370.456 748 639 567 040 37 0.682 310 461 997 985 80.68231046199798580.682\,310\,461\,997\,985\,80.682 310 461 997 985 8 0.821 078 419 685 363 80.82107841968536380.821\,078\,419\,685\,363\,80.821 078 419 685 363 8 0.542 106 330 394 744 90.54210633039474490.542\,106\,330\,394\,744\,90.542 106 330 394 744 9 0.909 403 681 755 065 90.90940368175506590.909\,403\,681\,755\,065\,90.909 403 681 755 065 9 0.863 811 075 687 408 40.86381107568740840.863\,811\,075\,687\,408\,40.863 811 075 687 408 4 0.878 898 799 419 403 10.87889879941940310.878\,898\,799\,419\,403\,10.878 898 799 419 403 1 0.793 505 370 616 912 80.79350537061691280.793\,505\,370\,616\,912\,80.793 505 370 616 912 8 0.784 444 877 079 5550.7844448770795550.784\,444\,877\,079\,5550.784 444 877 079 555
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.283 743 197 835 2022.2837431978352022.283\,743\,197\,835\,2022.283 743 197 835 202 0.696 750 879 287 719 70.69675087928771970.696\,750\,879\,287\,719\,70.696 750 879 287 719 7 0.769 607 841 968 536 40.76960784196853640.769\,607\,841\,968\,536\,40.769 607 841 968 536 4 0.558 246 970 176 696 80.55824697017669680.558\,246\,970\,176\,696\,80.558 246 970 176 696 8 0.913 990 855 216 980.913990855216980.913\,990\,855\,216\,980.913 990 855 216 98 0.854 475 557 804 107 70.85447555780410770.854\,475\,557\,804\,107\,70.854 475 557 804 107 7 0.850 157 082 080 841 10.85015708208084110.850\,157\,082\,080\,841\,10.850 157 082 080 841 1 0.767 954 409 122 4670.7679544091224670.767\,954\,409\,122\,4670.767 954 409 122 467 0.773 026 227 951 049 80.77302622795104980.773\,026\,227\,951\,049\,80.773 026 227 951 049 8
LoRA 0.692 192 750 875 576 50.69219275087557650.692\,192\,750\,875\,576\,50.692 192 750 875 576 5 0.678 700 387 477 874 80.67870038747787480.678\,700\,387\,477\,874\,80.678 700 387 477 874 8 0.789 215 683 937 072 80.78921568393707280.789\,215\,683\,937\,072\,80.789 215 683 937 072 8 0.488 315 135 240 554 80.48831513524055480.488\,315\,135\,240\,554\,80.488 315 135 240 554 8 0.913 990 855 216 980.913990855216980.913\,990\,855\,216\,980.913 990 855 216 98 0.868 753 433 227 539 10.86875343322753910.868\,753\,433\,227\,539\,10.868 753 433 227 539 1 0.858 443 200 588 226 30.85844320058822630.858\,443\,200\,588\,226\,30.858 443 200 588 226 3 0.786 114 215 850 830 10.78611421585083010.786\,114\,215\,850\,830\,10.786 114 215 850 830 1 0.769 076 130 219 868 20.76907613021986820.769\,076\,130\,219\,868\,20.769 076 130 219 868 2
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.398 792 708 674 0611.3987927086740611.398\,792\,708\,674\,0611.398 792 708 674 061 0.660 649 836 063 3850.6606498360633850.660\,649\,836\,063\,3850.660 649 836 063 385 0.752 451 002 597 808 80.75245100259780880.752\,451\,002\,597\,808\,80.752 451 002 597 808 8 0.566 903 471 946 716 30.56690347194671630.566\,903\,471\,946\,716\,30.566 903 471 946 716 3 0.910 550 475 120 544 40.91055047512054440.910\,550\,475\,120\,544\,40.910 550 475 120 544 4 0.862 479 388 713 836 70.86247938871383670.862\,479\,388\,713\,836\,70.862 479 388 713 836 7 0.871 308 445 930 4810.8713084459304810.871\,308\,445\,930\,4810.871 308 445 930 481 0.784 801 721 572 8760.7848017215728760.784\,801\,721\,572\,8760.784 801 721 572 876 0.772 734 905 992 235 50.77273490599223550.772\,734\,905\,992\,235\,50.772 734 905 992 235 5
LoRA 0.234 957 666 087 183 550.234957666087183550.234\,957\,666\,087\,183\,550.234 957 666 087 183 55 0.671 480 119 228 3630.6714801192283630.671\,480\,119\,228\,3630.671 480 119 228 363 0.799 019 634 723 663 30.79901963472366330.799\,019\,634\,723\,663\,30.799 019 634 723 663 3 0.550 651 848 316 192 60.55065184831619260.550\,651\,848\,316\,192\,60.550 651 848 316 192 6 0.909 403 681 755 065 90.90940368175506590.909\,403\,681\,755\,065\,90.909 403 681 755 065 9 0.527 375 936 508 178 70.52737593650817870.527\,375\,936\,508\,178\,70.527 375 936 508 178 7 0.865 866 959 095 001 20.86586695909500120.865\,866\,959\,095\,001\,20.865 866 959 095 001 2 0.786 634 087 562 5610.7866340875625610.786\,634\,087\,562\,5610.786 634 087 562 561 0.730 061 752 455 575 10.73006175245557510.730\,061\,752\,455\,575\,10.730 061 752 455 575 1
Conv1d Full 0.142 733 949 864 700 120.142733949864700120.142\,733\,949\,864\,700\,120.142 733 949 864 700 12 0.682 310 461 997 985 80.68231046199798580.682\,310\,461\,997\,985\,80.682 310 461 997 985 8 0.784 313 738 346 099 90.78431373834609990.784\,313\,738\,346\,099\,90.784 313 738 346 099 9 0.578 959 524 631 500 20.57895952463150020.578\,959\,524\,631\,500\,20.578 959 524 631 500 2 0.910 550 475 120 544 40.91055047512054440.910\,550\,475\,120\,544\,40.910 550 475 120 544 4 0.859 600 961 208 343 50.85960096120834350.859\,600\,961\,208\,343\,50.859 600 961 208 343 5 0.859 980 225 563 049 30.85998022556304930.859\,980\,225\,563\,049\,30.859 980 225 563 049 3 0.779 508 292 675 018 30.77950829267501830.779\,508\,292\,675\,018\,30.779 508 292 675 018 3 0.779 317 668 506 077 30.77931766850607730.779\,317\,668\,506\,077\,30.779 317 668 506 077 3
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.043 414 909 750 512 950.043414909750512950.043\,414\,909\,750\,512\,950.043 414 909 750 512 95 0.653 429 627 418 518 10.65342962741851810.653\,429\,627\,418\,518\,10.653 429 627 418 518 1 0.791 666 686 534 881 60.79166668653488160.791\,666\,686\,534\,881\,60.791 666 686 534 881 6 0.403 467 684 984 207 150.403467684984207150.403\,467\,684\,984\,207\,150.403 467 684 984 207 15 0.910 550 475 120 544 40.91055047512054440.910\,550\,475\,120\,544\,40.910 550 475 120 544 4 0.839 069 962 501 525 90.83906996250152590.839\,069\,962\,501\,525\,90.839 069 962 501 525 9 0.859 824 895 858 764 60.85982489585876460.859\,824\,895\,858\,764\,60.859 824 895 858 764 6 0.670 331 358 909 606 90.67033135890960690.670\,331\,358\,909\,606\,90.670 331 358 909 606 9 0.732 620 098 761 149 90.73262009876114990.732\,620\,098\,761\,149\,90.732 620 098 761 149 9
Table 6: Full benchmark results on the GLUE (Wang et al., 2019) benchmark using Mamba-I 130M. We report accuracy ($\uparrow$) for RTE, MRPC, SST-2, QNLI, QQP, and MNLI tasks. CoLA performance is measured using Matthews Correlation Coefficient ($\uparrow$). In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for the SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) METEOR BLEU
All All Full $100.0$ 71.008 300 781 2571.0083007812571.008\,300\,781\,2571.008 300 781 25 51.802 164 316 177 3751.8021643161773751.802\,164\,316\,177\,3751.802 164 316 177 37
LoRA 1.922 527 283 438 617 31.92252728343861731.922\,527\,283\,438\,617\,31.922 527 283 438 617 3 70.968 002 080 917 3670.9680020809173670.968\,002\,080\,917\,3670.968 002 080 917 36 49.521 303 176 879 8849.5213031768798849.521\,303\,176\,879\,8849.521 303 176 879 88
DoRA 2.019 972.019972.019\,972.019 97 70.943 58870.94358870.943\,58870.943 588 51.364 43751.36443751.364\,43751.364 437
Prompt Prompt Tuning 64 tokens 0.038 047 904 689 998 750.038047904689998750.038\,047\,904\,689\,998\,750.038 047 904 689 998 75 66.186 875 104 904 1766.1868751049041766.186\,875\,104\,904\,1766.186 875 104 904 17 39.826 393 127 441 40639.82639312744140639.826\,393\,127\,441\,40639.826 393 127 441 406
Prefix-Tuning 64 tokens 22.688 043 993 029 53522.68804399302953522.688\,043\,993\,029\,53522.688 043 993 029 535 66.589 879 989 624 0266.5898799896240266.589\,879\,989\,624\,0266.589 879 989 624 02 42.462 074 756 622 31442.46207475662231442.462\,074\,756\,622\,31442.462 074 756 622 314
[…]
Linear Projection Matrices All LoRA 1.017 230 519 270 278 41.01723051927027841.017\,230\,519\,270\,278\,41.017 230 519 270 278 4 71.176 528 930 664 0671.1765289306640671.176\,528\,930\,664\,0671.176 528 930 664 06 49.160 829 186 439 51449.16082918643951449.160\,829\,186\,439\,51449.160 829 186 439 514
DoRA 1.087 1041.0871041.087\,1041.087 104 71.190 07771.19007771.190\,07771.190 077 50.797 19150.79719150.797\,19150.797 191
$\bm{W}_{\text{in},x}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 70.250 225 067 138 6770.2502250671386770.250\,225\,067\,138\,6770.250 225 067 138 67 48.860 117 793 083 1948.8601177930831948.860\,117\,793\,083\,1948.860 117 793 083 19
DoRA 0.369 7360.3697360.369\,7360.369 736 70.811 09670.81109670.811\,09670.811 096 49.932 89449.93289449.932\,89449.932 894
$\bm{W}_{\text{in},z}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 70.425 111 055 374 1570.4251110553741570.425\,111\,055\,374\,1570.425 111 055 374 15 49.059 996 008 872 98649.05999600887298649.059\,996\,008\,872\,98649.059 996 008 872 986
DoRA 0.369 7360.3697360.369\,7360.369 736 70.197 67470.19767470.197\,67470.197 674 48.341 75948.34175948.341\,75948.341 759
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.680 460 965 049 587 40.68046096504958740.680\,460\,965\,049\,587\,40.680 460 965 049 587 4 70.937 132 835 388 1870.9371328353881870.937\,132\,835\,388\,1870.937 132 835 388 18 49.452 367 424 964 90549.45236742496490549.452\,367\,424\,964\,90549.452 367 424 964 905
DoRA 0.736 7480.7367480.736\,7480.736 748 70.709 28870.70928870.709\,28870.709 288 51.550 57951.55057951.550\,57951.550 579
$\bm{W}_{\text{out}}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 70.727 145 671 844 4870.7271456718444870.727\,145\,671\,844\,4870.727 145 671 844 48 46.977 826 952 934 26546.97782695293426546.977\,826\,952\,934\,26546.977 826 952 934 265
DoRA 0.355 5660.3555660.355\,5660.355 566 70.706 88970.70688970.706\,88970.706 889 46.038 35746.03835746.038\,35746.038 357
S6 All Full 4.310 565 285 913 9444.3105652859139444.310\,565\,285\,913\,9444.310 565 285 913 944 70.348 870 754 241 9470.3488707542419470.348\,870\,754\,241\,9470.348 870 754 241 94 48.673 865 199 089 0548.6738651990890548.673\,865\,199\,089\,0548.673 865 199 089 05
LoRA 0.923 912 723 244 536 30.92391272324453630.923\,912\,723\,244\,536\,30.923 912 723 244 536 3 69.898 468 255 996 769.898468255996769.898\,468\,255\,996\,769.898 468 255 996 7 50.779 956 579 208 37450.77995657920837450.779\,956\,579\,208\,37450.779 956 579 208 374
DoRA 0.953 3850.9533850.953\,3850.953 385 70.150 12270.15012270.150\,12270.150 122 50.006 47150.00647150.006\,47150.006 471
${\bm{A}}$ Full 0.456 748 639 567 040 370.456748639567040370.456\,748\,639\,567\,040\,370.456 748 639 567 040 37 69.333 219 528 198 2469.3332195281982469.333\,219\,528\,198\,2469.333 219 528 198 24 48.095 324 635 505 67648.09532463550567648.095\,324\,635\,505\,67648.095 324 635 505 676
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.283 743 197 835 2022.2837431978352022.283\,743\,197\,835\,2022.283 743 197 835 202 70.056 617 259 979 2570.0566172599792570.056\,617\,259\,979\,2570.056 617 259 979 25 49.983 921 647 071 8449.9839216470718449.983\,921\,647\,071\,8449.983 921 647 071 84
LoRA 0.692 192 750 875 576 50.69219275087557650.692\,192\,750\,875\,576\,50.692 192 750 875 576 5 68.773 013 353 347 7868.7730133533477868.773\,013\,353\,347\,7868.773 013 353 347 78 47.991 544 008 255 00547.99154400825500547.991\,544\,008\,255\,00547.991 544 008 255 005
DoRA 0.693 6590.6936590.693\,6590.693 659 68.280 24968.28024968.280\,24968.280 249 47.327 52747.32752747.327\,52747.327 527
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.398 792 708 674 0611.3987927086740611.398\,792\,708\,674\,0611.398 792 708 674 061 69.576 507 806 777 9569.5765078067779569.576\,507\,806\,777\,9569.576 507 806 777 95 47.242 343 425 750 7347.2423434257507347.242\,343\,425\,750\,7347.242 343 425 750 73
LoRA 0.234 957 666 087 183 550.234957666087183550.234\,957\,666\,087\,183\,550.234 957 666 087 183 55 68.856 698 274 612 4368.8566982746124368.856\,698\,274\,612\,4368.856 698 274 612 43 47.047 722 339 630 1347.0477223396301347.047\,722\,339\,630\,1347.047 722 339 630 13
DoRA 0.263 3620.2633620.263\,3620.263 362 68.419 20868.41920868.419\,20868.419 208 46.260 97346.26097346.260\,97346.260 973
Conv1d Full 0.142 733 949 864 700 120.142733949864700120.142\,733\,949\,864\,700\,120.142 733 949 864 700 12 68.621 623 516 082 7668.6216235160827668.621\,623\,516\,082\,7668.621 623 516 082 76 47.934 231 162 071 2347.9342311620712347.934\,231\,162\,071\,2347.934 231 162 071 23
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.043 414 909 750 512 950.043414909750512950.043\,414\,909\,750\,512\,950.043 414 909 750 512 95 67.027 038 335 800 1767.0270383358001767.027\,038\,335\,800\,1767.027 038 335 800 17 44.228 667 020 797 7344.2286670207977344.228\,667\,020\,797\,7344.228 667 020 797 73
Table 7: Full benchmark results on the DART (Nan et al., 2021) benchmark using Mamba-I 130M. We report METEOR ()($\uparrow$)( ↑ ) and BLEU ()($\uparrow$)( ↑ ) scores. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) R1 R2 RL
All All Full $100.0$ $51.2$ $27.3$ $42.9$
LoRA 0.972 818 142 718 422 30.97281814271842230.972\,818\,142\,718\,422\,30.972 818 142 718 422 3 50.844 812 3950.8448123950.844\,812\,3950.844 812 39 26.647 731 6626.6477316626.647\,731\,6626.647 731 66 42.692 086 142.692086142.692\,086\,142.692 086 1
Prompt Prompt Tuning 64 tokens 0.009 551 198 1530.0095511981530.009\,551\,198\,1530.009 551 198 153 $50.1$ $25.6$ $41.6$
Prefix-Tuning 64 tokens 12.807 268 9112.8072689112.807\,268\,9112.807 268 91 $50.6$ $26.5$ $42.1$
[…]
Linear Projection Matrices All LoRA 0.513 166 979 724 261 80.51316697972426180.513\,166\,979\,724\,261\,80.513 166 979 724 261 8 50.820 308 9250.8203089250.820\,308\,9250.820 308 92 26.873 826 9826.8738269826.873\,826\,9826.873 826 98 42.773 035 1742.7730351742.773\,035\,1742.773 035 17
$\bm{W}_{\text{in},x}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 49.831 70349.83170349.831\,70349.831 703 25.436 69325.43669325.436\,69325.436 693 41.155 55341.15555341.155\,55341.155 553
$\bm{W}_{\text{in},z}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 50.021 547 0850.0215470850.021\,547\,0850.021 547 08 26.050 487 1626.0504871626.050\,487\,1626.050 487 16 41.673 398 0241.6733980241.673\,398\,0241.673 398 02
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.342 697 523 326 188 860.342697523326188860.342\,697\,523\,326\,188\,860.342 697 523 326 188 86 50.872 570 2850.8725702850.872\,570\,2850.872 570 28 26.967 412 2326.9674122326.967\,412\,2326.967 412 23 42.283 281 6842.2832816842.283\,281\,6842.283 281 68
$\bm{W}_{\text{out}}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 49.860 50149.86050149.860\,50149.860 501 25.443 37825.44337825.443\,37825.443 378 41.459 15541.45915541.459\,15541.459 155
S6 All Full 4.456 059 545 468 7924.4560595454687924.456\,059\,545\,468\,7924.456 059 545 468 792 51.134 622 151.134622151.134\,622\,151.134 622 1 26.891 985 5426.8919855426.891\,985\,5426.891 985 54 42.242 348 1942.2423481942.242\,348\,1942.242 348 19
LoRA 0.464 394 215 127 723 30.46439421512772330.464\,394\,215\,127\,723\,30.464 394 215 127 723 3 50.520 539 2850.5205392850.520\,539\,2850.520 539 28 26.363 494 9926.3634949926.363\,494\,9926.363 494 99 42.184 033 9942.1840339942.184\,033\,9942.184 033 99
${\bm{A}}$ Full 0.229 250 651 856 915 340.229250651856915340.229\,250\,651\,856\,915\,340.229 250 651 856 915 34 50.090 897 0850.0908970850.090\,897\,0850.090 897 08 25.944 900 5125.9449005125.944\,900\,5125.944 900 51 41.724 771 2641.7247712641.724\,771\,2641.724 771 26
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.292 506 518 569 153 82.29250651856915382.292\,506\,518\,569\,153\,82.292 506 518 569 153 8 50.457 620 6250.4576206250.457\,620\,6250.457 620 62 25.973 382 5925.9733825925.973\,382\,5925.973 382 59 41.786 766 0541.7867660541.786\,766\,0541.786 766 05
LoRA 0.347 144 240 365 408 70.34714424036540870.347\,144\,240\,365\,408\,70.347 144 240 365 408 7 50.416 523 2250.4165232250.416\,523\,2250.416 523 22 25.998 947 0225.9989470225.998\,947\,0225.998 947 02 41.848 915 8241.8489158241.848\,915\,8241.848 915 82
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.848 333 380 596 381.848333380596381.848\,333\,380\,596\,381.848 333 380 596 38 50.262 951 8550.2629518550.262\,951\,8550.262 951 85 25.663 086 7725.6630867725.663\,086\,7725.663 086 77 41.611 900 9341.6119009341.611\,900\,9341.611 900 93
LoRA 0.118 067 802 522 655 750.118067802522655750.118\,067\,802\,522\,655\,750.118 067 802 522 655 75 50.183 439 2550.1834392550.183\,439\,2550.183 439 25 25.424 981 1225.4249811225.424\,981\,1225.424 981 12 41.260 328 8941.2603288941.260\,328\,8941.260 328 89
Conv1d Full 0.071 640 828 705 286 060.071640828705286060.071\,640\,828\,705\,286\,060.071 640 828 705 286 06 50.085 21750.08521750.085\,21750.085 217 25.710 83025.71083025.710\,83025.710 830 41.920 88541.92088541.920\,88541.920 885
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.021 641 500 338 055 160.021641500338055160.021\,641\,500\,338\,055\,160.021 641 500 338 055 16 49.580 90849.58090849.580\,90849.580 908 24.796 23024.79623024.796\,23024.796 230 41.105 80741.10580741.105\,80741.105 807
Table 8: Full benchmark results on the SAMSum (Gliwa et al., 2019) benchmark using Mamba-I 1.4B. R1, R2, and RL represent ROUGE-1 ($\uparrow$), ROUGE-2 ($\uparrow$), and ROUGE-L ($\uparrow$), respectively. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) All Easy Medium Hard Extra
All All Full $100.0$ 66.150 867 938 995 3666.1508679389953666.150\,867\,938\,995\,3666.150 867 938 995 36 84.274 190 664 291 3884.2741906642913884.274\,190\,664\,291\,3884.274 190 664 291 38 69.506 728 649 139 469.506728649139469.506\,728\,649\,139\,469.506 728 649 139 4 53.448 277 711 868 28653.44827771186828653.448\,277\,711\,868\,28653.448 277 711 868 286 43.373 495 340 347 2943.3734953403472943.373\,495\,340\,347\,2943.373 495 340 347 29
LoRA 0.972 818 142 718 422 30.97281814271842230.972\,818\,142\,718\,422\,30.972 818 142 718 422 3 56.382 977 962 493 956.382977962493956.382\,977\,962\,493\,956.382 977 962 493 9 76.209 676 265 716 5576.2096762657165576.209\,676\,265\,716\,5576.209 676 265 716 55 56.950 670 480 728 1556.9506704807281556.950\,670\,480\,728\,1556.950 670 480 728 15 47.701 150 178 909 347.701150178909347.701\,150\,178\,909\,347.701 150 178 909 3 34.337 350 726 127 62534.33735072612762534.337\,350\,726\,127\,62534.337 350 726 127 625
DoRA 1.022 521.022521.022\,521.022 52 55.705 99655.70599655.705\,99655.705 996 77.016 12977.01612977.016\,12977.016 129 56.950 67356.95067356.950\,67356.950 673 47.126 43747.12643747.126\,43747.126 437 29.518 07229.51807229.518\,07229.518 072
Prompt Prompt Tuning 64 tokens 0.009 551 198 1530.0095511981530.009\,551\,198\,1530.009 551 198 153 43.617 022 0443.6170220443.617\,022\,0443.617 022 04 65.322 577 9565.3225779565.322\,577\,9565.322 577 95 42.376 682 1642.3766821642.376\,682\,1642.376 682 16 33.333 334 $3$.3333343333.333\,334\,$3$.333 334 33 25.301 206 1125.3012061125.301\,206\,1125.301 206 11
Prefix-Tuning 64 tokens 12.807 268 9112.8072689112.807\,268\,9112.807 268 91 39.651 837 9439.6518379439.651\,837\,9439.651 837 94 65.725 809 3465.7258093465.725\,809\,3465.725 809 34 38.565 021 7538.5650217538.565\,021\,7538.565 021 75 31.034 481 5331.0344815331.034\,481\,5331.034 481 53 15.060 241 5215.0602415215.060\,241\,5215.060 241 52
[…]
Linear Projection Matrices All LoRA 0.513 166 979 724 261 80.51316697972426180.513\,166\,979\,724\,261\,80.513 166 979 724 261 8 54.738 879 203 796 3954.7388792037963954.738\,879\,203\,796\,3954.738 879 203 796 39 $75.0$ 55.605 381 727 218 6355.6053817272186355.605\,381\,727\,218\,6355.605 381 727 218 63 45.977 011 322 975 1645.9770113229751645.977\,011\,322\,975\,1645.977 011 322 975 16 31.325 301 527 976 9931.3253015279769931.325\,301\,527\,976\,9931.325 301 527 976 99
DoRA 0.548 6080.5486080.548\,6080.548 608 57.156 67357.15667357.156\,67357.156 673 79.435 48479.43548479.435\,48479.435 484 58.744 39558.74439558.744\,39558.744 395 45.977 01145.97701145.977\,01145.977 011 31.325 30131.32530131.325\,30131.325 301
$\bm{W}_{\text{in},x}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 60.831 719 636 917 11460.83171963691711460.831\,719\,636\,917\,11460.831 719 636 917 114 76.612 901 687 622 0776.6129016876220776.612\,901\,687\,622\,0776.612 901 687 622 07 63.452 917 337 417 663.452917337417663.452\,917\,337\,417\,663.452 917 337 417 6 52.873 563 766 479 4952.8735637664794952.873\,563\,766\,479\,4952.873 563 766 479 49 38.554 215 431 213 3838.5542154312133838.554\,215\,431\,213\,3838.554 215 431 213 38
DoRA 0.185 920.185920.185\,920.185 92 58.413 92658.41392658.413\,92658.413 926 80.241 93580.24193580.241\,93580.241 935 60.089 68660.08968660.089\,68660.089 686 49.425 28749.42528749.425\,28749.425 287 30.722 89230.72289230.722\,89230.722 892
$\bm{W}_{\text{in},z}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 46.324 950 456 619 2646.3249504566192646.324\,950\,456\,619\,2646.324 950 456 619 26 68.548 387 289 047 2468.5483872890472468.548\,387\,289\,047\,2468.548 387 289 047 24 45.739 910 006 523 1345.7399100065231345.739\,910\,006\,523\,1345.739 910 006 523 13 36.781 609 058 380 1336.7816090583801336.781\,609\,058\,380\,1336.781 609 058 380 13 24.698 795 378 208 1624.6987953782081624.698\,795\,378\,208\,1624.698 795 378 208 16
DoRA 0.185 920.185920.185\,920.185 92 59.767 89259.76789259.767\,89259.767 892 83.870 96883.87096883.870\,96883.870 968 60.089 68660.08968660.089\,68660.089 686 50.574 71350.57471350.574\,71350.574 713 32.530 1232.5301232.530\,1232.530 12
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.342 697 523 326 188 860.342697523326188860.342\,697\,523\,326\,188\,860.342 697 523 326 188 86 57.543 522 119 522 09557.54352211952209557.543\,522\,119\,522\,09557.543 522 119 522 095 77.419 352 531 433 177.419352531433177.419\,352\,531\,433\,177.419 352 531 433 1 58.744 394 779 205 3258.7443947792053258.744\,394\,779\,205\,3258.744 394 779 205 32 45.402 297 377 586 36545.40229737758636545.402\,297\,377\,586\,36545.402 297 377 586 365 37.349 396 944 046 0237.3493969440460237.349\,396\,944\,046\,0237.349 396 944 046 02
DoRA 0.371 150.371150.371\,150.371 15 60.735 0160.7350160.735\,0160.735 01 78.629 03278.62903278.629\,03278.629 032 62.107 62362.10762362.107\,62362.107 623 52.873 56352.87356352.873\,56352.873 563 38.554 21738.55421738.554\,21738.554 217
$\bm{W}_{\text{out}}$ LoRA 0.171 642 869 594 627 820.171642869594627820.171\,642\,869\,594\,627\,820.171 642 869 594 627 82 61.798 840 761 184 6961.7988407611846961.798\,840\,761\,184\,6961.798 840 761 184 69 81.854 838 132 858 2881.8548381328582881.854\,838\,132\,858\,2881.854 838 132 858 28 65.246 635 675 430 365.246635675430365.246\,635\,675\,430\,365.246 635 675 430 3 45.402 297 377 586 36545.40229737758636545.402\,297\,377\,586\,36545.402 297 377 586 365 39.759 036 898 612 97639.75903689861297639.759\,036\,898\,612\,97639.759 036 898 612 976
DoRA 0.178 7820.1787820.178\,7820.178 782 61.315 2861.3152861.315\,2861.315 28 79.435 48479.43548479.435\,48479.435 484 63.901 34563.90134563.901\,34563.901 345 $50$ 39.156 62739.15662739.156\,62739.156 627
S6 All Full 4.456 059 545 468 7924.4560595454687924.456\,059\,545\,468\,7924.456 059 545 468 792 56.673 115 491 867 06556.67311549186706556.673\,115\,491\,867\,06556.673 115 491 867 065 76.612 901 687 622 0776.6129016876220776.612\,901\,687\,622\,0776.612 901 687 622 07 57.847 535 610 198 97557.84753561019897557.847\,535\,610\,198\,97557.847 535 610 198 975 45.977 011 322 975 1645.9770113229751645.977\,011\,322\,975\,1645.977 011 322 975 16 34.939 759 969 711 30434.93975996971130434.939\,759\,969\,711\,30434.939 759 969 711 304
LoRA 0.464 394 215 127 723 30.46439421512772330.464\,394\,215\,127\,723\,30.464 394 215 127 723 3 56.286 269 426 345 82556.28626942634582556.286\,269\,426\,345\,82556.286 269 426 345 825 $75.0$ 56.502 240 896 224 97656.50224089622497656.502\,240\,896\,224\,97656.502 240 896 224 976 50.574 713 945 388 79450.57471394538879450.574\,713\,945\,388\,79450.574 713 945 388 794 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
DoRA 0.479 1420.4791420.479\,1420.479 142 58.897 48558.89748558.897\,48558.897 485 77.419 35577.41935577.419\,35577.419 355 62.107 62362.10762362.107\,62362.107 623 47.126 43747.12643747.126\,43747.126 437 34.939 75934.93975934.939\,75934.939 759
${\bm{A}}$ Full 0.229 250 651 856 915 340.229250651856915340.229\,250\,651\,856\,915\,340.229 250 651 856 915 34 51.063 829 660 415 6551.0638296604156551.063\,829\,660\,415\,6551.063 829 660 415 65 71.370 965 242 385 8671.3709652423858671.370\,965\,242\,385\,8671.370 965 242 385 86 52.466 368 675 231 93452.46636867523193452.466\,368\,675\,231\,93452.466 368 675 231 934 42.528 736 591 339 1142.5287365913391142.528\,736\,591\,339\,1142.528 736 591 339 11 25.903 615 355 491 63825.90361535549163825.903\,615\,355\,491\,63825.903 615 355 491 638
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.292 506 518 569 153 82.29250651856915382.292\,506\,518\,569\,153\,82.292 506 518 569 153 8 47.195 357 084 274 2947.1953570842742947.195\,357\,084\,274\,2947.195 357 084 274 29 72.177 422 046 661 3872.1774220466613872.177\,422\,046\,661\,3872.177 422 046 661 38 46.860 986 948 013 30646.86098694801330646.860\,986\,948\,013\,30646.860 986 948 013 306 35.632 184 147 834 7835.6321841478347835.632\,184\,147\,834\,7835.632 184 147 834 78 22.891 566 157 341 00322.89156615734100322.891\,566\,157\,341\,00322.891 566 157 341 003
LoRA 0.347 144 240 365 408 70.34714424036540870.347\,144\,240\,365\,408\,70.347 144 240 365 408 7 55.029 010 772 705 0855.0290107727050855.029\,010\,772\,705\,0855.029 010 772 705 08 73.790 323 734 283 4573.7903237342834573.790\,323\,734\,283\,4573.790 323 734 283 45 56.726 455 688 476 5656.7264556884765656.726\,455\,688\,476\,5656.726 455 688 476 56 44.252 872 467 041 01644.25287246704101644.252\,872\,467\,041\,01644.252 872 467 041 016 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
DoRA $0.3477$ 55.319 14955.31914955.319\,14955.319 149 78.225 80678.22580678.225\,80678.225 806 57.847 53457.84753457.847\,53457.847 534 41.379 3141.3793141.379\,3141.379 31 28.915 66328.91566328.915\,66328.915 663
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.848 333 380 596 381.848333380596381.848\,333\,380\,596\,381.848 333 380 596 38 56.769 824 028 015 1456.7698240280151456.769\,824\,028\,015\,1456.769 824 028 015 14 77.016 127 109 527 5977.0161271095275977.016\,127\,109\,527\,5977.016 127 109 527 59 59.417 039 155 960 0859.4170391559600859.417\,039\,155\,960\,0859.417 039 155 960 08 43.678 161 501 884 4643.6781615018844643.678\,161\,501\,884\,4643.678 161 501 884 46 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
LoRA 0.118 067 802 522 655 750.118067802522655750.118\,067\,802\,522\,655\,750.118 067 802 522 655 75 58.027 076 721 191 40658.02707672119140658.027\,076\,721\,191\,40658.027 076 721 191 406 78.629 034 757 614 1478.6290347576141478.629\,034\,757\,614\,1478.629 034 757 614 14 59.417 039 155 960 0859.4170391559600859.417\,039\,155\,960\,0859.417 039 155 960 08 48.850 575 089 454 6548.8505750894546548.850\,575\,089\,454\,6548.850 575 089 454 65 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
DoRA 0.132 360.132360.132\,360.132 36 55.319 14955.31914955.319\,14955.319 149 76.209 67776.20967776.209\,67776.209 677 59.192 82559.19282559.192\,82559.192 825 42.528 73642.52873642.528\,73642.528 736 27.108 43427.10843427.108\,43427.108 434
Conv1d Full 0.071 640 828 705 286 060.071640828705286060.071\,640\,828\,705\,286\,060.071 640 828 705 286 06 53.191 488 981 246 9553.1914889812469553.191\,488\,981\,246\,9553.191 488 981 246 95 74.596 774 578 094 4874.5967745780944874.596\,774\,578\,094\,4874.596 774 578 094 48 52.914 798 259 735 1152.9147982597351152.914\,798\,259\,735\,1152.914 798 259 735 11 43.678 161 501 884 4643.6781615018844643.678\,161\,501\,884\,4643.678 161 501 884 46 31.927 710 771 560 6731.9277107715606731.927\,710\,771\,560\,6731.927 710 771 560 67
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.021 641 500 338 055 160.021641500338055160.021\,641\,500\,338\,055\,160.021 641 500 338 055 16 49.613 153 934 478 7649.6131539344787649.613\,153\,934\,478\,7649.613 153 934 478 76 70.564 514 398 574 8370.5645143985748370.564\,514\,398\,574\,8370.564 514 398 574 83 50.448 429 584 503 17450.44842958450317450.448\,429\,584\,503\,17450.448 429 584 503 174 40.229 883 790 016 17440.22988379001617440.229\,883\,790\,016\,17440.229 883 790 016 174 25.903 615 355 491 63825.90361535549163825.903\,615\,355\,491\,63825.903 615 355 491 638
(a) Full benchmark results on Spider using Mamba-I 1.4B.
Layer Method # Params (%) All Easy Medium Hard Extra
All All Full $100.0$ 71.760 153 770 446 7871.7601537704467871.760\,153\,770\,446\,7871.760 153 770 446 78 $87.5$ 73.542 600 870 132 4573.5426008701324573.542\,600\,870\,132\,4573.542 600 870 132 45 63.793 104 887 008 6763.7931048870086763.793\,104\,887\,008\,6763.793 104 887 008 67 51.807 230 710 983 27651.80723071098327651.807\,230\,710\,983\,27651.807 230 710 983 276
LoRA 0.804 873 122 224 144 40.80487312222414440.804\,873\,122\,224\,144\,40.804 873 122 224 144 4 70.889 747 142 791 7570.8897471427917570.889\,747\,142\,791\,7570.889 747 142 791 75 90.725 809 335 708 6290.7258093357086290.725\,809\,335\,708\,6290.725 809 335 708 62 73.991 030 454 635 6273.9910304546356273.991\,030\,454\,635\,6273.991 030 454 635 62 58.620 691 299 438 4858.6206912994384858.620\,691\,299\,438\,4858.620 691 299 438 48 45.783 132 314 682 0145.7831323146820145.783\,132\,314\,682\,0145.783 132 314 682 01
Prompt Prompt Tuning 64 tokens 0.005 917 985 961 427 6770.0059179859614276770.005\,917\,985\,961\,427\,6770.005 917 985 961 427 677 50.676 983 594 894 4150.6769835948944150.676\,983\,594\,894\,4150.676 983 594 894 41 75.403 225 421 905 5275.4032254219055275.403\,225\,421\,905\,5275.403 225 421 905 52 53.811 657 428 741 45553.81165742874145553.811\,657\,428\,741\,45553.811 657 428 741 455 37.356 323 003 768 9237.3563230037689237.356\,323\,003\,768\,9237.356 323 003 768 92 19.277 107 715 606 6919.2771077156066919.277\,107\,715\,606\,6919.277 107 715 606 69
Prefix-Tuning 1 token 10.820 948 510 794 74110.82094851079474110.820\,948\,510\,794\,74110.820 948 510 794 741 45.067 697 763 442 9945.0676977634429945.067\,697\,763\,442\,9945.067 697 763 442 99 $75.0$ 45.067 265 629 768 3745.0672656297683745.067\,265\,629\,768\,3745.067 265 629 768 37 32.183 909 416 198 7332.1839094161987332.183\,909\,416\,198\,7332.183 909 416 198 73 13.855 421 543 121 33813.85542154312133813.855\,421\,543\,121\,33813.855 421 543 121 338
[…]
Linear Projection Matrices All LoRA 0.424 312 127 242 072 460.424312127242072460.424\,312\,127\,242\,072\,460.424 312 127 242 072 46 58.220 505 714 416 50458.22050571441650458.220\,505\,714\,416\,50458.220 505 714 416 504 74.596 774 578 094 4874.5967745780944874.596\,774\,578\,094\,4874.596 774 578 094 48 58.295 965 194 702 1558.2959651947021558.295\,965\,194\,702\,1558.295 965 194 702 15 51.724 135 875 701 90451.72413587570190451.724\,135\,875\,701\,90451.724 135 875 701 904 40.361 446 142 196 65540.36144614219665540.361\,446\,142\,196\,65540.361 446 142 196 655
$\bm{W}_{\text{in},x}$ LoRA 0.141 838 601 338 417 10.14183860133841710.141\,838\,601\,338\,417\,10.141 838 601 338 417 1 66.731 142 997 741 766.731142997741766.731\,142\,997\,741\,766.731 142 997 741 7 87.903 225 421 905 5287.9032254219055287.903\,225\,421\,905\,5287.903 225 421 905 52 67.713 004 350 662 2367.7130043506622367.713\,004\,350\,662\,2367.713 004 350 662 23 56.896 549 463 272 09556.89654946327209556.896\,549\,463\,272\,09556.896 549 463 272 095 42.771 083 116 531 3742.7710831165313742.771\,083\,116\,531\,3742.771 083 116 531 37
$\bm{W}_{\text{in},z}$ LoRA 0.141 838 601 338 417 10.14183860133841710.141\,838\,601\,338\,417\,10.141 838 601 338 417 1 65.377 175 807 952 8865.3771758079528865.377\,175\,807\,952\,8865.377 175 807 952 88 86.693 549 156 188 9686.6935491561889686.693\,549\,156\,188\,9686.693 549 156 188 96 68.834 078 311 920 1768.8340783119201768.834\,078\,311\,920\,1768.834 078 311 920 17 54.597 699 642 181 454.597699642181454.597\,699\,642\,181\,454.597 699 642 181 4 35.542 169 213 294 9835.5421692132949835.542\,169\,213\,294\,9835.542 169 213 294 98
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.283 275 408 799 057 940.283275408799057940.283\,275\,408\,799\,057\,940.283 275 408 799 057 94 65.183 752 775 192 2665.1837527751922665.183\,752\,775\,192\,2665.183 752 775 192 26 89.112 901 687 622 0789.1129016876220789.112\,901\,687\,622\,0789.112 901 687 622 07 67.264 574 766 159 0667.2645747661590667.264\,574\,766\,159\,0667.264 574 766 159 06 51.724 135 875 701 90451.72413587570190451.724\,135\,875\,701\,90451.724 135 875 701 904 37.951 806 187 629 737.951806187629737.951\,806\,187\,629\,737.951 806 187 629 7
$\bm{W}_{\text{out}}$ LoRA 0.141 838 601 338 417 10.14183860133841710.141\,838\,601\,338\,417\,10.141 838 601 338 417 1 67.021 274 566 650 3967.0212745666503967.021\,274\,566\,650\,3967.021 274 566 650 39 87.096 774 578 094 4887.0967745780944887.096\,774\,578\,094\,4887.096 774 578 094 48 69.058 293 104 171 7569.0582931041717569.058\,293\,104\,171\,7569.058 293 104 171 75 52.873 563 766 479 4952.8735637664794952.873\,563\,766\,479\,4952.873 563 766 479 49 46.385 541 558 265 68646.38554155826568646.385\,541\,558\,265\,68646.385 541 558 265 686
S6 All Full 4.438 752 155 800 24.43875215580024.438\,752\,155\,800\,24.438 752 155 800 2 65.667 313 337 326 0565.6673133373260565.667\,313\,337\,326\,0565.667 313 337 326 05 81.854 838 132 858 2881.8548381328582881.854\,838\,132\,858\,2881.854 838 132 858 28 68.834 078 311 920 1768.8340783119201768.834\,078\,311\,920\,1768.834 078 311 920 17 58.045 977 354 049 6858.0459773540496858.045\,977\,354\,049\,6858.045 977 354 049 68 40.963 855 385 780 33440.96385538578033440.963\,855\,385\,780\,33440.963 855 385 780 334
LoRA 0.383 804 949 263 572 40.38380494926357240.383\,804\,949\,263\,572\,40.383 804 949 263 572 4 63.926 500 082 015 9963.9265000820159963.926\,500\,082\,015\,9963.926 500 082 015 99 86.290 323 734 283 4586.2903237342834586.290\,323\,734\,283\,4586.290 323 734 283 45 68.161 433 935 165 468.161433935165468.161\,433\,935\,165\,468.161 433 935 165 4 49.425 286 054 611 20649.42528605461120649.425\,286\,054\,611\,20649.425 286 054 611 206 34.337 350 726 127 62534.33735072612762534.337\,350\,726\,127\,62534.337 350 726 127 625
${\bm{A}}$ Full 0.189 386 758 647 475 230.189386758647475230.189\,386\,758\,647\,475\,230.189 386 758 647 475 23 56.576 400 995 254 5256.5764009952545256.576\,400\,995\,254\,5256.576 400 995 254 52 77.016 127 109 527 5977.0161271095275977.016\,127\,109\,527\,5977.016 127 109 527 59 58.071 750 402 450 5658.0717504024505658.071\,750\,402\,450\,5658.071 750 402 450 56 45.977 011 322 975 1645.9770113229751645.977\,011\,322\,975\,1645.977 011 322 975 16 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.272 641 103 769 702 72.27264110376970272.272\,641\,103\,769\,702\,72.272 641 103 769 702 7 58.800 774 812 698 36458.80077481269836458.800\,774\,812\,698\,36458.800 774 812 698 364 79.032 260 179 519 6579.0322601795196579.032\,260\,179\,519\,6579.032 260 179 519 65 60.986 548 662 185 6760.9865486621856760.986\,548\,662\,185\,6760.986 548 662 185 67 50.574 713 945 388 79450.57471394538879450.574\,713\,945\,388\,79450.574 713 945 388 794 31.325 301 527 976 9931.3253015279769931.325\,301\,527\,976\,9931.325 301 527 976 99
LoRA 0.286 806 195 766 252 20.28680619576625220.286\,806\,195\,766\,252\,20.286 806 195 766 252 2 60.251 450 538 635 25460.25145053863525460.251\,450\,538\,635\,25460.251 450 538 635 254 82.661 288 976 669 3182.6612889766693182.661\,288\,976\,669\,3182.661 288 976 669 31 63.004 481 792 449 9563.0044817924499563.004\,481\,792\,449\,9563.004 481 792 449 95 46.551 725 268 363 9546.5517252683639546.551\,725\,268\,363\,9546.551 725 268 363 95 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.905 704 258 890 219 31.90570425889021931.905\,704\,258\,890\,219\,31.905 704 258 890 219 3 62.185 686 826 705 9362.1856868267059362.185\,686\,826\,705\,9362.185 686 826 705 93 82.258 063 554 763 882.258063554763882.258\,063\,554\,763\,882.258 063 554 763 8 65.695 065 259 933 4765.6950652599334765.695\,065\,259\,933\,4765.695 065 259 933 47 51.724 135 875 701 90451.72413587570190451.724\,135\,875\,701\,90451.724 135 875 701 904 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
LoRA 0.097 557 280 258 231 170.097557280258231170.097\,557\,280\,258\,231\,170.097 557 280 258 231 17 62.185 686 826 705 9362.1856868267059362.185\,686\,826\,705\,9362.185 686 826 705 93 80.241 936 445 236 280.241936445236280.241\,936\,445\,236\,280.241 936 445 236 2 66.591 930 389 404 366.591930389404366.591\,930\,389\,404\,366.591 930 389 404 3 49.425 286 054 611 20649.42528605461120649.425\,286\,054\,611\,20649.425 286 054 611 206 36.746 987 700 462 3436.7469877004623436.746\,987\,700\,462\,3436.746 987 700 462 34
Conv1d Full 0.059 183 362 077 336 010.059183362077336010.059\,183\,362\,077\,336\,010.059 183 362 077 336 01 62.475 824 356 079 162.475824356079162.475\,824\,356\,079\,162.475 824 356 079 1 81.854 838 132 858 2881.8548381328582881.854\,838\,132\,858\,2881.854 838 132 858 28 66.143 494 844 436 6566.1434948444366566.143\,494\,844\,436\,6566.143 494 844 436 65 51.149 427 890 777 5951.1494278907775951.149\,427\,890\,777\,5951.149 427 890 777 59 35.542 169 213 294 9835.5421692132949835.542\,169\,213\,294\,9835.542 169 213 294 98
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.017 847 482 626 446 640.017847482626446640.017\,847\,482\,626\,446\,640.017 847 482 626 446 64 50.967 115 163 803 150.967115163803150.967\,115\,163\,803\,150.967 115 163 803 1 70.967 739 820 480 3570.9677398204803570.967\,739\,820\,480\,3570.967 739 820 480 35 51.121 073 961 257 93551.12107396125793551.121\,073\,961\,257\,93551.121 073 961 257 935 42.528 736 591 339 1142.5287365913391142.528\,736\,591\,339\,1142.528 736 591 339 11 29.518 070 816 993 71329.51807081699371329.518\,070\,816\,993\,71329.518 070 816 993 713
(b) Full benchmark results on Spider using Mamba-I 2.8B.
Table 9: Full benchmark results on Spider (Yu et al., 2018) dataset using Mamba-I. We report the accuracy ($\uparrow$) for Spider and its subsets. We consider two models in our experiments: Mamba-I 1.4B and Mamba-I 2.8B. In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) Accuracy
Pretrained $0.0$ 0.081 5000.0815000.081\,5000.081 500
All All Full $100.0$ 59.960 001 707 077 02659.96000170707702659.960\,001\,707\,077\,02659.960 001 707 077 026
LoRA 1.922 527 283 438 617 31.92252728343861731.922\,527\,283\,438\,617\,31.922 527 283 438 617 3 60.350 000 858 306 88560.35000085830688560.350\,000\,858\,306\,88560.350 000 858 306 885
[…]
Linear Projection Matrices All LoRA 1.017 230 519 270 278 41.01723051927027841.017\,230\,519\,270\,278\,41.017 230 519 270 278 4 62.790 000 438 690 18662.79000043869018662.790\,000\,438\,690\,18662.790 000 438 690 186
$\bm{W}_{\text{in},x}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 53.490 000 963 211 0653.4900009632110653.490\,000\,963\,211\,0653.490 000 963 211 06
$\bm{W}_{\text{in},z}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 58.149 999 380 111 69458.14999938011169458.149\,999\,380\,111\,69458.149 999 380 111 694
$\bm{W}_{\text{in},x},\bm{W}_{\text{in},z}$ LoRA 0.680 460 965 049 587 40.68046096504958740.680\,460\,965\,049\,587\,40.680 460 965 049 587 4 61.040 002 107 620 2461.0400021076202461.040\,002\,107\,620\,2461.040 002 107 620 24
$\bm{W}_{\text{out}}$ LoRA 0.341 392 002 181 115 60.34139200218111560.341\,392\,002\,181\,115\,60.341 392 002 181 115 6 52.039 998 769 760 1352.0399987697601352.039\,998\,769\,760\,1352.039 998 769 760 13
S6 All Full 4.310 565 285 913 9444.3105652859139444.310\,565\,285\,913\,9444.310 565 285 913 944 55.510 002 374 649 0555.5100023746490555.510\,002\,374\,649\,0555.510 002 374 649 05
LoRA 0.923 912 723 244 536 30.92391272324453630.923\,912\,723\,244\,536\,30.923 912 723 244 536 3 43.959 999 084 472 65643.95999908447265643.959\,999\,084\,472\,65643.959 999 084 472 656
${\bm{A}}$ Full 0.456 748 639 567 040 370.456748639567040370.456\,748\,639\,567\,040\,370.456 748 639 567 040 37 61.210 000 514 984 1361.2100005149841361.210\,000\,514\,984\,1361.210 000 514 984 13
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ Full 2.283 743 197 835 2022.2837431978352022.283\,743\,197\,835\,2022.283 743 197 835 202 49.509 999 155 998 2349.5099991559982349.509\,999\,155\,998\,2349.509 999 155 998 23
LoRA 0.692 192 750 875 576 50.69219275087557650.692\,192\,750\,875\,576\,50.692 192 750 875 576 5 52.270 001 173 019 4152.2700011730194152.270\,001\,173\,019\,4152.270 001 173 019 41
${\bm{W}}_{\bm{\Delta},\uparrow}$ Full 1.398 792 708 674 0611.3987927086740611.398\,792\,708\,674\,0611.398 792 708 674 061 34.540 000 557 899 47534.54000055789947534.540\,000\,557\,899\,47534.540 000 557 899 475
LoRA 0.234 957 666 087 183 550.234957666087183550.234\,957\,666\,087\,183\,550.234 957 666 087 183 55 56.489 998 102 188 1156.4899981021881156.489\,998\,102\,188\,1156.489 998 102 188 11
Conv1d Full 0.142 733 949 864 700 120.142733949864700120.142\,733\,949\,864\,700\,120.142 733 949 864 700 12 55.650 001 764 297 48555.65000176429748555.650\,001\,764\,297\,48555.650 001 764 297 485
Others ${\bm{D}},\texttt{LayerNorm}$ Full 0.043 414 909 750 512 950.043414909750512950.043\,414\,909\,750\,512\,950.043 414 909 750 512 95 58.090 001 344 680 78658.09000134468078658.090\,001\,344\,680\,78658.090 001 344 680 786
Table 10: Full benchmark results on the CIFAR-10 (Krizhevsky et al., 2009) dataset using Mamba-I 130M. We report accuracy ($\uparrow$). In each Mamba block, $\bm{W}_{\text{in},x}$ and $\bm{W}_{\text{in},z}$ are input projections that preprocess the input for SSM modules and the gating branch, respectively. $\bm{W}_{\text{out}}$ denotes the output projection after the gating mechanism. $\bm{W}_{{\bm{B}}}$ and $\bm{W}_{{\bm{C}}}$ are weight matrices for computing input-dependent ${\bm{B}}_{n}$ and ${\bm{C}}_{n}$. ${\bm{W}}_{\bm{\Delta},\downarrow}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$ represent down and up projections of low-rank weight matrices in the linear layer computing input-dependent step size $\bm{\Delta}_{n}$. $\bm{\beta}_{\bm{\Delta}}$ represents the bias in this linear layer. ${\bm{D}}$ denotes the weight of residual connections.
Layer Method # Params (%) METEOR BLEU
All All Full $100.0$ 66.565 710 306 167 666.565710306167666.565\,710\,306\,167\,666.565 710 306 167 6 34.870 043 396 949 7734.8700433969497734.870\,043\,396\,949\,7734.870 043 396 949 77
LoRA 1.393 771 181 403 076 71.39377118140307671.393\,771\,181\,403\,076\,71.393 771 181 403 076 7 66.923 677 921 295 1766.9236779212951766.923\,677\,921\,295\,1766.923 677 921 295 17 45.405 793 190 002 4445.4057931900024445.405\,793\,190\,002\,4445.405 793 190 002 44
Linear Projection Matrices $\bm{W}_{\text{in}},\bm{W}_{\text{out}}$ LoRA 1.018 368 047 523 842 11.01836804752384211.018\,368\,047\,523\,842\,11.018 368 047 523 842 1 67.109 721 899 032 5967.1097218990325967.109\,721\,899\,032\,5967.109 721 899 032 59 44.712 835 550 308 2344.7128355503082344.712\,835\,550\,308\,2344.712 835 550 308 23
$\bm{W}_{\text{in}}$ LoRA 0.681 224 489 192 509 40.68122448919250940.681\,224\,489\,192\,509\,40.681 224 489 192 509 4 67.064 607 143 402 167.064607143402167.064\,607\,143\,402\,167.064 607 143 402 1 43.031 200 766 563 41643.03120076656341643.031\,200\,766\,563\,41643.031 200 766 563 416
$\bm{W}_{\text{out}}$ LoRA 0.341 776 376 784 720 950.341776376784720950.341\,776\,376\,784\,720\,950.341 776 376 784 720 95 66.796 505 451 202 3966.7965054512023966.796\,505\,451\,202\,3966.796 505 451 202 39 42.278 078 198 432 9242.2780781984329242.278\,078\,198\,432\,9242.278 078 198 432 92
S6 All Full 4.169 116 475 966 0684.1691164759660684.169\,116\,475\,966\,0684.169 116 475 966 068 65.723 252 296 447 7565.7232522964477565.723\,252\,296\,447\,7565.723 252 296 447 75 39.696 556 329 727 1739.6965563297271739.696\,556\,329\,727\,1739.696 556 329 727 17
LoRA 0.383 152 556 698 126 50.38315255669812650.383\,152\,556\,698\,126\,50.383 152 556 698 126 5 64.183 932 542 800 964.183932542800964.183\,932\,542\,800\,964.183 932 542 800 9 40.119 516 849 517 8240.1195168495178240.119\,516\,849\,517\,8240.119 516 849 517 82
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},\bm{W}_{\bm{\Delta}}$ Full 4.001 065 760 075 972 54.00106576007597254.001\,065\,760\,075\,972\,54.001 065 760 075 972 5 65.965 962 409 973 1465.9659624099731465.965\,962\,409\,973\,1465.965 962 409 973 14 36.188 501 119 613 6536.1885011196136536.188\,501\,119\,613\,6536.188 501 119 613 65
LoRA 0.383 152 556 698 126 50.38315255669812650.383\,152\,556\,698\,126\,50.383 152 556 698 126 5 64.831 387 996 673 5864.8313879966735864.831\,387\,996\,673\,5864.831 387 996 673 58 39.472 505 450 248 7239.4725054502487239.472\,505\,450\,248\,7239.472 505 450 248 72
Table 11: Full benchmark results of LoRA on DART (Nan et al., 2021) dataset using Mamba-II 130M.
Layer Method # Params (%) All Easy Medium Hard Extra
All All Full $100.0$ 64.796 906 709 671 0264.7969067096710264.796\,906\,709\,671\,0264.796 906 709 671 02 85.887 098 312 377 9385.8870983123779385.887\,098\,312\,377\,9385.887 098 312 377 93 65.695 065 259 933 4765.6950652599334765.695\,065\,259\,933\,4765.695 065 259 933 47 54.022 985 696 792 654.022985696792654.022\,985\,696\,792\,654.022 985 696 792 6 42.168 673 872 947 6942.1686738729476942.168\,673\,872\,947\,6942.168 673 872 947 69
LoRA 0.706 416 999 407 535 90.70641699940753590.706\,416\,999\,407\,535\,90.706 416 999 407 535 9 64.506 769 180 297 8564.5067691802978564.506\,769\,180\,297\,8564.506 769 180 297 85 81.048 387 289 047 2481.0483872890472481.048\,387\,289\,047\,2481.048 387 289 047 24 66.367 715 597 152 7166.3677155971527166.367\,715\,597\,152\,7166.367 715 597 152 71 56.896 549 463 272 09556.89654946327209556.896\,549\,463\,272\,09556.896 549 463 272 095 42.771 083 116 531 3742.7710831165313742.771\,083\,116\,531\,3742.771 083 116 531 37
Linear Projection Matrices $\bm{W}_{\text{in}},\bm{W}_{\text{out}}$ LoRA 0.523 963 841 037 011 80.52396384103701180.523\,963\,841\,037\,011\,80.523 963 841 037 011 8 50.386 846 065 521 2450.3868460655212450.386\,846\,065\,521\,2450.386 846 065 521 24 68.548 387 289 047 2468.5483872890472468.548\,387\,289\,047\,2468.548 387 289 047 24 52.017 939 090 728 7652.0179390907287652.017\,939\,090\,728\,7652.017 939 090 728 76 44.827 586 412 429 8144.8275864124298144.827\,586\,412\,429\,8144.827 586 412 429 81 24.698 795 378 208 1624.6987953782081624.698\,795\,378\,208\,1624.698 795 378 208 16
$\bm{W}_{\text{in}}$ LoRA 0.349 920 379 444 911 50.34992037944491150.349\,920\,379\,444\,911\,50.349 920 379 444 911 5 57.543 522 119 522 09557.54352211952209557.543\,522\,119\,522\,09557.543 522 119 522 095 76.209 676 265 716 5576.2096762657165576.209\,676\,265\,716\,5576.209 676 265 716 55 59.417 039 155 960 0859.4170391559600859.417\,039\,155\,960\,0859.417 039 155 960 08 48.850 575 089 454 6548.8505750894546548.850\,575\,089\,454\,6548.850 575 089 454 65 33.734 938 502 311 7133.7349385023117133.734\,938\,502\,311\,7133.734 938 502 311 71
$\bm{W}_{\text{out}}$ LoRA 0.175 266 836 912 838 980.175266836912838980.175\,266\,836\,912\,838\,980.175 266 836 912 838 98 57.930 368 185 043 33557.93036818504333557.930\,368\,185\,043\,33557.930 368 185 043 335 81.048 387 289 047 2481.0483872890472481.048\,387\,289\,047\,2481.048 387 289 047 24 56.726 455 688 476 5656.7264556884765656.726\,455\,688\,476\,5656.726 455 688 476 56 51.724 135 875 701 90451.72413587570190451.724\,135\,875\,701\,90451.724 135 875 701 904 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
S6 All Full 2.419 408 304 585 2852.4194083045852852.419\,408\,304\,585\,2852.419 408 304 585 285 55.125 725 269 317 6355.1257252693176355.125\,725\,269\,317\,6355.125 725 269 317 63 76.209 676 265 716 5576.2096762657165576.209\,676\,265\,716\,5576.209 676 265 716 55 56.053 811 311 721 856.053811311721856.053\,811\,311\,721\,856.053 811 311 721 8 42.528 736 591 339 1142.5287365913391142.528\,736\,591\,339\,1142.528 736 591 339 11 34.337 350 726 127 62534.33735072612762534.337\,350\,726\,127\,62534.337 350 726 127 625
LoRA 0.184 378 487 046 726 70.18437848704672670.184\,378\,487\,046\,726\,70.184 378 487 046 726 7 54.061 895 608 901 9854.0618956089019854.061\,895\,608\,901\,9854.061 895 608 901 98 74.193 549 156 188 9674.1935491561889674.193\,549\,156\,188\,9674.193 549 156 188 96 58.071 750 402 450 5658.0717504024505658.071\,750\,402\,450\,5658.071 750 402 450 56 45.977 011 322 975 1645.9770113229751645.977\,011\,322\,975\,1645.977 011 322 975 16 21.686 747 670 173 64521.68674767017364521.686\,747\,670\,173\,64521.686 747 670 173 645
${\bm{A}_{\text{log}}}$ Full 0.000 228 612 709 494 971 640.000228612709494971640.000\,228\,612\,709\,494\,971\,640.000 228 612 709 494 971 64 21.470 019 221 305 84721.47001922130584721.470\,019\,221\,305\,84721.470 019 221 305 847 45.967 742 800 712 58545.96774280071258545.967\,742\,800\,712\,58545.967 742 800 712 585 18.834 081 292 152 40518.83408129215240518.834\,081\,292\,152\,40518.834 081 292 152 405 11.494 252 830 743 7911.4942528307437911.494\,252\,830\,743\,7911.494 252 830 743 79 2.409 638 464 450 8362.4096384644508362.409\,638\,464\,450\,8362.409 638 464 450 836
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},\bm{W}_{\bm{\Delta}}$ Full 2.340 994 145 228 512.340994145228512.340\,994\,145\,228\,512.340 994 145 228 51 50.290 137 529 373 1750.2901375293731750.290\,137\,529\,373\,1750.290 137 529 373 17 72.983 872 890 472 4172.9838728904724172.983\,872\,890\,472\,4172.983 872 890 472 41 52.242 153 882 980 3552.2421538829803552.242\,153\,882\,980\,3552.242 153 882 980 35 39.655 172 824 859 6239.6551728248596239.655\,172\,824\,859\,6239.655 172 824 859 62 22.289 156 913 757 32422.28915691375732422.289\,156\,913\,757\,32422.289 156 913 757 324
LoRA 0.184 378 487 046 726 70.18437848704672670.184\,378\,487\,046\,726\,70.184 378 487 046 726 7 55.512 571 334 838 8755.5125713348388755.512\,571\,334\,838\,8755.512 571 334 838 87 77.419 352 531 433 177.419352531433177.419\,352\,531\,433\,177.419 352 531 433 1 55.156 952 142 715 45455.15695214271545455.156\,952\,142\,715\,45455.156 952 142 715 454 46.551 725 268 363 9546.5517252683639546.551\,725\,268\,363\,9546.551 725 268 363 95 33.132 529 258 728 0333.1325292587280333.132\,529\,258\,728\,0333.132 529 258 728 03
Table 12: Full benchmark results on the Spider (Yu et al., 2018) dataset using Mamba-II 1.3B.
Layer Method # Params (%) METEOR BLEU
All All Full $100.0$ $70.79$ $45.04$
Attention All LoRA 0.016 7060.0167060.016\,7060.016 706 $63.47$ $19.67$
MLP All LoRA 1.369 0841.3690841.369\,0841.369 084 $70.87$ $46.2$
Linear Projection Matrices + S6 All LoRA 0.308 3140.3083140.308\,3140.308 314 $70.16$ $39.99$
Linear Projection Matrices $\bm{W}_{\text{in}}$ LoRA 0.107 8460.1078460.107\,8460.107 846 $68.85$ $37.76$
$\bm{W}_{\text{out}}$ LoRA 0.053 9520.0539520.053\,9520.053 952 $67.67$ $31.85$
S6 All Full 0.535 3140.5353140.535\,3140.535 314 $69.23$ $35.49$
$\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\downarrow}$ LoRA 0.147 1070.1471070.147\,1070.147 107 $66.55$ $24.16$
Table 13: Full benchmark results on DART (Nan et al., 2021) dataset using Jamba-Tiny-319M.
C.3 Limitations of Applying Input-injection Methods on SSMs
Task RTE MRPC CoLA SST-2 QNLI QQP MNLI Avg. Score
Prompt Tuning 56.0 71.6 12.0 89.4 76.8 79.6 61.5 63.8
Prefix-Tuning 69.5 75.7 43.4 91.5 83.4 83.1 35.6 68.6
Initial State Tuning 66.8 75.1 52.4 92.4 86.4 86.1 78.5 76.8
LoRA (Linear Projection Matrices) 70.4 82.8 60.6 92.4 88.4 87.7 81.5 80.5
Table 14: Comparison of prompt-tuning, prefix-tuning, initial state tuning, and LoRA on seven tasks from the GLUE benchmark. We report Matthews correlation ($\uparrow$) for CoLA, overall (matched and mismatched) accuracy ()($\uparrow$)( ↑ ) for MNLI, and accuracy for other tasks. Initial state tuning and LoRA are constrained to use less than 0.5% trainable parameters. Bold numbers indicate the best performance across all three methods, while underlined numbers show the highest score among input-injection methods (prefix-tuning and initial state tuning). Initial state tuning outperforms prefix-tuning and prompt-tuning on five out of seven tasks, while LoRA consistently outperforms all input-injection methods.
We start by introducing the necessary notations. Denote the space of S4 mechanisms with $D$ channels as ${\mathcal{F}}_{\text{S4},D}$. Let ${\bm{H}}_{0}=({\bm{h}}_{0}^{(1)},{\bm{h}}_{0}^{(2)},\dots,{\bm{h}}_{0}^{(D)})\in{\mathbb{R}}^{H\times D}$ represent the initial hidden state, and ${\bm{X}}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$ denote the input sequence. The output of the S4 mechanism is represented as $f({\bm{X}};{\bm{H}}_{0})$. Furthermore, for $d$-th channel, let state transition matrix $\overline{{\bm{A}}}^{(d)}=\operatorname{diag}{(a_{1}^{(d)},\cdots,a_{H}^{(d)})}$ and input transition vector $\overline{{\bm{B}}}^{(d)}=(b_{1},\cdots,b_{H})^{\top}$, where $d=1,\ldots,D$. For any vector ${\bm{v}}\in{\mathbb{R}}^{n}$, we use ${\bm{v}}_{i:j}\in{\mathbb{R}}^{j-i}$ to denote the subvector of ${\bm{v}}$ containing elements from $i\in{\mathbb{N}}^{+}$ to $j\in{\mathbb{N}}^{+}$, where $i<j$. Similarly, for any matrix ${\bm{M}}\in{\mathbb{R}}^{m\times n}$, we use ${\bm{M}}_{i_{1}:j_{1},i_{2}:j_{2}}$ to denote the submatrix containing rows $i_{1}\in{\mathbb{N}}^{+}$ to $j_{1}\in{\mathbb{N}}^{+}$ and columns $i_{2}\in{\mathbb{N}}^{+}$ to $j_{2}\in{\mathbb{N}}^{+}$, where $i_{1}<j_{1}$, $i_{2}<j_{2}$.
Proposition 1 (Expressivity of Prefix-Tuning on SSMs).
Let $f\in{\mathcal{F}}_{\text{S4},D}$ be an S4 mechanism. Consider prefix-tuning that prepends a sequence ${\bm{P}}=({\bm{p}}_{1},\ldots,{\bm{p}}_{M})\in{\mathbb{R}}^{D\times M}$ to the input sequence ${\bm{X}}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$. For any prefix ${\bm{P}}\in{\mathbb{R}}^{D\times M}$, there exists an initial hidden state ${\bm{H}}_{0}^{\star}\in{\mathbb{R}}^{H\times D}$ such that the output of S4 after prefix-tuning and that after initial state tuning are identical, i.e., $f({\bm{X}};{\bm{H}}_{0}^{\star})\equiv f([{\bm{P}},{\bm{X}}];{\bm{H}}_{0})_{1:D,M+1:M+N}$ for all ${\bm{X}}\in{\mathbb{R}}^{D\times N}$.
Furthermore, assume that $\prod_{0\leq i<j\leq H}(a_{j}^{(d)}-a_{i}^{(d)})\neq 0$ and $\prod_{k=1}^{H}b_{k}^{(d)}\neq 0$ for all channels $d=1,\ldots,D$. Then the converse (i.e., for any ${\bm{H}}_{0}\in{\mathbb{R}}^{H\times D}$, there exists a ${\bm{P}}^{\star}\in{\mathbb{R}}^{D\times M}$ such that $f([{\bm{P}}^{\star},{\bm{X}}];{\bm{H}}_{0})_{1:D,M+1:M+N}\equiv f({\bm{X}};{\bm{H}}_{0}^{\star})$ for all ${\bm{X}}\in{\mathbb{R}}^{D\times N}$) holds if and only if $M\geq H$.
Proof of Proposition 1.
Given that operations in S4 are independent across all channels, we can, without loss of generality, consider the case where the number of channels $D=1$. Consequently, we can simplify our notation: the initial hidden states ${\bm{H}}_{0}\in{\mathbb{R}}^{H\times D}$ become ${\bm{h}}_{0}\in{\mathbb{R}}^{H}$, the input sequence ${\bm{X}}\in{\mathbb{R}}^{D\times N}$ becomes ${\bm{x}}\in{\mathbb{R}}^{N}$, and the prefix ${\bm{P}}\in{\mathbb{R}}^{D\times M}$ becomes ${\bm{p}}\in{\mathbb{R}}^{M}$. We omit the superscript $(d)$ denoting the channel index. To differentiate between the hidden states and output of prefix-tuned S4 (i.e., $f([{\bm{P}},{\bm{X}}];{\bm{H}}_{0})_{1:D,M+1:M+N}$) and initial state tuned S4 (i.e., $f({\bm{X}};{\bm{H}}_{0}^{\star})$), we introduce superscripts “PT” and “IST” respectively. The “PT” superscript denotes hidden states and output of S4 after prefix-tuning, while “IST” indicates those after initial state tuning.
We divide the proposition into two statements:
  1. 1.
    For any prefix ${\bm{p}}\in{\mathbb{R}}^{M}$, there exists an initial hidden state ${\bm{h}}_{0}^{\star}\in{\mathbb{R}}^{H}$ such that the output of S4 after prefix-tuning and that after initial state tuning are identical, i.e., $f({\bm{x}};{\bm{h}}_{0}^{\star})\equiv f([{\bm{p}},{\bm{x}}];{\bm{h}}_{0})_{M+1:N+M}$ for all ${\bm{x}}\in{\mathbb{R}}^{N}$.
  2. 2.
    Furthermore, assume that $\prod_{0\leq i<j\leq H}(a_{j}-a_{i})\neq 0$ and $\prod_{k=1}^{H}b_{k}\neq 0$. Then the converse (i.e., for any ${\bm{h}}_{0}\in{\mathbb{R}}^{H}$, there exists a ${\bm{p}}^{\star}\in{\mathbb{R}}^{M}$ such that $f([{\bm{p}}^{\star},{\bm{x}}];{\bm{h}}_{0})_{M+1:N+M}\equiv f({\bm{x}};{\bm{h}}_{0}^{\star})$ for all ${\bm{x}}\in{\mathbb{R}}^{N}$) holds if and only if $M\geq H$.
We will first prove the first statement and then proceed to prove the second statement.
Statement 1. The recurrent computation formulation of S4 in (2) implies that for each position $i$, the output $y_{i}$ depends solely on the previous hidden state $h_{i-1}$ and the current input $x_{i}$. Thus, to demonstrate that $f({\bm{x}};{\bm{h}}_{0}^{\star})\equiv f([{\bm{p}},{\bm{x}}];{\bm{h}}_{0})_{M+1:N+M}$ for all ${\bm{x}}\in{\mathbb{R}}^{N}$, it suffices to show that the hidden state for predicting output $y^{\text{IST}}_{1}$ equals that for predicting output $y^{\text{PT}}_{M+1}$, where $y^{\text{IST}}_{1}$ and $y^{\text{PT}}_{M+1}$ are outputs corresponding to the input $x_{1}$ for initial state tuning and prefix-tuning, respectively. In other words, it is sufficient to show that the initial state of initial-state-tuned model ${\bm{h}}^{\text{IST}}_{0}={\bm{h}}_{0}^{\star}$ is equal to the $(M+1)$-th hidden state of prefix-tuned model ${\bm{h}}^{\text{PT}}_{M+1}=\sum_{m=1}^{M}\overline{{\bm{A}}}^{M-m}\overline{{\bm{B}}}p_{m}$. When this equality holds, the subsequent hidden states and outputs for both versions of S4 will be identical, as the input sequence from that point onward is the same. Therefore, we prove the first statement by letting
${\bm{h}}_{0}^{\star}=\sum_{m=1}^{M}\overline{{\bm{A}}}^{M-m}\overline{{\bm{B}}}p_{m}.$ (15)
Statement 2. We aim to investigate the conditions under which there exists a ${\bm{h}}_{0}^{\star}\in{\mathbb{R}}^{H}$ such that for any ${\bm{p}}\in{\mathbb{R}}^{M}$, $f([{\bm{p}}^{\star},{\bm{x}}];{\bm{h}}_{0})_{M+1:N+M}\neq f({\bm{x}};{\bm{h}}_{0}^{\star})$. This is equivalent to demonstrating the existence of ${\bm{h}}_{0}^{\star}\in{\mathbb{R}}^{H}$ such that
${\bm{h}}_{0}^{\star}\neq\sum_{m=1}^{M}\overline{{\bm{A}}}^{M-m}\overline{{\bm{B}}}p_{m},\quad\text{for all }{\bm{p}}\in{\mathbb{R}}^{M}.$ (16)
This condition can be further reformulated as
${\mathbb{R}}^{H}\setminus\text{span}(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})\neq\emptyset,$ (17)
which is equivalent to
$\text{span}(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})\subsetneq{\mathbb{R}}^{H}.$ (18)
To determine when this condition holds, we analyze three distinct cases: (i) $M<H$, (ii) $M=H$, and (iii) $M>H$.
(Case 1: When $M<H$). In this scenario, it is obvious that (18) holds. The existence of such a ${\bm{h}}_{0}^{\star}$ is guaranteed because the dimension of the span is at most $M$, which is strictly less than $H$. This choice of ${\bm{h}}_{0}^{\star}$ ensures that it cannot be represented as a linear combination of the vectors in the span, thereby establishing the inequality.
(Case 2: When $M=H$). In this scenario, $\text{span}(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})={\mathbb{R}}^{H}$ if and only if $(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})$ are linearly independent. Note that
$\det(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})=\det(\overline{{\bm{A}}}^{M},\overline{{\bm{A}}}^{M-1},\ldots,{\bm{1}})\prod_{k=1}^{H}b_{k},$ (19)
where
[…]
Combining (19) and (21) yields
$\det(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})=(-1)^{\frac{H(H-1)}{2}}\prod_{0\leq i<j\leq H}^{H}(a_{j}-a_{i})\prod_{k=1}^{H}b_{k}.$ (22)
Therefore, if and only if $\prod_{1\leq i<j\leq H}(a_{j}-a_{i})\neq 0$ and $\prod_{k=1}^{H}b_{k}\neq 0$, we have
$\det(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})\neq 0,$ (23)
which is both necessary and sufficient for the linear independence of $(\overline{{\bm{A}}}^{M}\overline{{\bm{B}}},\overline{{\bm{A}}}^{M-1}\overline{{\bm{B}}},\ldots,\overline{{\bm{B}}})$, and consequently, for the condition in (18) to be satisfied.
(Case 3: When $M>H$). The analysis presented in case 2 extends naturally to this scenario.
The combination of the three cases above completes the proof of statement 2. ∎
C.4 Optimal Application of LoRA in SSM-based Models
Several studies (Hu et al., 2023; He et al., 2021) present findings on Transformers, indicating that applying LoRA to linear projection matrices yields performance comparable to or marginally superior to that of attention layers. In contrast, our experimental results on SSMs reveal that applying LoRA to linear projection matrices is more effective than applying it to S6. To elucidate this phenomenon, we examine the influence of updating linear projection matrices on the model’s output.
Notations.
To make the analysis tractable, we consider a simplified SSM-based architecture composed of the following components:
  • Two input projection matrices: ${\bm{W}}_{\text{in,1}},{\bm{W}}_{\text{in,2}}\in{\mathbb{R}}^{D\times D}$;
  • The S6 module parameterized by diagonal state transition matrices $\{{\bm{A}}^{(d)}\}_{d=1}^{D}$ with ${\bm{A}}^{(d)}\in{\mathbb{R}}^{H\times H}$, the weight matrices $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}\in{\mathbb{R}}^{H\times D}$ for computing input-dependent input transition vectors ${\bm{B}}_{n}\in{\mathbb{R}}^{H}$ and output mapping vectors ${\bm{C}}_{n}\in{\mathbb{R}}^{H}$, the down and up projection matrices ${\bm{W}}_{\bm{\Delta},\downarrow}\in{\mathbb{R}}^{D\times R},{\bm{W}}_{\bm{\Delta},\uparrow}\in{\mathbb{R}}^{R\times D}$ (where $R$ is the rank) for low-rank weight matrices for computing the input-dependent step size $\bm{\Delta}_{n}=(\Delta^{(1)}_{n},\ldots,\Delta^{(D)}_{n})\in{\mathbb{R}}_{D}$, for $n=1,\ldots,N$.
Define $\bm{W}_{\text{S6}}=[\bm{W}_{{\bm{B}}}^{\top},\bm{W}_{{\bm{C}}}^{\top},{\bm{W}}_{\bm{\Delta},\uparrow}^{\top}]^{\top}\in{\mathbb{R}}^{(2H+R)\times D}$. In the Mamba implementation, $\bm{W}_{\text{S6}}$ is implemented as the weight matrix of a single linear layer, referred to as x_proj in the codebase. Let the input sequence be ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{N})\in{\mathbb{R}}^{D\times N}$. At each time step $n$, the S6 module uses two differently projected versions of the input: (i) the input projected via ${\bm{W}}_{\text{in,1}}$ is used to compute the input-dependent parameters $\overline{{\bm{A}}}_{n}$, $\overline{{\bm{B}}}_{n}$, and ${\bm{C}}_{n}$, and (ii) the input projected via ${\bm{W}}_{\text{in,2}}$ serves as the actual input to the S6 module. We note that this formulation generalizes the standard case, which uses a single input projection matrix before the S6 module. In particular, it reduces to the standard case when ${\bm{W}}_{\text{in,1}}={\bm{W}}_{\text{in,2}}$. Then, the output at time step $N$ is given by:
$y^{(d)}_{N}=\underbrace{\overbrace{{\bm{C}}({\bm{W}}_{\text{in,1}}{\bm{x}}_{N})^{\top}}^{\text{Input-dependent }{\bm{C}}_{N}}\sum_{n=1}^{N}\left(\prod_{m=1}^{n}\overbrace{\overline{{\bm{A}}}({\bm{W}}_{\text{in,1}}{\bm{x}}_{m})}^{\text{Input-dependent }\overline{{\bm{A}}}_{m}}\right)\overbrace{\overline{{\bm{B}}}_{n}({\bm{W}}_{\text{in,1}}{\bm{x}}_{n})}^{\text{Input-dependent }\overline{{\bm{B}}}_{n}}}_{\text{Parameters depending on input after projection }{\bm{W}}_{\text{in,1}}}\underbrace{({\bm{W}}_{\text{in,2}}{\bm{x}}_{n})^{(d)}}_{\text{Input after projection}{\bm{W}}_{\text{in,2}}}.$ (24)
To be more specific, the definitions for the relevant terms are:
$\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\text{in, 1}}\bm{x}_{n}+\bm{\beta}_{\bm{\Delta}}),$ (25)
$\overline{{\bm{A}}}_{n}^{(d)}=\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)}),\quad\overline{{\bm{B}}}_{n}^{(d)}=\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}{\bm{W}}_{\text{in, 1}}\bm{x}_{n},\quad{\bm{C}}_{n}=\bm{W}_{{\bm{C}}}{\bm{W}}_{\text{in, 1}}\bm{x}_{n}.$ (26)
When $\bm{\beta}_{\bm{\Delta}}={\bm{0}}$, the output at time step $N$ can be further written as
$y^{(d)}_{N}=(\bm{W}_{{\bm{C}}}{\bm{W}}_{\text{in, 1}}\bm{x}_{n})^{\top}\sum_{n=1}^{N}\left(\prod_{m=1}^{n}\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)})\right)\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}{\bm{W}}_{\text{in, 1}}\bm{x}_{n}({\bm{W}}_{\text{in,2}}^{\star}{\bm{x}}_{n})^{(d)},$ (27)
where $\text{where }\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\text{in, 1}}\bm{x}_{n}).$ (28)
Theoretical Analysis.
Assume none of the parameters are zero and $D>2H+R$, where $R$ is the rank of ${\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}$. Lemma 1 in the main body shows that applying LoRA solely to ${\bm{W}}_{\text{in,1}}$ is equivalent to applying it to $\bm{W}_{\text{S6}}$. For completeness, we provide the proof below and restate the lemma for the reader’s convenience.
Lemma 0 (Expressivity of Fine-Tuning Projection Matrices).
Consider two models with the architecture described above. Let:
  • A target model $f^{\star}$ parameterized by ($(\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, $\bm{W}_{{\bm{B}}}^{\star}$, $\bm{W}_{{\bm{C}}}^{\star}$, ${\bm{W}}_{\bm{\Delta},\uparrow}^{\star}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, ${\bm{W}}_{\text{in,1}}^{\star}$, ${\bm{W}}_{\text{in,2}}^{\star})$;
  • A frozen model $f_{0}$ parameterized by ($(\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, ${\bm{W}}_{\text{in,1}}$, ${\bm{W}}_{\text{in,2}}^{\star})$.
The two models share $\{{\bm{A}}^{\star(d)}\}_{d=1}^{D}$, ${\bm{W}}_{\bm{\Delta},\downarrow}^{\star}$, and ${\bm{W}}_{\text{in,2}}^{\star}$, while differing in $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$, and ${\bm{W}}_{\text{in,1}}$. Then, there exists a projection matrix $\widehat{{\bm{W}}}_{\text{in,1}}$ such that the frozen model matches the output of the target model for any input sequence, i.e.,
$f(\cdot;\{{\bm{A}}^{\star(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}},{\bm{W}}_{\bm{\Delta},\uparrow},{\bm{W}}_{\bm{\Delta},\downarrow}^{\star},\widehat{{\bm{W}}_{\text{in,1}}},{\bm{W}}_{\text{in,2}}^{\star})=f^{\star}(\cdot;\{{\bm{A}}^{\star(d)}\}_{d=1}^{D},\bm{W}_{{\bm{B}}}^{\star},\bm{W}_{{\bm{C}}}^{\star},{\bm{W}}_{\bm{\Delta},\uparrow}^{\star},{\bm{W}}_{\bm{\Delta},\downarrow}^{\star},{\bm{W}}_{\text{in,1}}^{\star},{\bm{W}}_{\text{in,2}}^{\star}).$ (29)
Proof of Lemma 1.
To prove (29), we substitute (27) into (29), simplify the expression, and show that the equality holds under the following conditions:
$\bm{W}_{{\bm{C}}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}=\bm{W}_{{\bm{C}}}{\bm{W}}_{\text{in,1}}$ (30)
${\bm{W}}_{\bm{\Delta},\uparrow}^{\star}{\bm{W}}_{\text{in,1}}^{\star}={\bm{W}}_{\bm{\Delta},\uparrow}{\bm{W}}_{\text{in,1}}$ (31)
$\bm{W}_{{\bm{B}}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}=\bm{W}_{{\bm{B}}}{\bm{W}}_{\text{in,1}}.$ (32)
Since $\bm{W}_{\text{S6}}=\begin{bmatrix}\bm{W}_{{\bm{B}}}\\ \bm{W}_{{\bm{C}}}\\ {\bm{W}}_{\bm{\Delta},\uparrow}\end{bmatrix}$, the three conditions (30) can be compactly written as
$\bm{W}_{\text{S6}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}=\bm{W}_{\text{S6}}{\bm{W}}_{\text{in,1}}.$ (33)
We now show that for any $\bm{W}_{\text{S6}}$, there exists a matrix ${\bm{W}}_{\text{in,1}}$ that satisfies (33). By applying Singular Value Decomposition (SVD) to $\bm{W}_{\text{S6}}$ , we obtain:
$\bm{W}_{\text{S6}}={\bm{U}}\begin{bmatrix}{\bm{\Sigma}}&{\bm{O}}_{(2H+R)\times(D-2H-R)}\end{bmatrix}{\bm{V}}^{\top},$ (34)
where ${\bm{U}}\in{\mathbb{R}}^{(2H+R)\times(2H+R)}$, ${\bm{\Sigma}}\in{\mathbb{R}}^{(2H+R)\times(2H+R)}$, and ${\bm{V}}\in{\mathbb{R}}^{D\times D}$. The diagonal elements of ${\bm{\Sigma}}$ are in decreasing order. We let
${\bm{W}}_{\text{in,1}}={\bm{V}}\begin{bmatrix}{\bm{\Sigma}}^{-1}{\bm{U}}^{\top}\bm{W}_{\text{S6}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}\\ {\bm{Q}}\end{bmatrix},$ (35)
where ${\bm{Q}}\in{\mathbb{R}}^{(D-2H-R)\times D}$ is an arbitrary matrix to be determined later. Plugging (34) and (35) back to $\bm{W}_{\text{S6}}{\bm{W}}_{\text{in,1}}$ and simplifying results in
$\bm{W}_{\text{S6}}{\bm{W}}_{\text{in,1}}$ (36)
$={\bm{U}}\begin{bmatrix}{\bm{\Sigma}}&{\bm{O}}_{(2H+R)\times(D-2H-R)}\end{bmatrix}{\bm{V}}^{\top}{\bm{V}}\begin{bmatrix}{\bm{\Sigma}}^{-1}{\bm{U}}^{\top}\bm{W}_{\text{S6}}^{\star}{\bm{W}}_{\text{in,1}}^{\star}\\ {\bm{Q}}\end{bmatrix}$ ((34) & (35)) (37)
$=\bm{W}_{\text{S6}}^{\star}{\bm{W}}_{\text{in,1}}^{\star},$ (Simplifying) (38)
which demonstrates that (33) is satisfied and completes the proof. ∎
Empirical Validation.
To experimentally verify Lemma 1, we conduct a small-scale experiment. Specifically, we train Mamba 130M on three GLUE tasks—RTE, MRPC, and CoLA—for ten epoch under two settings: (1) training only the linear projection layer ($\bm{W}_{\text{in}}$), and (2) training the S6 modules ($\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$). We experiment with various learning rates and, for each configuration, repeated the best-performing setting five times to ensure robustness. As shown in Fig. 3 (training loss) and Table 15 (validation metrics), our results confirm that optimizing only the linear projection layer is as expressive as training the S6 layers. In fact, in all cases, training only the linear projection not only matches, but even outperforms S6 layer training and converges more quickly.
Figure 3: Training cross-entropy loss, with the shaded area indicating the standard deviation, for fine-tuning linear projection layers versus S6 layers. The results show that $\bm{W}_{\text{in}}$ matches the expressivity of $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\uparrow}$, while also achieving faster convergence.
Layers RTE MRPC CoLA
$\bm{W}_{\text{in}}$ 69.9 ± 1.2 82.4 ± 0.7 61.1 ± 2.3
$\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, ${\bm{W}}_{\bm{\Delta},\uparrow}$ 68.3 ± 0.8 79.9 ± 1.2 54.9 ± 1.5
Table 15: Mean and standard deviation of validation metrics for fine-tuning linear projection layers versus S6 layers. The results demonstrate that $\bm{W}_{\text{in}}$ effectively captures the expressivity of $\bm{W}_{{\bm{B}}}$, $\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\uparrow}$, and achieves superior performance on validation metrics—accuracy for RTE and MRPC, and Matthews correlation coefficient for CoLA.
Appendix D Details of Sec. 5: SDT
D.1 Understanding the Roles of State Matrix A, Input Transition Vector B, and Output Mapping Vector C for a Single Channel in S4 Modules
Problem Setting.
Inspired by Zeng & Lee (2024)’s theoretical analysis of LoRA’s expressive power, we adopt a similar framework to explore the expressive potential of various parameters in the S4 model. Specifically, we assume a target model that performs well on the intended task and a frozen model, which may be either pretrained or randomly initialized. Our goal is to identify a parameter-efficient method to update the frozen model so that it becomes functionally equivalent to the target model. In alignment with Zeng & Lee (2024), we assume that the frozen model’s capacity is equal to or exceeds that of the target model. This assumption is based on two main considerations: (i) analytical tractability, which necessitates that the frozen model must have the potential to match the functionality of the target model, and (ii) a practical rationale, given that the models typically used in practice are often overparameterized. Assume that both the target model and the frozen model are S4, with the target model having a hidden state dimension $H_{\star}$ and the frozen model having a hidden state dimension $H\geq H_{\star}$. Meanwhile, suppose that all the hidden dimensions of both models are valid, meaning that none of the parameter elements are zero. The target model, frozen model, and the updated model after tuning the parameters on the frozen model can be formulated using discretized parameters $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$ as follows:
(Target model) $f^{\star}({\bm{x}})_{n}=\sum_{m=1}^{n}{\bm{C}}_{\star}\overline{{\bm{A}}}_{\star}^{m-n}\overline{{\bm{B}}}_{\star}x_{m},\text{ where }\operatorname{diag}(\overline{{\bm{A}}}_{\star}),\overline{{\bm{B}}}_{\star},{\bm{C}}_{\star}\in{\mathbb{R}}^{H_{\star}},$ (39)
(Frozen model) $f_{0}({\bm{x}})_{n}=\sum_{m=1}^{n}{\bm{C}}\overline{{\bm{A}}}^{m-n}\overline{{\bm{B}}}x_{m},\text{ where }\operatorname{diag}(\overline{{\bm{A}}}),\overline{{\bm{B}}},{\bm{C}}\in{\mathbb{R}}^{H},$ (40)
(Updated model) $\hat{f}({\bm{x}})_{n}=\sum_{m=1}^{n}\widehat{{\bm{C}}}\widehat{\overline{{\bm{A}}}}^{m-n}\widehat{\overline{{\bm{B}}}}x_{m},\text{ where }\operatorname{diag}(\widehat{\overline{{\bm{A}}}}),\widehat{\overline{{\bm{B}}}},\widehat{{\bm{C}}}\in{\mathbb{R}}^{H}.$ (41)
Parameter Efficiency Analysis on S4.
Let ${\mathcal{P}}^{H}$ denote the set of all $H\times H$ permutation matrices. Given this formulation, we present our first analysis of parameter efficiency for the S4 model in the following lemma. This analysis is based on the parameters after necessary discretization $(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})$. For the reader’s convenience, we restate Lemma 2 below with minor notational changes to facilitate the proof.
Lemma 0 (Minimal Parameter Adjustment for S4 Fine-Tuning).
Consider the parameters after discretization, i.e., $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, the minimum number of tunable parameters is:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligningused dimensions with target model}}.$
Proof of Lemma 2.
The key idea of this proof is straightforward. To facilitate the analysis and update the frozen model to be equivalent to the target model, we first equalize the number of hidden state dimensions between the two models. This is achieved by expanding the target model’s ${\bm{A}}_{\star}$, ${\bm{B}}_{\star}$, and ${\bm{C}}_{\star}$ to match the $H$ hidden state dimensions of the frozen model, padding the additional $H-H^{\star}$ dimensions with zeros.
Define $\odot$ as the element-wise product. We can express the target model as:
$f^{\star}(\bm{x})_{n}$ $=\sum_{m=1}^{n}\begin{bmatrix}{\bm{C}}_{\star}&{\bm{0}}^{\top}\end{bmatrix}\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}^{n-m}\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}x_{m}$ (43)
$=\sum_{m=1}^{n}\operatorname{diag}\left(\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}\right)^{n-m}\left(\begin{bmatrix}{\bm{C}}_{\star}^{\top}\\ {\bm{0}}\end{bmatrix}\odot\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}\right)x_{m}.$ (44)
Consider any permutation matrix ${\bm{P}}\in{\mathcal{P}}^{H}$. Applying ${\bm{P}}$ to permute the frozen model leaves the model functionally unchanged:
$f_{0}(\bm{x})_{n}$ $=\sum_{m=1}^{n}{\bm{C}}\overline{{\bm{A}}}^{n-m}\overline{{\bm{B}}}x_{m}=\sum_{m=1}^{n}{\bm{C}}{\bm{P}}\left({\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right)^{n-m}{\bm{P}}^{\top}\overline{{\bm{B}}}x_{m}$ (45)
$=\sum_{m=1}^{n}\operatorname{diag}\left({\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right)^{n-m}\left(({\bm{P}}^{\top}{\bm{C}}^{\top})\odot({\bm{P}}^{\top}\overline{{\bm{B}}})\right)x_{m}.$ (46)
Due to the convolution structure of $\overline{{\bm{A}}}$, two models are functionally equivalent if and only if ${\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}$ aligns with $\begin{bmatrix}\overline{{\bm{A}}}_{\star}&{\bm{O}}\\ {\bm{O}}&{\bm{O}}\\ \end{bmatrix}$, and $({\bm{P}}^{\top}{\bm{C}}^{\top})\odot({\bm{P}}^{\top}\overline{{\bm{B}}})$ align with $\begin{bmatrix}{\bm{C}}_{\star}^{\top}\\ {\bm{0}}\end{bmatrix}\odot\begin{bmatrix}\overline{{\bm{B}}}_{\star}\\ {\bm{0}}\end{bmatrix}$ for some $P\in{\mathcal{P}}^{H}$. If they are already matching or partially matched for certain entries, no updates are required for those entries; only the unmatched entries need to be updated. Then, the required trainable parameters for this permutation matrix ${\bm{P}}$ are:
$\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}+\left\lVert\left[{\bm{P}}^{\top}\overline{{\bm{A}}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}+\left\lVert\left[{\bm{P}}^{\top}(\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}.$
Optimizing the permutation matrix ${\bm{P}}\in{\mathcal{P}}^{H}$ yields the desired results. ∎
This lemma highlights the significance of identifying essential hidden state dimensions. The term $\left\lVert\left[{\bm{P}}^{\top}(\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}$ underscores the importance of excluding redundant dimensions. This can be achieved by either directly removing these dimensions from the state matrix $\overline{{\bm{A}}}$, or by updating $\overline{{\bm{B}}}$ or ${\bm{C}}$ to ensure that only the selected hidden state dimensions are utilized during the input transition or output mapping phases. Once redundant dimensions are filtered out, tuning only the essential dimensions is sufficient to align the updated model with the target model.
Furthermore, based on the lemma, the roles of the input transition vector $\overline{{\bm{B}}}$ and ${\bm{C}}^{\top}$ are nearly identical, as they consistently appear together as the combined term $\overline{{\bm{B}}}\odot{\bm{C}}^{\top}$, which is also discussed in Gupta et al. (2022). Consequently, one could opt to tune either $\overline{{\bm{B}}}$ or ${\bm{C}}$ exclusively or alternatively, split the indices into two groups, tuning $\overline{{\bm{B}}}$ for the first group and ${\bm{C}}$ for the second. Both vectors indicate how information from different hidden state dimensions is integrated, whereas $\overline{{\bm{A}}}$ plays a distinct role, determining how the hidden states are stored.
In practice, instead of directly using the discretized parameters $\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}$, S4 is implemented using the continuous parameters ${\bm{A}},{\bm{B}},{\bm{C}}$ with step size $\Delta$. To provide further practical guidance on parameter tuning, the following two lemmas analyze the parameter efficiency of continuous parameters under different discretization methods. Two exemplary methods of discretization are bilinear and zero-order hold (ZOH):
[…]
Lemma 3 (Essential Continuous Parameter Set for S4 with Bilinear Discritization).
Consider the parameters before discretization, i.e., ${\bm{A}},{\bm{B}},{\bm{C}}$, which are subsequently discretized using bilinear discretization. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[\Delta{\bm{P}}^{\top}(\operatorname{diag}({\bm{I}}+\Delta/2{\bm{A}})\odot{\bm{B}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}{\bm{A}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-{\bm{A}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}({\bm{B}}\odot{\bm{C}}^{\top})\right]_{1:H_{\star}}-{\bm{B}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
Proof of Lemma 3.
Combining Lemma 2 and the Bilinear discretization method in (48) yields the desired results. ∎
Lemma 4 (Essential Continuous Parameter Set for S4 with ZOH Discritization).
Consider the parameters before discretization, i.e., ${\bm{A}},{\bm{B}},{\bm{C}}$, which are subsequently discretized using ZOH discretization. To achieve functional equivalence between the updated model and the target model, i.e., $\hat{f}\equiv f^{\star}$, it is sufficient to tune the following number of parameters:
$\min_{{\bm{P}}\in{\mathcal{P}}^{H}}\overbrace{\left\lVert\left[\Delta{\bm{P}}^{\top}(\operatorname{diag}(\exp(\Delta{\bm{A}})-{\bm{I}})\odot{\bm{B}}\odot{\bm{C}}^{\top})\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminating redundant dimensions}}+\overbrace{\underbrace{\left\lVert\left[{\bm{P}}^{\top}{\bm{A}}{\bm{P}}\right]_{1:H_{\star},1:H_{\star}}-{\bm{A}}_{\star}\right\rVert_{0}}_{\text{aligning the state matrix}}+\underbrace{\left\lVert\left[{\bm{P}}^{\top}({\bm{B}}\odot{\bm{C}}^{\top})\right]_{1:H_{\star}}-{\bm{B}}_{\star}\odot{\bm{C}}_{\star}^{\top}\right\rVert_{0}}_{\text{aligning input-output interactions}}}^{\text{aligning used dimensions with target model}}.$
Proof of Lemma 4.
Combining Lemma 2 and the ZOH discretization method in (48) yields the desired results. ∎
The insights provided by Lemma 3 and Lemma 4 are the same as those provided by Lemma 2. The analysis here supports the second step of SDT-P presented in Sec. 5.
D.2 Extension to Deep S4 Models
Our previous analysis focused on single-channel S4 models. We now expand our investigation to more complex scenarios involving deep S4 models for both target and frozen architectures, incorporating $D$ channels and varying layer depths. In this section, in addition to SDT-P, we introduce SDT+. The key distinction between SDT+ and SDT-P lies in their treatment of linear projection matrices. SDT-P only operate on SSM modules and additionally requires applying LoRA to modify the linear projection matrices. In contrast, SDT+ applies SDT-P on SSM modules while also updates the columns of weight matrices corresponding to the updatable channels identified through Alg. 2. It is worth noting that the linear projection matrix updates in SDT+ are inherently low-rank, making it a specialized case of SDT-P combined with LoRA. Our analysis starts with SDT+, and it automatically applies to SDT-P combined with LoRA.
In this analysis, we assume that each input token $x_{t}$ belongs to ${\mathcal{X}}$, a bounded subset of ${\mathbb{R}}^{D}$, and that the length of the input sequence is finite. Let the frozen model have $L$ layers, and the target model have $L^{\star}$ layers, where $L\geq L^{\star}$. Similar to the technique used in Zeng & Lee (2024) and Giannou et al. (2023). The basic idea of updating the frozen model to match the functionality of the target model is to utilize every $\lceil L/L^{\star}\rceil$ layers of the frozen model to approximate every layer of the target model. We start introducing this proof idea from the simplest case where $L^{\star}=1,L=D$. In this scenario, we can simply choose one different channel to tune and maintain all other channels at zero at every layer. The outputs from the various channels of the deep S4 layers are then combined through a residual connection. This proof idea inspires us to perform channel selection and make use of the residual connections, which is the first and third step of SDT-P presented in Sec. 5. Building on this idea, we present the following results for when the target model has only $L^{\star}=1$ layer, and $L=D=2$.
Lemma 5.
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT+, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target one-layer deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$. Then this can be achieved by selectively fine-tuning at most $\lceil D/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
Proof of Lemma 5.
In this proof, we start by considering the case where $L=D$. In this case, we update a single distinct channel for each layer while setting the other channels to zero. Essentially, we modify the frozen model so that each layer corresponds to and functions as an individual channel in the target model. To be more specific, we fully update the first channel in the first layer to match the first channel of the target model, second channel in the second layer to match the second channel of the target model, so on and so forth.
For the $l$-th layer of the frozen model , we append subscript $l$ to all parameters of the deep S4 layer as introduced in (4). For the $d$-th channel, corresponding notations are denoted with a superscript $(d)$. We define the $t$-th intermediate output token of the $l$-th deep S4 layer as ${\bm{z}}_{l,t}\in{\mathbb{R}}^{D}$. Additionally, the updated S4 module in layer $l$ is denoted as $\widehat{\operatorname{S4}}_{l}$, with $\widehat{\operatorname{S4}}_{l,t}$ referring specifically to the sub-function that outputs the $t$-th token. Therefore, for the $t$-th intermediate output token of the $l$-th deep S4 layer of the updated model can be written as
$\begin{split}{\bm{z}}_{l,t}&=\widehat{{\bm{W}}}_{l}\cdot\widehat{\operatorname{S4}}_{l,t}({\bm{z}}_{l-1,1},\ldots,{\bm{z}}_{l-1,t})+\widehat{\bm{\beta}}_{l}+\widehat{{\bm{u}}}_{l}\odot{\bm{z}}_{l-1,t}\\ &=\widehat{{\bm{W}}}_{l}\cdot\begin{bmatrix}\widehat{\operatorname{S4}}_{l,t}^{(1)}(z_{l-1,1}^{(1)},\ldots,z_{l-1,t}^{(1)})\\ \vdots\\ \widehat{\operatorname{S4}}_{l,t}^{(D)}(z_{l-1,1}^{(D)},\ldots,z_{l-1,t}^{(D)})\\ \end{bmatrix}+\widehat{\bm{\beta}}_{l}+\widehat{{\bm{u}}}_{l}\odot{\bm{z}}_{l-1,t},\end{split}$ (51)
where $\widehat{{\bm{W}}}_{l}\in{\mathbb{R}}^{D\times D},\widehat{\bm{\beta}}_{l}\in{\mathbb{R}}^{D}$ are the updated weight and biases of the $l$-th layer of the frozen model, and $\widehat{{\bm{u}}}_{l}\in{\mathbb{R}}^{D}$ is the updated residual connection weight of the frozen model.
For layers $l<L=D$.
We follow the steps provided in Sec. 5 to update the $l$-th layer of the frozen model such that it functionally equivalent to the $l$-th channel of the target model. For the reader’s convinence, we restate our strategies here:
  • (Channel Selection) Select $D^{\prime}\leq D$ ($D^{\prime}=1$ here) important channels for making predictions. Any channel $d$ that is not utilized will have their corresponding ${\bm{C}}^{(d)}$ set to zero, eliminating the need to update parameters for ${\bm{A}}^{(d)}$ and the $d$-th column of ${\bm{W}}$. To be more specific, we let ${\bm{C}}^{(d)}={\bm{0}}$ for all $d\neq l$ in this scenario.
  • (Hidden State Selection) Within the selected channels, select $H^{\prime}\leq H$ important hidden states. For any hidden state that is not used within a selected channel $d$, the corresponding element in ${\bm{C}}^{(d)}$ will be set to zero, thus eliminating the need to tune the corresponding element in ${\bm{A}}^{(d)}$. To be more specific, we can achieve $\widehat{\operatorname{S4}}_{l,t}^{(l)}(\cdot)=\operatorname{S4}_{\star,t}^{(l)}(\cdot)$ by Lemma 2.
  • (Residual and Bias Tuning) Regardless of other selections, SDT consistently tunes the coefficients of residual connections and biases in linear projections, as these components contain a negligible number of parameters. In this scenario, we let $\widehat{\bm{\beta}}_{l}={\bm{0}},\leavevmode\nobreak\ \widehat{{\bm{u}}}_{l}=\begin{bmatrix}\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{l-1\text{ elements}}}&0&\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{D-l\text{ elements}}}\end{bmatrix}^{\top}.$
This construction yields
${\bm{z}}_{l,t}=\begin{bmatrix}z_{l-1,t}^{(1)}&\ldots&z_{l-1,t}^{(l-1)}&\operatorname{S4}_{\star,t}^{(l)}(z_{l,1}^{(l)},\ldots,z_{l,t}^{(l)})&z_{l-1,t}^{(l+1)}&\ldots&z_{l-1,t}^{(D)}\end{bmatrix}^{\top}.$ (52)
Consequently, only the $l$-th channel is active in the $l$-th layer, while all other layers function as identity mappings, propagating the output of the preceding layer without modification.
For layer $l=L=D$.
Based on the setup of the first $L-1$ layers, we have
${\bm{z}}_{L-1,t}=\begin{bmatrix}\operatorname{S4}_{\star,t}^{(1)}(x^{(1)})&\cdots&\operatorname{S4}_{\star,t}^{(L-1)}(x^{(L-1)})&x^{(L)}\end{bmatrix}^{\top}.$ (53)
For the last layer, we let
$\widehat{{\bm{W}}}_{L}={\bm{W}}_{\star},\quad\widehat{\bm{\beta}}_{L}=\bm{\beta}_{\star},\quad\widehat{{\bm{u}}}_{L}={\bm{0}},$ (54)
[…]
It is easy to verify that the output of the updated frozen model is identical to the output of the target model, i.e.,
$\bm{y}_{t}={\bm{z}}_{L,t}={\bm{W}}_{\star}\begin{bmatrix}\operatorname{S4}_{\star,t}^{(1)}(x^{(1)})&\cdots&\operatorname{S4}_{\star,t}^{(L-1)}(x^{(L-1)})&\operatorname{S4}_{\star,t}^{(L)}(x^{(L)})\end{bmatrix}^{\top}+\bm{\beta}_{\star}.$ (56)
Thus far, we have demonstrated that the statement holds when $L=D$. This analysis can be readily extended to cases where $L\neq D$ by tuning $\lceil D/L\rceil$ channels at each layer. For example, when $L=D/2$, we can tune two channels per layer using a construction similar to the one described above. This generalization completes the proof. ∎
Theorem 2 (Expressive Power of SDT+ on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT+, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer.
Proof of Theorem 2.
We update every $\lceil D/L\rceil$ layers of the frozen model to approximate each layer of the target model. By applying Lemma 5 iteratively to each set of $\lceil D/L\rceil$ layers, we obtain the desired result. ∎
Theorem 2 leads to the following result, which represents the deep S4 model case of Theorem 1.
Theorem 3 (Expressive Power of SDT-P on Deep S4 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT-P, any deep S4 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S4 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states on SSM modules, applying rank-$\lceil\frac{L}{L^{\star}}\rceil$ updates on linear projection matrices and updating residual connections and biases at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
Proof of Theorem 1.
Since SDT+ is a special case of SDT-P, Theorem 2 directly implies the desired statement. ∎
D.3 Extension to S6
In this section, we extend the discussion of SDT-P and SDT+ to S6, following the same logic. We begin by proving results for SDT+ in the scenario where the target model consists of only a single layer. In doing so, we extend Theorem 3 to apply to deep S6 models by first generalizing Lemma 5 to Lemma 6.
Lemma 6.
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT+, any deep S6 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target one-layer deep S6 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$. Then this can be achieved by selectively fine-tuning at most $\lceil D/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
Proof of Lemma 6.
To demonstrate this, we can follow the same proof strategy as in the proof of Lemma 5. In particular, the $t$-th intermediate output token of the $l$-th deep S6 layer of the updated model can be similarly written as
$\begin{split}{\bm{z}}_{l,t}&=\widehat{{\bm{W}}}_{l}\cdot\widehat{\operatorname{S6}}_{l,t}({\bm{z}}_{l-1,1},\ldots,{\bm{z}}_{l-1,t})+\widehat{\bm{\beta}}_{l}+\widehat{{\bm{u}}}_{l}\odot{\bm{z}}_{l-1,t}\\ &=\widehat{{\bm{W}}}_{l}\cdot\begin{bmatrix}\widehat{\operatorname{S6}}_{l,t}^{(1)}(z_{l-1,1}^{(1)},\ldots,z_{l-1,t}^{(1)})\\ \vdots\\ \widehat{\operatorname{S6}}_{l,t}^{(D)}(z_{l-1,1}^{(D)},\ldots,z_{l-1,t}^{(D)})\\ \end{bmatrix}+\widehat{\bm{\beta}}_{l}+\widehat{{\bm{u}}}_{l}\odot{\bm{z}}_{l-1,t},\end{split}$ (57)
where $\widehat{{\bm{W}}}_{l}\in{\mathbb{R}}^{D\times D},\widehat{\bm{\beta}}_{l}\in{\mathbb{R}}^{D}$ are the updated weight and biases of the $l$-th layer of the frozen model, and $\widehat{{\bm{u}}}_{l}\in{\mathbb{R}}^{D}$ is the updated residual connection weight of the frozen model.
For layers $l<L=D$.
We follow the steps provided in Sec. 5 to update the $l$-th layer of the frozen model such that it functionally equivalent to the $l$-th channel of the target model. For the reader’s convinence, we restate our strategies here:
  • (Channel Selection) Select $D^{\prime}\leq D$ ($D^{\prime}=1$ here) important channels for making predictions. For any channel $d$ that is not utilized, rather than directly setting the corresponding ${\bm{C}}^{(d)}$ to zero as in the deep S4 model, we instead set $\bm{\beta}^{(d)}_{\bm{\Delta}}$ to be sufficiently large. According to the computation of SSM parameters described in (58), this ensures that $\overline{{\bm{B}}}_{n}^{(d)}$ is set to zero for all $d\neq l$ in this scenario. This approach is equivalent to setting ${\bm{C}}^{(d)}$ to zero, as both result in the channel producing all zeros.
  • (Hidden State Selection) Within the selected channels, select $H^{\prime}\leq H$ important hidden states. For any hidden state that is not used within a selected channel $d$, the corresponding entries in ${\bm{A}}^{(d)}$ will be set to sufficiently small. To be more specific, we can achieve $\widehat{\operatorname{S6}}_{l,t}^{(l)}(\cdot)=\operatorname{S6}_{\star,t}^{(l)}(\cdot)$ by leveraging the discretized parameters. Lemma 2 provides the conditions for this equality to hold by updating $\overline{{\bm{A}}}$, $\overline{{\bm{B}}}$, and ${\bm{C}}$ for the $l$-th channel, where these parameters are computed as follows:
    $\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}\bm{x}_{n}+\bm{\beta}_{\bm{\Delta}}),$ (58)
    $\overline{{\bm{A}}}_{n}^{(d)}=\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)}),\quad\overline{{\bm{B}}}_{n}^{(d)}=\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}\bm{x}_{n},\quad{\bm{C}}_{n}=\bm{W}_{{\bm{C}}}\bm{x}_{n},$ (59)
    Therefore, we can achieve $\widehat{\operatorname{S6}}_{l,t}^{(l)}(\cdot)=\operatorname{S6}_{\star,t}^{(l)}(\cdot)$ by only updating the corresponding values or columns of the weight matrices for each channel and dimension.
  • (Residual and Bias Tuning) Regardless of other selections, SDT+ consistently tunes the coefficients of residual connections and biases in linear projections, as these components contain a negligible number of parameters. In this scenario, we let $\widehat{\bm{\beta}}_{l}={\bm{0}},\leavevmode\nobreak\ \widehat{{\bm{u}}}_{l}=\begin{bmatrix}\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{l-1\text{ elements}}}&0&\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{D-l\text{ elements}}}\end{bmatrix}^{\top}.$
This construction yields
${\bm{z}}_{l,t}=\begin{bmatrix}z_{l-1,t}^{(1)}&\ldots&z_{l-1,t}^{(l-1)}&\operatorname{S4}_{\star,t}^{(l)}(z_{l,1}^{(l)},\ldots,z_{l,t}^{(l)})&z_{l-1,t}^{(l+1)}&\ldots&z_{l-1,t}^{(D)}\end{bmatrix}^{\top}.$ (60)
For the remaining layers, following the same steps leads to the desired results. ∎
Therefore, we similarly obtain the following two results.
Theorem 4 (Expressive Power of SDT+ on Deep S6 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT+, any deep S6 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S6 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer.
Theorem 5 (Expressive Power of SDT-P on Deep S6 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT-P, any deep S6 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S6 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states on SSM modules, applying rank-$\lceil\frac{L}{L^{\star}}\rceil$ updates on linear projection matrices and updating residual connections and biases at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
Combining Theorem 3 and 5 leads to Theorem 1.
D.4 Sparse Dimension Tuning and Pruning (SDT-P)
Algorithm 2 is our extended algorithm, which includes setting dimensions to zero. However, in practical settings, setting channels to zero is not necessary and omitting it reduces number of hyperparameters, as pruning parameters is effectively equivalent to training them to zero.
Algorithm 2 Dimension Selection Algorithm of SDT-P
Input: A small subset of dataset ${\mathcal{D}}$, warmup epochs $E$, number of layers $L$, total channels $D$, total states $H$, state sparsity $\beta_{0}$, channel sparsity $\alpha_{0}$, state update fraction $\beta$, channel update fraction $\alpha$
/* Warmup Epochs */
Perform full update on SSM modules using ${\mathcal{D}}$ for $E$ epochs  /* Categorize dimensions */
for $l=1$ to $L$ do
       /* Set dimensions as zero */ Sort channels based on $\|\overline{{\bm{A}}}^{(d)}\|$   Select final $(1-\beta_{0})D$ channels as zero channels and denote non-zero channels as set ${\mathbb{D}}$for $d\in{\mathbb{D}}$ do
             Sort states based on magnitude of $\bar{A}^{(d)}_{h}$ at each state dimension  Select final $(1-\alpha_{0})H$ states as zero states and denote non-zero states as set ${\mathbb{H}}$
      /* Unfreeze dimensions */ Sort non-zero channels ${\mathbb{D}}$ based on changes of $\|\overline{{\bm{A}}}^{(d)}\|$  Select the top $\beta|{\mathbb{D}}|$ channels as updatable, denoted by ${\mathbb{D}}^{\prime}$for $d\in{\mathbb{D}}^{\prime}$ do
             Sort non-zero state dimensions based on changes of $\|\overline{{\bm{A}}}^{(d)}\|$  Select the top $\alpha|{\mathbb{H}}|$ states as updatable at the $d$-th channel
D.5 Extension to S5
In this part, we extend Lemma 2 and Theorem 1 to two corresponding results for S5. The extension follows the same procedure as the previous proof, so we omit the details here.
Lemma 7 (Minimal Parameter Adjustment for S5 Fine-Tuning).
Assume all hidden dimensions of the target model $f^{\star}$ are non-zero, i.e., all elements of $\operatorname{diag}(\overline{{\bm{A}}}_{\star})\odot\overline{{\bm{B}}}_{\star}^{(d)}\odot{\bm{C}}_{\star}^{(d)}$ are non-zero. To update frozen model $f_{0}$ such that it becomes functionally equivalent to the target model $f^{\star}$, the minimum number of tunable parameters is: $\min_{\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}}\left\lVert\left[\overline{{\bm{A}}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}+\sum_{d=1}^{D}\left(\overbrace{\left\lVert\left[\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}^{(d)}\odot{\bm{C}}^{(d)\top}\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminatingredundant dimensions}}+\overbrace{\left\lVert\left[\overline{{\bm{B}}}^{(d)}\odot{\bm{C}}^{(d)\top}\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}^{(d)}\odot{\bm{C}}_{\star}^{(d)\top}\right\rVert_{0}}^{\text{ aligning remaining dimensions with target model}}\right),$
subject $\textnormal{subject to}\quad(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})\in\{({\bm{P}}^{\top}\overline{{\bm{A}}}_{0}{\bm{P}},{\bm{P}}^{\top}\overline{{\bm{B}}}_{0},{\bm{C}}_{0}{\bm{P}}):{\bm{P}}\text{ is a permutation matrix}\}.$ (61)
Theorem 6 (Expressive Power of SDT-P with LoRA on Simplified SSM-based Models).
Assume all layers use linear activations. Let $f_{0}$ be a frozen deep S4 S5, or S6 model with $L$ layers, each containing $H$ hidden states per channel. Let $f^{\star}$ be a smaller target model of the same type (S4, S5 or S6), with no residual connections, $L^{\star}<L$ layers, and $H^{\star}<H$ hidden states per channel. Then, there exists a set of parameter updates to $f_{0}$ satisfying the following conditions such that for any finite-length input sequence ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{N})$ with ${\bm{x}}_{n}\in{\mathcal{X}}\subset{\mathbb{R}}^{D}$, where ${\mathcal{X}}$ is bounded, the resulting model $f$ satisfies $f({\bm{X}})=f^{\star}({\bm{X}})$:
  1. 1.
    (SDT-P on SSM) In each SSM module, update at most $\lceil DL^{\star}/L\rceil$ channels. Within each updated channel, fine-tune at most $H^{\star}$ hidden states and set the rest to zero.
  2. 2.
    (LoRA on Linear Projections) Apply rank-$\lceil L/L^{\star}\rceil$ updates to each linear projection matrix.
  3. 3.
    (Minimal Additional Updates) Modify only the residual connections, per-layer biases, and the final-layer output projection.
D.6 Memory Usage and Runtime Analysis of SDT
Memory Usage Analysis.
To assess the memory usage of SDT and LoRA, we conducted experiments on four different models, including both SSM and hybrid architectures. For each model and method, a dataset was generated with 2,500 batches of data samples, each batch comprising a random sequence of 1,500 tokens. The simulation was repeated four times, including dataset generation. All experiments were carried out on a single H100 GPU, and the reported metrics represent averages across the four simulations. Consistent with our previous experiments, we used the original hyperparameter settings, ensuring that SDT includes similar trainable parameters than LoRA. The memory usage of LoRA and SDT is presented in Table 16. Our observations indicate that SDT requires less memory than LoRA. This difference can be attributed to the design of the LoRA adapters, which involve matrix multiplication of two low-rank matrices. In contrast, tuning SSM with the same number of parameters does not require any matrix multiplication, resulting in lower memory usage.
Memory Usage (GB) Mamba-130M Mamba-1.4B Jamba-Tiny-319M Jamba-Mini-52B
LoRA 7.753 37.167 7.207 71.986
LoRA & SDT 5.738 26.491 6.605 67.193
Table 16: Memory usage comparison between SDT and LoRA on various models. Bold numbers indicate the lowest memory usage for each model.
To provide a more fine-grained view, we further analyze how sequence length affects peak memory usage for different Mamba model sizes, as shown in Fig. 4. We measure the memory required to process a single training batch with varying context lengths using randomly generated data. Each batch contains four examples, with 90% of tokens used as input and 10% as output (loss is computed only on the output tokens). The experimental settings for both LoRA and SDT follow the setup described in Section 6.2. We evaluate three configurations for each method, matched in parameter budget. In the plot, each line represents the average across the three configurations, and the shaded region for LoRA shows the range (minimum to maximum). SDT shows negligible variance across configurations, so no shading is included. All models are trained for 500 iterations, and results are averaged over these iterations. Experiments were conducted on an NVIDIA H100 80GB GPU.
Figure 4: Peak memory usage during training as a function of context length for different Mamba model sizes. Each line represents the mean across three configurations; shaded regions indicate min–max ranges. SDT is consistently more memory-efficient than LoRA when applying to SSM module.
Runtime Analysis.
We similarly analyze the latency of LoRA and SDT, using the same experimental setup as in Table 16. Fine-tuning with SDT consists of two stages: (1) dimension selection and (2) standard training. In this study, we first compare the runtime of SDT and LoRA during stage 2 (training) and then evaluate the additional runtime introduced by SDT during stage 1 (dimension selection). Our results show that the dimension selection stage adds only marginal runtime overhead, and SDT is more efficient than LoRA in standard training.
Training: When the channels and states have been selected, the training of SDT is faster than LoRA when the same number of trainable parameters are considered.
The runtimes are reported in Table 17. We observe that, despite having similar trainable parameters, SDT is faster than LoRA. We attribute this to the fact that LoRA introduces additional FLOPs due to the extra matrix multiplication operations required for each update (specifically, the multiplication of two low-rank matrices).
Avg. Runtime (Seconds) Mamba-130M Mamba-1.4B Jamba-Tiny-319M Jamba-Mini-52B
LoRA 410.0 $\pm$ 80.0 2060.0 $\pm$ 135.0 352.5 $\pm$ 107.5 3427.5 $\pm$ 185.0
LoRA & SDT 330.0 $\pm$ 77.5 1697.5 $\pm$ 87.5 257.5 $\pm$ 72.5 3065.0 $\pm$ 232.5
Table 17: Runtime comparison of SDT and LoRA during stage 2 (training).
Dimension Selection: For dimension selection, our method first performs Initial Subset Training, and then selects the dimensions based on the magnitude of parameter changes across different dimensions.
  1. 1.
    Initial Subset Training: We update the model by going through only a subset of the dataset (e.g., 3% of batches in DART experiments), which is sufficient in practice.
  2. 2.
    Magnitude-Based Dimension Selection: After the subset training, we select dimensions based on the magnitude of parameter changes observed.
In this experiment, we simulate a real scenario using a dataset with 2,500 batches, considering a small subset containing 125 batches (5% of the full dataset). We repeat the experiments 80 times, and the reported numbers are averaged across these simulations. Table 18 demonstrates that the dimension selection stage adds only negligible runtime.
Avg. Runtime (Seconds) Mamba-130M Mamba-1.4B Jamba-Tiny-319M Jamba-Mini-52B
Initial Subset Training 16.250 $\pm$ 3.880 85.250 $\pm$ 5.130 15.750 $\pm$ 1.000 163.630 $\pm$ 10.120
Magnitude-Based Dimension Selection 0.280 $\pm$ 0.000 0.520 $\pm$ 0.120 0.090 $\pm$ 0.000 0.240 $\pm$ 0.040
Total Time 16.530 $\pm$ 3.880 85.770 $\pm$ 5.250 15.840 $\pm$ 1.000 163.870 $\pm$ 10.160
Proportion of Training 1 Epoch 0.050$\times$ 0.051$\times$ 0.062$\times$ 0.053$\times$
Proportion of Training 5 Epoch 0.010$\times$ 0.010$\times$ 0.012$\times$ 0.011$\times$
Table 18: Runtime comparison of SDT and LoRA during stage 1 (dimension selection).
We further examine how runtime varies with sequence length, using the same experimental setup as in the memory analysis (Fig. 4). Our results in Fig. 5 show that SDT consistently outperforms LoRA in training speed when applied to SSM modules.
Figure 5: Average training time per batch across different sequence lengths and Mamba model sizes. Each line represents the mean across three configurations; shaded regions indicate min–max ranges. SDT is consistently faster than LoRA when applying to SSM module.
Appendix E Expanded Sec. 6: Evaluation of SDT
E.1 Experiments on Deep S4 Models
Synthetic.
For selecting channels and hidden states, we initiate with a warmup learning rate between $21e-$ 2 and $31e-$ 3 and conduct 20 warmup iterations. Learning rates are adjusted between $25e-$ 2, $21e-$ 2, $35e-$ 3, and $31e-$ 3. We apply LoRA with ranks of 2 and 4 to the SSM and with ranks of 4, 8, and 16 to the linear projection matrices. Non-zero states are selected from the sets {4, 8}, and non-zero channels from {8, 16}.
In addition, we compare the convergence speed of LoRA and SDT in terms of training loss for sequence lengths in $\{100,500,1000\}$. We plot the MSE of both methods against wall-clock time. As shown in Fig. 6, SDT consistently converges to a lower loss faster than LoRA across all tested sequence lengths.
Figure 6: Comparison of SDT and LoRA for tuning S4 in deep S4 models, where LoRA is applied to linear projection matrices. Results are shown across varying sequence lengths under different time budgets in synthetic experiments.
CIFAR-10 (Krizhevsky et al., 2009).
Previous work (Dinh et al., 2022) demonstrates that large language models can be fine-tuned for image classification tasks. In this study, we investigate the adaptation of SSMs for computer vision, focusing on experiments conducted with the CIFAR-10 dataset (Krizhevsky et al., 2009). We employ an eight-layer deep S4 model with a hidden state dimension of 16 and a model dimension of 64. Since pretrained deep S4 models are not available, we simulate a pretrained scenario by fully updating the model for 50 epochs first, then subsequently evaluating the PEFT methods over an additional 5 epochs. We adhere to the preprocessing steps for CIFAR-10 as outlined by Gu et al. (2022a). The LoRA ranks for linear projection matrices are tuned among {1, 2, 4, 8, 16}, and for the S4 component, ranks are set from {1, 2, 4}. Non-zero states are chosen from {8, 12, 16}, and non-zero channels from {48, 64}. The warmup phase includes 1 epoch with a learning rate of $21e-$ 2. For linear projection matrices, LoRA ranks are explored at {2, 4, 8, 16}, and for the SSM, ranks at {2, 4, 8}. All state dimensions are updated, and channel dimensions considered for updates are {4, 8, 16, 32}. The results, as reported in Table 19, indicate that SDT outperforms LoRA with fewer trainable parameters.
Method # Params (%) Accuracy
Frozen 0.00 73.9
LoRA (Proj) 16.00 77.6
LoRA (S4+Proj) 15.52 77.6
LoRA & SDT 11.17 78.0
Full Fine-Tuning 100.00 77.6
Table 19: Accuracy comparison between SDT and LoRA on deep S4 models for CIFAR-10 (Krizhevsky et al., 2009).
E.2 Experiments on Mamba-II, Jamba, and LoRA$+$
Additional Experimental Details.
In this paragraph, we provide further experimental details. Unless otherwise stated, our experiment setting is identical to Sec. C.1. For LoRA, we consider three different LoRA configurations at each layer, targeting the primary parameters of Mamba. Specifically, we focus on the following matrices: $\bm{W}_{\text{out}}$ (output linear projection), $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$ (weight matrices for computing input-dependent ${\bm{B}}_{n},{\bm{C}}_{n}$), and ${\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\bm{\Delta},\uparrow}$ (down and up projection matrices of LoRA adapters for computing $\bm{\Delta}$). The three LoRA application methods are: (i) $\bm{W}_{\text{out}}$, $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\bm{\Delta},\uparrow}$; (ii) $\bm{W}_{\text{out}},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$ and ${\bm{W}}_{\bm{\Delta},\downarrow}$; and (iii) $\bm{W}_{\text{out}}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$. For SDT, we set the channel freeze ratio at 99% across all scenarios. We select the state freeze ratio $\alpha$ from the set $\{75\%,90\%,95\%\}$ and apply LoRA exclusively to $\bm{W}_{\text{out}}$ to maintain a comparable number of trainable parameters. Residual connections and biases are frozen in this experiment. For the warmup, we employ 500 data batches to fully train the SSM modules prior to dimension selection, except for the RTE task in GLUE, where we use 250 batches due to its limited dataset size. Note that the parameters are reverted back after the warmup stage.
Additional Results on Mamba-II.
For Mamba-II, applying SDT is not straightforward because Mamba-II further constrains $A$ such that all (non-zero) entries must have the same value. Therefore, our original dimension selection approach cannot be directly applied here. We consider a naive extension of SDT by selecting dimensions in the projection matrices for input mapping vector $B$ and the projection matrices for output mapping vector $C$ using their respective magnitude, and fine-tune the selected dimensions and all elements of state transition matrix $A$.
Tables 20 and 21 compare the performance on Mamba-II. The results demonstrate that SDT consistently outperforms LoRA on Mamba-II models.
Model Mamba-II-130M Mamba-II-1.3B
Dataset Params (%) DART Params (%) SAMSum Spider
Metric ($\uparrow$) METEOR BLEU R1 R2 RL Acc.
LoRA 0.3354 68.71 48.09 0.1614 49.73 26.14 41.53 72.36
LoRA & SDT 0.3393 70.60 48.93 0.1767 50.72 27.21 42.54 84.15
Table 20: Performance comparison between SDT and LoRA on Mamba-II-130M and Mamba-II-1.3B. Bold numbers indicate the best performance for each task.
Model Mamba-II-130M
Dataset Params (%) GLUE
Accuracy ($\uparrow$) RTE MRPC SST2 QNLI QQP MNLI
LoRA 0.3354 63.4 80.9 89.1 85.3 87.1 78.6
LoRA & SDT 0.3393 64.3 82.3 94.1 87.0 88.3 81.1
Table 21: Performance comparison between SDT and LoRA on the GLUE (Wang et al., 2019) benchmark using Mamba-II-130M. Bold numbers indicate the best performance for each task.
Additional Results on Jamba.
Table 22 shows results for SDT and LoRA on additional datasets. Even though the performance improvement is smaller, our method outperforms pure LoRA in most cases. Mamba layers make up only a small part of Jamba, which is a possible reason for smaller performance gains.
LinProj S6 GLUE DART CelebA SAMSum Spider
Avg. BLEU MET. Acc. R1 R2 RL Acc.
LoRA LoRA 65.5 52.9 73.0 88.5 56.4 33.5 47.9 90.7
SDT 67.7 53.1 73.0 88.4 56.5 33.5 48.0 89.8
Table 22: Performance comparison between SDT and LoRA on pretrained Jamba models. Bold numbers indicate the best performance for each task. We use Jamba-Tiny-319M to compare the performance of SDT and LoRA on GLUE (Wang et al., 2019), and CelebA (Liu et al., 2015) benchmarks. For all other datasets, we employ Jamba-Mini-52B. We report only the best setting out of three for each method.
Additional Results for LoRA+.
We extend our investigation to include LoRA+ (Hayou et al., 2024) with SDT and evaluate its performance against LoRA+ across various datasets on both Mamba-I and Mamba-II. The results, presented in Table 23, show that integrating SDT with LoRA+ enhances its effectiveness and achieves superior performance compared to using LoRA+ alone.
Model Mamba-I-130M Mamba-II-130M Mamba-II-1.3B
Dataset DART DART SAMSum Spider
Metric ($\uparrow$) METEOR BLEU METEOR BLEU R1 R2 RL Acc.
LoRA+ 70.06 50.91 69.78 49.14 49.83 26.09 41.66 73.75
LoRA+ & SDT 70.58 51.93 70.48 49.99 50.81 27.19 42.4 84.22
Table 23: Performance comparison between LoRA+ and SDT on Mamba-I and Mamba-II. Bold numbers indicate the best performance for each task. We test all experiments under various parameter settings (<0.4%) for both LoRA+ and SDT, and report the best values.